### v1

# Phase transition on the fluctuation of the structure of random walk ranges

Abstract.
We investigate fluctuation phenomena for the graph distance and the number of cut points associated with random media arising from the range of a random walk. Our results demonstrate a sequence of dimension-dependent phase transitions in the scaling behavior of these fluctuations, leading to qualitatively different regimes across dimensions lower than six, equal to six, and higher than six. In particular, in six dimensions, we show that the convergence to the normal distribution occurs under a scaling different from that of the standard central limit theorem. We also give the related results for the effective resistance of the random walk range.
Research supported in part by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists (No. JP24K16931) (I.O.).
Research partially supported by NSF grant DMS 2102842
1. Random walk on random walk range
Consider the simple random walk $(S_{n})_{n=0}^{\infty}$ on ${\mathbb{Z}}^{d}$ for $d\geq 4$. We define the graph of the random walk range as follows:
$V({\mathcal{G}}):=\{S_{m}:m\geq 0\},$
$E({\mathcal{G}}):=\{\{S_{m},S_{m+1}\}:m\geq 0\}.$
Originally, Lawler and Burdzy [5] discussed the exponents for the fluctuation of the random walk range and proposed many open problems. Croydon and Shiraishi [11, 20, 21] estimated the graph distance, the effective resistance of $(V({\mathcal{G}}),E({\mathcal{G}}))$ and the number of cut points of the random walk range between $S_{0}$ and $S_{n}$ in $d=4$, and showed that their typical order is $n/(\log n)^{1/2+o(1)}$. In addition, Croydon [7] did similar work in $d\geq 5$, and showed that their typical order is linear. In [22], he discussed the loop erased random walk and other related concepts for the graph of the random walk range in $d=2,3$.
As our motivation, random walks in random media have attracted considerable interest as mathematical models for transport phenomena in disordered environments. Typical examples include random conductance models, random walks in random environments, and random walks on percolation clusters. One of the central themes in this area is to understand how the geometry and randomness of the underlying medium influence the long-time behavior of the walk (e.g., see [8, 9]). For example, for supercritical percolation clusters on ${\mathbb{Z}}^{d}$ with $d\geq 2$, the walk behaves diffusively, in close analogy with the simple random walk on ${\mathbb{Z}}^{d}$ (e.g., see [3]). The situation is markedly different at criticality. In highly irregular media, such as large critical percolation clusters, it is widely believed that the random walk exhibits subdiffusive behavior (e.g., [19]). In addition, [8, 10] deal with time changes of processes on resistance spaces, which cover the case when the resistance has a non-trivial scaling limit. We refer the reader to [14] for further references on these topics.
An important tool for analyzing random walks in random media is the effective resistance or graph distance associated with the underlying random media. In particular, the behavior of the effective resistance is closely connected to the spectral dimension of the random subset, which characterizes the long-time decay of return probabilities and provides a key indicator of whether the associated random walk exhibits diffusive or anomalous scaling.
Independent of the study of random walks on random media, understanding the graph distance in random media is of intrinsic interest. The graph distance captures the geometry induced by the random environment itself and provides a natural notion of distance that reflects how connectivity and paths are shaped by randomness.
The behavior of graph distance plays a central role in a variety of problems, including volume growth, metric scaling, and universality phenomena in random geometry. In particular, in light of recent developments, it is conjectured that graph distances in several models of random media exhibit non-trivial scaling behavior, potentially related to general universality phenomena. These considerations provide strong motivation for a detailed analysis of graph distances in random media, which is a focus of the present work.
Motivated by these considerations, we aim to understand the fluctuation behavior of the effective resistance and the graph distance in random media. In this paper, we take the range of a random walk as a concrete example of a random medium. According to the results of Croydon [7], in $d\geq 5$ the leading-order behavior is linear regardless of the dimension and does not possess a phase transition, which naturally raises the question of how fluctuations behave around this leading term.
Our analysis reveals that the fluctuation behavior depends sensitively on the dimension. In particular, we find distinct regimes in dimensions four, five, six, and seven and higher. A major phase transition occurs at dimension six: for $d\geq 6$, the fluctuations, when normalized by the square root of the variance, converge in distribution to a Gaussian random variable, whereas in dimensions four and five, the same normalization leads to convergence to zero. We also note that the variance of the normal random variable in $d=6$ that appears in the central limit theorem is non-standard in the sense that it is actually not equal to the variance of the original random variable. Moreover, our proofs suggest that in dimensions four and five, if the fluctuations are rescaled by the appropriate order, they are expected to converge to non-Gaussian random variables.
We will give the definitions necessary to state the main results. We consider the graph ${\mathcal{G}}[a,b]$ as the set of edges constructed by the random walk trace:
$E({\mathcal{G}}[a,b]):=\{\{S_{m},S_{m+1}\}:a\leq m\leq b-1\}.$
The graph distance between two points, as is standard, counts the number of edges in the shortest path connecting them in the graph. In addition, we define the number of cut points on $\mathcal{G}[1,n]$ by
$\sum_{i=2}^{n}1\{S[1,i-1]\cap S[i,n]=\emptyset\}.$
We define a quadratic form $\mathcal{E}$ on functions $f,g:V(G)\to\mathbb{R}$ by
$\mathcal{E}(f,g)=\frac{1}{2}\sum_{\begin{subarray}{c}x,y\in V\\ \{x,y\}\in E\end{subarray}}(f(x)-f(y))(g(x)-g(y)).$
This corresponds to the energy dissipation of the resistance electrical network on $G$. We also introduce the space
$H^{2}:=\{\,f\in\mathbb{R}^{V}:\mathcal{E}(f,f)<\infty\,\},$
which serves as the natural domain of the energy form. The effective resistance between $A$ and $B$ is defined via the energy minimization
$R_{G}(A,B)^{-1}\;=\;\inf\Biggl\{\mathcal{E}(f,f)\;\Big|\;\quad f|_{A}=1,\ f|_{B}=0,f\in H^{2}\Biggr\}.$
This quantity represents the voltage difference required to send the current from $A$ to $B$ and therefore corresponds to the usual notion of effective resistance in an electrical network. For vertices $x,y\in V$, we write $R_{G}(x,y)$ to denote $R_{G}(\{x\},\{y\})$.
We will now define the notation used to represent the graph distance, the effective resistance, and the number of cut points. Let $X^{1}_{[a,b]}(x,y):=d_{\mathcal{G}_{[a,b]}}(x,y)$, $X^{2}_{[a,b]}(i,j)$ be the number of cut points on $\mathcal{G}_{[a,b]}$ between the time $i$ and $j$, and $X^{3}_{[a,b]}(x,y):=R_{\mathcal{G}_{[a,b]}}(x,y)$. A trivial relationship among these variables is $X^{3}_{[a,b]}(x,y)\leq X^{1}_{[a,b]}(x,y)$ and $X^{2}_{[a,b]}(i,j)\leq X^{3}_{[a,b]}(S_{i},S_{j})$. We sometimes write $X^{i}[a,b]$ as $X^{i}_{[a,b]}(S_{a},S_{b})$ and $X^{i}_{n}=X^{i}_{[0,n]}(S_{0},S_{n})$ for $i=1,3$ and $X^{2}[a,b]$ as $X^{2}_{[a,b]}(a,b)$ and $X^{2}_{n}=X^{2}_{[1,n]}(1,n)$. Since the following proof is the same, we omit the symbol “ii”, wherever it appears. To ease the presentation, we omit hereafter the integer-part symbol $\lceil\cdot\rceil$. If there exists $c\in(0,\infty)$ such that $ca_{n}\leq b_{n}\leq c^{-1}a_{n}$, we write $a_{n}\asymp b_{n}$. In addition, if $a_{n}/b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\sim b_{n}$. If $\log a_{n}/\log b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\approx b_{n}$. $a_{n}\gg b_{n}$ means $a_{n}/b_{n}\to\infty$ as $n\to\infty$ and $a_{n}\ll b_{n}$ means $b_{n}/a_{n}\to\infty$ as $n\to\infty$. We write $f(n)\lesssim g(n)$ if there exists a (deterministic) constant $c>0$ such that $f(n)\leq cg(n)$ for all $n$, and $f(n)\gtrsim g(n)$ if $g(n)\lesssim f(n)$.
First, we will give our main results for the graph distance and the number of cut points:
Proposition 1.1.
It holds that, as $n\to\infty$,
$\mathrm{Var}(X_{n})\quad\begin{cases}\sim\alpha^{i}_{d}n&\text{for }d\geq 7,\\ \asymp n\log n&\text{for }d=6,\\ \asymp n^{3/2}&\text{for }d=5,\\ =\frac{n^{2}}{(\log n)^{2\pm o(1)}}&\text{for }d=4,\end{cases}$
for some constant $\alpha^{i}_{d}>0$ for $i=1,2$.
Theorem 1.2.
For $d\geq 7$, as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
and for $d=6$, there is a deterministic function $\phi(n)\asymp n\log n$ such that as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\phi(n)}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
For $d=4,5$, as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
Next, we will give the result for the effective resistance.
Proposition 1.3.
For $d\geq 7$ and $\alpha^{3}_{d}>0$, as $n\to\infty$,
$\mathrm{Var}(X_{n})\sim\alpha^{3}_{d}n.$
For $d=6$,
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n\log n}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n(\log n)^{2}}<\infty.$
For $d=5$,
$\mathrm{Var}(X_{n})\approx n^{3/2}.$
For $d=4$ and $\epsilon>0$,
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{2+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{1-\epsilon}}<\infty.$
Theorem 1.4.
For $d\geq 7$, as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
for $d=6$, there is a subsequence $(n_{k})$ and $\phi(n)\geq cn(\log n)$ for some $c>0$ such that as $n\to\infty$,
$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\phi(n_{k})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
For $d=5$, as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0,$
and for $d=4$, there is a subsequence $(n_{k})$ such that as $n\to\infty$,
$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\mathrm{Var}(X_{n_{k}})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
Open problem 1.5.
Does a strong law of large numbers hold for $X_{n}$? (In [11], a weak law has been proved on $(V({\mathcal{G}}),E({\mathcal{G}}))$. In [21, Remark 2.2.2], the upper bound for a strong law for the effective resistance was essentially shown.) One of the difficulties in the proof is that $X_{n}$ does not satisfy monotonicity with respect to $n$. Even if this obstacle could be overcome, especially in the case $d=4$, the variance is not much smaller than that of $\mathbb{E}[X_{n}]^{2}$ (they differ only by a factor of $\log n$" in $d=4$), so it seems unlikely that the result can be proved using standard concentration inequalities.
Remark 1.6.
We believe that the stronger results of Propositions 1.1 and 1.2 are true for the effective resistance as well, but we avoid writing the extra technical details for simplicity. Instead, we present a weaker version that holds under weaker assumptions, such as without any correlation assumptions on the cross terms; we believe that the analysis performed in this section is of independent interest and can be readily applicable to any random variable on the random walk range that is governed by heavy tailed behavior.
As a key ingredient of our approach to the analysis of fluctuations, we employ a decomposition of the random walk path. More precisely, we split the trajectory $S[0,n]$ into two subpaths $S[0,n/2]$ and $S[n/2,n]$, and compare the corresponding effective resistances or graph distances. The quantity
X[0,n/2]+X[n/2,n]X[0,n]X[0,n/2]+X[n/2,n]-X[0,n]
naturally arises as error terms.
By iterating this decomposition procedure, we obtain a representation consisting of a sum of finely decomposed random variables, which behave approximately as independent and identically distributed terms, together with accumulated error terms. A key observation is that a phase transition occurs depending on the dimension: while in dimensions $d\geq 6$ the sum of the i.i.d.-type terms dominates the fluctuations, in lower dimensions the error terms become dominant and significantly affect the scaling behavior. Consequently, it is expected that for $d\geq 6$ the fluctuations converge in distribution to a Gaussian random variable, whereas such Gaussian behavior does not persist in lower dimensions.
We outline the main technical difficulties in the proof of our results. For the graph distance and the number of cut points, establishing the validity or failure of the central limit theorem, which constitutes our main theorem, requires a precise and delicate analysis of the variance. In particular, this analysis necessitates a careful control of the correlations among the error terms arising from the decomposition employed in the proof. As a first step, it is necessary to improve the result of [4, Lemma 1.1] by identifying the optimal strategy for events in the large deviation regime that give rise to large error terms. We show that such events correspond to situations in which the random walk forms a large loop. The second step consists in estimating the correlations of these error terms. More precisely, one must show that these correlations are sufficiently small when the time intervals associated with the decomposition that generates the error terms are well separated.
A technical difficulty to show the validity or failure of the central limit theorem is the fact that the large deviation probabilities governing these error terms exhibit heavy-tailed behavior. In fact, a related phenomenon for the decomposition was already observed in earlier work of the authors [1, 2, 12] concerning the capacity of the range of a random walk. In that setting, when analyzing fluctuations via a similar decomposition scheme, a phase transition occurred between dimensions four and five, depending on which term became dominant in the decomposition.
While this qualitative behavior is closely related to the phenomena studied in the present paper, the analysis in our current setting requires substantially different techniques. More specifically, in the previous work the error terms arising from the decomposition became negligible in dimensions $d\geq 5$, whereas in the present case the error terms remain of polynomial order even for $d\geq 6$, necessitating a much more delicate analysis. Moreover, although in dimension four the error terms are leading in both settings, their control in the present work requires global information on the entire path, leading to more intricate estimates.
We also obtained partial results for the effective resistance; however, the resulting bounds are weaker than those for the other quantities. This is due to technical difficulties arising in the estimation of the correlations of the error terms. Nevertheless, even without a detailed analysis of these correlations, once an estimate of the error terms corresponding to an improvement of [4, Lemma 1.1] is available, we are able to obtain an analogy that yields a phase transition for the central limit theorem of the effective resistance under weaker assumptions on the error terms. We believe that this analogy is useful for a broad class of random variables for which a decomposition of the type considered in this work can be carried out.
2. Preparations in $d\geq 5$
In this section, we will establish various estimates that will allow us to understand the cross terms that appear in the course of our analysis. We will give the definitions here. Consider the error term
$E_{n}:=X^{i}[0,n]+X^{i}[n,2n]-X^{i}[0,2n].$
In general, we may write
$E_{n}^{(k,l)}(i)$= $X^{i}\bigg[\frac{2l}{2^{k+1}}n,\frac{2l+1}{2^{k+1}}n\bigg]+X^{i}\bigg[\frac{2l+1}{2^{k+1}}n,\frac{2l+2}{2^{k+1}}n\bigg]-X^{i}\bigg[\frac{l}{2^{k}}n,\frac{l+1}{2^{k}}n\bigg].$
Since the following proofs do not depend on the quantity of interest, we omit the symbol“ii”, wherever it appears.
In the next lemma, we improve the result of [4, Lemma 1.1]. We will use the following asymptotics on the tail probabilities of $E_{n}$ in the following sections.
Lemma 2.1.
Let $d\geq 5$. There exist some constants $0<c,C<\infty$ such that
(2.1) $\mathbb{P}(E_{n}\geq l)\leq Cl^{-d/2+2},\quad\forall l,$
and
(2.2) $\mathbb{P}(E_{n}\geq l)\geq cl^{-d/2+2},\quad an\leq\forall l\leq bn.$
Remark 2.2.
By the same proof, we also have for $l\leq\max(m,n)$,
$\mathbb{P}(X^{i}[0,n]+X^{i}[n,n+m]-X^{i}[0,n+m]\geq l)\leq Cl^{-d/2+2}.$
Proof of (2.1).
For $I\subset{\mathbb{Z}}^{d}$, let
$N(I):=\sum_{n\in I}1\{S(-\infty,n-1]\cap S[n,\infty)=\emptyset\},$
for $I\subset[a,b]\subset{\mathbb{Z}}^{d}$,
$N(I,[a,b]):=\sum_{n\in I}1\{S[a,n-1]\cap S[n,b]=\emptyset\}$
and $T^{1}$ (resp. $T^{-1}$) be the first global cut point after (resp. before) the origin (see the details in [4, Lemma 1.1]). Note that $T^{1}$ is the same distribution as that of $T^{-1}$ and
$\mathbb{P}(E_{n}\geq l)\leq$ $\mathbb{P}(X_{[1,2n]}(n+1,T^{1}(\theta^{n}\omega))\geq l/2-1)+\mathbb{P}(X_{[1,2n]}(n,T^{-1}(\theta^{n-1}\omega))\geq l/2-1)$
$\leq$ $\mathbb{P}(T^{1}\geq l/2-1)+\mathbb{P}(T^{-1}\geq l/2-1),$
then it suffices to show the result for $T^{1}$. In the equation above, we used the notation of [4], where $\omega$ is a variable in the state space that represents the sequence of steps taken by the random walk. $\theta^{n}\omega$ is the translation of the walk by $n$ time steps; namely, the new walk $\hat{S}$ given by $\theta^{n}\omega$ satisfies $\hat{S}_{i}=S_{i-n}$.
We now fix a parameter $J$, whose value will be set sufficiently large later. We then define $L=n/(3J)$ and set $k_{j}:=Lj$ for $0\leq j\leq 3J$.
If we let
$a_{1}:=\sum_{0\leq j\leq 3J-1}\mathbb{P}(S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)\neq\emptyset),$
$a_{2}:=\mathbb{P}(\cap_{j=0}^{3J-1}\{N([1,n])=0,S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)=\emptyset\}),$
we have
(2.3) $\mathbb{P}(T^{1}>n)\leq a_{1}+a_{2}.$
By the same proof in [4, Lemma 1.1] (which essentially computes the probability of having a large loop), we can obtain
(2.4) $a_{1}\leq C(J)n^{-\frac{d}{2}+2}.$
Note that if $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$, $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$ and $N([1,n])=0$, then $N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0$. To see the reason for this, note that $N([1,n])=0$ means that no point in the interval $k_{3j+1},k_{3j+2}-1]$ can be a cut point; thus, for $t\in[k_{3j+1},k_{3j+2}-1]$, we must have that $S(-\infty,t-1]\cap S[t,\infty)\neq\emptyset$. However, since $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$ and $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$, $S[k_{3j},t-1]\cap S[t,k_{3j+3}-1]\neq\emptyset$ holds instead. Hence,
$a_{2}\leq\mathbb{P}\bigg(\bigcap_{j=1}^{J}\{N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\}\bigg).$
For $0\leq j<J$ and $0<\alpha<1$, we have:
$A_{1}^{j,\alpha}=$ $\Big\{S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+n^{\alpha},k_{3j+3}-1]\neq\emptyset\Big\}$
$\cup\Big\{S[k_{3j},\,k_{3j+2}-n^{\alpha}]\cap S[k_{3j+2},\,k_{3j+3}-1]\neq\emptyset\Big\}.$
Then,
$a_{2}$ $\leq\mathbb{P}\left(\bigcap_{j=1}^{J}(A_{1}^{j,\alpha}\cup(A_{1}^{j,\alpha})^{c})\cap\{N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\}\right)$
$\leq\prod_{j=1}^{J}\mathbb{P}(A_{1}^{j,\alpha}\cup((A_{1}^{j,\alpha})^{c}\cap\{N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\})).$
Note that
$\mathbb{P}(A_{1}^{j,\alpha})\leq C(n^{\alpha})^{-\frac{d}{2}+2}.$
Now, we can split $k_{3j+1},k_{3j+2}-1]$” like in (2.3). That is, let $L^{\prime}=n^{\alpha}$, $J^{\prime}=\frac{k_{3j+2}-k_{3j+1}}{2L^{\prime}}$, $k^{\prime}_{\ell}=\ell L^{\prime}$ for $1\leq\ell\leq 2J^{\prime}$. Then, we have
$\mathbb{P}\Big(N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0,(A_{1}^{j,\alpha})^{c}\Big)$
$\quad\leq\bigg\{\sum_{\ell=1}^{2J-1}\mathbb{P}\Big(S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell}-1]\cap S[k_{3j+1}+k^{\prime}_{\ell+1},\,k_{3j+2}-1]\neq\emptyset\Big)\bigg\}(:=\hat{a}_{1})$
$\qquad+\mathbb{P}\Big(\text{for all }1\leq\ell\leq 2J^{\prime},\,S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+2}-1]=\emptyset,\,$
$\quad\quad N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0,$
$\quad\quad S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+k^{\prime}_{1},\,k_{3j+3}-1]=\emptyset,$
$\quad\quad S[k_{3j},\,k_{3j+2}+k^{\prime}_{2J^{\prime}-1}-1]\cap S[k_{3j+2},\,k_{3j+3}-1]=\emptyset\Big)(:=\hat{a}_{2}).$
Note that
$\hat{a}_{2}\leq$ $\mathbb{P}\Big(\text{for all }0\leq\ell\leq 2J^{\prime}+1,\,S(k_{3j},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+3})=\emptyset,\,$
$\quad\quad N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\Big).$
By the same proof as that of [4, Lemma 1.1], we have
$\hat{a}_{1}\leq C\,n^{1-\alpha}\,(n^{\alpha})^{-\frac{d}{2}+2},\quad\hat{a}_{2}\leq C\,n^{-\alpha}.$
Then if we choose $\alpha$ close to $1$, for some $\delta>0$ (it should not depend on $J$),
$\hat{a}_{1}+\hat{a}_{2}\ll n^{-\delta}$
and hence
$a_{2}\leq C(J)(n^{-\delta})^{J}.$
Therefore, if we pick sufficiently large $J$, we obtain $a_{2}\leq Cn^{-\frac{d}{2}+2}$ and in addition the desired result follows by combining (2.3) and (2.4). ∎
Proof of (2.2).
We start by defining a collection of sets which, when they hold together, will imply $E_{n}>cn$ for an appropriate constant $c$. Let
$D_{1}:=\{\ X[n,2n]\geq c_{d}n\},$ $D_{2}:=\{{\rm Cap}(S[2n-sn,2n])\geq c^{\prime}_{d}sn\},$
$D_{3}:=\{\max_{2n-sn\leq i\leq 2n}\|S_{i}-S_{n}\|\leq C_{1}n^{1/2}\},$ $D_{4}:=\{\|S_{sn}-S_{n}\|\leq C_{1}n^{1/2}\}.$
Here, $c_{d}$ and $c_{d}^{\prime}$ are some dimension-dependent constants, (by the law of large numbers for the capacity of the random walk range, we choose $c_{d}^{\prime}$ to satisfy ${\rm Cap}(S[0,n])\sim 2c^{\prime}_{d}n$). Recall the (slightly modified) definition of $T^{1}$ as the first cut point after time $n$. By [7, Lemma 2.2], we know that there is $c_{d}>0$ such that $\mathbb{P}(X^{2}_{(-\infty,\infty)}[T^{1},n+T^{1}]\geq c_{d}n)=1-o(1)$. Noting that $X^{2}_{(-\infty,\infty)}[a,b]$ will be a lower bound for $X[a,b]$ for any times $a$ and $b$ and that $X^{2}_{(-\infty,\infty)}[a,b]$ is monotonically increasing as we increase the interval $[a,b]$, we obtain
$\mathbb{P}(D_{1})\geq$ $\mathbb{P}(T^{1}\leq n+f(n),X^{2}_{(-\infty,\infty)}[T^{1},2n]\geq c_{d}n/2)-\mathbb{P}(T^{1}\geq n+f(n))$
$\geq$ $\mathbb{P}(X^{2}_{(-\infty,\infty)}(0,n-f(n))\geq c_{d}n/2|0\text{ is cut point})-\mathbb{P}(T^{1}\geq n+f(n))=1-o(1)$
if $f(n)=o(n)$ and $f(n)\to\infty$ as $n\to\infty$ with the aid of the estimate in the proof of (2.1). Moreover, we have $\mathbb{P}(D_{2})=1-o(1)$ (see [13]) and for any $\epsilon>0$, there is $C_{1}<\infty$ such that $\mathbb{P}(D_{i})\geq 1-\epsilon$ for $i=3,4$ by the local central limit theorem and the reflection principle for random walks (see [15, Theorem 1.2.1]). Hence, if we pick sufficiently large $C_{1}<\infty$, we have $\mathbb{P}(\cap_{i=1}^{4}D_{i})\geq 1/2$. Let us define the set A:=S[2nsn,2n]A:=S[2n-sn,2n] and $z:=S_{sn}$.
On the event $D_{1}\cap D_{2}$, we see that we have $\max_{x\in A}\|x-z\|\leq 2C_{1}n^{1/2}$. We now reverse the random walk $S_{0},\ldots,S_{sn}$, to create a new random walk $\tilde{S}_{0},\ldots,\tilde{S}_{sn}$, defined as,
$\tilde{S}_{i}:=S_{sn-i}-S_{sn}.$
$\tilde{S}$ is distributed as a random walk starting at the point $z$ and, conditional on the value of $z$, it is independent of $A$.
We now seek to compute the probability that the random walk $\tilde{S}$ intersects the set $A$. To do this, we derive a last hitting time decomposition of the random walk $\tilde{S}$ with the set $A$ before time $sn$, by considering the last time of intersection between $\tilde{S}$ and the set $A$. Then,
$\mathbb{P}^{z}(T_{A}<sn)$ $=\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}>sn-i)$
$\geq\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}=\infty)$
$\geq\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a)\sum_{x\in A}\mathbb{P}^{x}(T_{A}=\infty)$
$=\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a){\rm Cap}(A)\geq c(s)n^{-d/2+1}{\rm Cap}(A).$
In the last line, we used the local central limit theorem. It is important that $\|x-z\|<Cn^{1/2}$, so for times $\frac{sn}{2}\leq i\leq sn$, we know that $\min_{a\in A}\mathbb{P}^{z}(\tilde{S}_{i}=a)\geq\frac{c}{n^{d/2}}$, for some constant $c$ (cf. [15, Theorem 1.2.1]).
On the event $D_{2}$, we assumed that the capacity of $A$ is greater than $sn$. Thus, conditionally on the intersection of the events $D_{2}$, $D_{3}$, and $D_{4}$, we can derive a lower bound on the probability of intersection between $\tilde{S}$ (which is also $S[1,sn]$) and A (which is S[2nsn,2n]S[2n-sn,2n]) as
$\mathbb{P}(S[1,sn]\cap S[2n-sn,2n]\neq\emptyset|\cap_{i=1}^{4}D_{i})\geq c(s)n^{-d/2+2}.$
We remark now that on this intersection event, we can derive the following lower bound on the cross term $E_{n}$. First, notice that,
$X[n,2n]\geq c_{d}n,$
$X_{n}\geq 0,$
$X_{2n}\leq 2sn.$
The first line comes from the event $D_{1}$ and the second line is trivial. To see the last line (for the graph distance), notice that if we let $S_{t_{1}}$ and $S_{t_{2}}$ be the time indices of the intersection $\{S_{1},\ldots,S_{sn}\}$ with $\{S_{2n-sn},\ldots,S_{2n}\}$ with $t_{1}\in[0,sn]$ and $t_{2}\in[2n-sn,2n]$, the shortest geodesic must be smaller than $t_{1}+(2n-t_{2})$. Namely, shortest geodesic must be shorter than the path $S_{0},\ldots,S_{t_{1}},S_{t_{2}+1},\ldots,S_{2n}$. Now, $t_{1}$ and $2n-t_{2}$ are both less than $sn$. The same argument applies for the number of cut points. To see how the proof holds for the effective resistance, notice that since the effective resistance decreases upon adding edges, we must have that $X_{2n}$ is less than the effective resistance of the graph consisting only of the points $\{S_{1},\ldots,S_{t_{1}}\}\cup\{S_{t_{2}},\ldots,S_{2n}\}$. The effective resistance of this graph is less than $t_{1}+(2n-t_{2})$, the number of points along this graph.
Thus, we have that $E_{n}=X_{n}+X[n+1,2n]-X_{2n}\geq(c_{d}-2s)n$. Therefore,
$\mathbb{P}(E_{n}\geq 2(c_{d}-s)n)\geq\mathbb{P}(\{S[1,sn]\cap S[2n-sn,2n]\neq\emptyset\}\cap(\cap_{i=1}^{4}A_{i}))\geq c(s)n^{-d/2+2}.$
Remark 2.3.
By (2.1), we obtain $\mathbb{E}[X_{n}]/n\to c_{0}$ for some constant $c_{0}<\infty$ with the aid of Hammersley’s lemma for $d\geq 5$. Since $\mathbb{P}(S[1,\infty)\cap\tilde{S}[0,\infty)\neq\emptyset)<1$ for independent two random walks in $d\geq 5$ (e.g., see [15, Theorem 1.2.1]), we have
$\mathbb{E}[T_{\mathcal{G}_{n}}]=\sum_{i=2}^{n}\mathbb{P}(S[1,i-1]\cap S[i,n]=\emptyset)\gtrsim n.$
Hence, we have $c_{0}>0$.
As a consequence, we obtain the following lemma.
Lemma 2.4 ($d=6$).
We have the following estimates for the error terms $E_{n}$.
$\mathbb{E}[E_{n}]\lesssim(\log n),\quad\mathbb{E}[E_{n}^{2}]\asymp n.$
As a consequence of the first and last inequality above, we have that $\mathrm{Var}(E_{n})\gtrsim n$.
Let $\Omega_{n}$ be the event that $\mathbb{E}[E_{n}]-E_{n}\geq 0$, then, we have that
(2.5) $\mathbb{E}[(E_{n}-\mathbb{E}[E_{n}])^{2}\mathbbm{1}[\Omega_{n}]]\lesssim(\log n)^{2}\ll n.$
The previous equation implies that the largest contribution to the variance of $E_{n}$ occurs when $E_{n}-\mathbb{E}[E_{n}]$ is positive.
Proof.
Most of the previous estimates are a consequence of the layer cake representation of the expectation. For example, we have
$\mathbb{E}[E_{n}]=\sum_{l=1}^{n}\mathbb{P}(E_{n}\geq l)\lesssim\sum_{l=1}^{n}\frac{1}{l}\lesssim(\log n).$
In addition, we have
$\mathbb{E}[E_{n}^{2}]=\sum_{l=1}^{n}l\mathbb{P}(E_{n}\geq l)\lesssim\sum_{l=1}^{n}\frac{1}{l}l\lesssim\sum_{l=1}^{n}1=n.$
Finally, we have,
$\mathbb{E}[E_{n}^{2}]=\sum_{l=1}^{n}l\mathbb{P}(E_{n}\geq l)\gtrsim\sum_{l=an}^{bn}l\frac{1}{l}\gtrsim n.$
Since $E_{n}\geq 0$, we have that $\mathbb{E}[E_{n}]-E_{n}\lesssim\log n$. This gives us the final estimate we desire. ∎
3. Proof of CLT for $d\geq 7$
In this section, we will show the central limit theorem for $d\geq 7$.
3.1. Proof of variance
In this subsection, we will give the estimate of the variance.
Lemma 3.1.
For $d\geq 7$,
$\mathrm{Var}(X^{(i)}_{n})\sim\alpha_{d}^{i}n$
with $\alpha_{d}^{i}>0$ for $i=1,2,3$.
To show it, we first prove the following lemma.
Lemma 3.2.
For $d\geq 7$, the limit
$\lim_{n\to\infty}\frac{\operatorname{Var}(X_{n})}{n}$
exists.
For simplicity, we write $\overline{X}:=X-\mathbb{E}[X]$.
Proof.
We refer to [18, Lemma 6.2].
Define
$a_{k}=\sup\left\{\|\overline{X}_{n}\|_{2}:2^{k}\leq n<2^{k+1}\right\},$
where $\|X\|_{m}:=\mathbb{E}[|X|^{m}]^{1/m}$ for $m\geq 1$. For $k\geq 2$, let $n$ be such that $2^{k}\leq n<2^{k+1}$. Taking expectations and using the triangle inequality, we obtain
$\|\overline{X}_{2n}\|_{2}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2}+2\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right).$
Since $X_{n}$ and $X[n+1,2n]$ are independent, we have
$\|\overline{X}_{n}+\overline{X}[n+1,2n]\|_{2}=\left(\|\overline{X}_{n}\|_{2}^{2}+\|\overline{X}[n,2n]\|_{2}^{2}\right)^{1/2}.$
Using (2.1) and that $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq C(\log n)^{M}$ for $d\geq 8$ and $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq n^{1/4}$ for $d=7$, we find
$\|\overline{X}_{2n}\|_{2}\leq\left(\|\overline{X}_{n}\|_{2}^{2}+\|\overline{X}[n,2n]\|_{2}^{2}\right)^{1/2}+c_{1}n^{1/4}.$
From the definition of $a_{k}$, it follows that
$a_{k}\leq\sqrt{2}a_{k-1}+c_{2}2^{k/4},$
for some constant $c_{2}$. Setting $b_{k}=\frac{a_{k}}{2^{k/2}}$, we get a recursive inequality $b_{k}\leq b_{k-1}+c_{1}2^{-k/4}$. This implies $a_{k}\lesssim 2^{k/2}$, and hence
$\operatorname{Var}(X_{n})\lesssim n.$
Letting $\gamma_{n}=\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{2}]$, we have
$\gamma_{n+m}\leq\gamma_{n}+\gamma_{m}+\delta_{n+m},\quad\gamma_{n+m}\geq\gamma_{n}+\gamma_{m}-\delta_{n+m}^{\prime},$
where $\delta_{n+m},\delta_{n+m}^{\prime}=\mathcal{O}((n+m)^{3/4})$. Indeed, if $n\leq m$ and $E_{n,m}:=X_{n}+X[n,n+m]-X_{n+m}$,
$\gamma_{n+m}=$ $\gamma_{n}+\gamma_{m}+2\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])(E_{n,m}-\mathbb{E}[E_{n,m}])]$
$+$ $\mathbb{E}[(X[n,n+m]-\mathbb{E}[X[n,n+m]])(E_{n,m}-\mathbb{E}[E_{n,m}])]+\mathbb{E}[(E_{n,m}-\mathbb{E}[E_{n,m}])^{2}].$
Then, the Cauchy-Schwarz inequality would give,
$\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])(E_{n,m}-\mathbb{E}[E_{n,m}])]$
$\leq\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{2}]^{1/2}\mathbb{E}[(E_{n,m}-\mathbb{E}[E_{n,m}])^{2}]^{1/2}\lesssim n^{1/2}m^{1/4},$
$\mathbb{E}[(X[n,n+m]-\mathbb{E}[X[n,n+m]])(E_{n,m}-\mathbb{E}[E_{n,m}])]\lesssim m^{1/2}m^{1/4},$
$\mathbb{E}[(E_{n,m}-\mathbb{E}[E_{n,m}])^{2}]\lesssim m^{1/2}$
and hence
$\max(\delta_{n+m},\delta_{n+m}^{\prime})\lesssim(m^{1/2}m^{1/4}+m^{1/2}).$
Therefore, by Hammersley’s lemma, the existence of the limit
$\lim_{n\to\infty}\frac{\gamma_{n}}{n}$
follows. ∎
To show that the above limit is positive, we will give a toy model to understand the proof.
Lemma 3.3 (Toy model).
[…]
$\hat{N}_{1}$ $:=|\{i\in[1,T_{m}]:i-1,i+1\not\in\hat{D}_{m},i\in\hat{D}_{m}\}|,$
$\hat{N}_{2}$ $:=|\{i\in[1,T_{m}]:i-1,i+2\not\in\hat{D}_{m},i,i+1\in\hat{D}_{m}\}|.$
Moreover, let
$N_{1}$ $:=|\{i\in[1,n]:i-1,i+1\not\in D_{n},i\in D_{n}\}|,$
$N_{2}$ $:=|\{i\in[1,n]:i-1,i+2\not\in D_{n},i,i+1\in D_{n}\}|.$
Namely, $N_{1}$ denotes the number of events in which $Y_{i}=1$ occurs exactly once and $N_{2}$ the number of events in which $Y_{i}=1$ appears exactly twice in a row. $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge n})_{l=1}^{\infty}$. For large $M$, there exists $c>0$ such that
$A_{n}:=\{|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\quad|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\},$
(3.1) $\mathbb{P}(A_{n})>c>0.$
In addition, for $|a-\mathbb{E}_{m}[\hat{N}_{1}]|\leq\sqrt{m}$, $|b-\mathbb{E}_{m}[\hat{N}_{2}]|\leq\sqrt{m}$ and $c\leq\sqrt{m}$,
[…]
Eventually, for $|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$, $|b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$,
(3.3) $\mathbb{P}^{n}(N_{1}=a,N_{2}=b+\sqrt{n})\leq C\mathbb{P}^{n-\sqrt{n}}(N_{1}=a+\sqrt{n},N_{2}=b).$
Proof.
It is easy to see that $\mathbb{P}(||G_{n}|-\mathbb{E}[|G_{n}|]|\leq\sqrt{n})>c$. Since $(T_{l+1}-T_{l})_{l=1}^{m}$ is i.i.d. and $\mathbb{E}_{m}[\hat{N}_{1}]=p(1-p)m$, $\mathbb{E}_{m}[\hat{N}_{2}]=p^{2}(1-p)m$, by the central limit theorem for the sum of Bernoulli random variables, there exists $c>0$ such that
$\hat{A}_{m}:=\{|\hat{N}_{1}-\mathbb{E}_{m}[\hat{N}_{1}]|\leq\sqrt{m},\quad|\hat{N}_{2}-\mathbb{E}_{m}[\hat{N}_{2}]|\leq\sqrt{m}\},$
$\mathbb{P}(\hat{A}_{m})>c>0.$
We now define $m$ to be the largest value of $l$ such that $T_{l}\leq n$. We consider values of $m$ that satisfy $|m-\mathbb{E}[|G_{n}|]|\leq\sqrt{n}$. Under these conditions, note that if $|\hat{N}_{1}-\mathbb{E}_{m}[\hat{N}_{1}]|\leq\sqrt{m}$, then $|\mathbb{E}[N_{1}]-\mathbb{E}_{m}[\hat{N}_{1}]|+\sqrt{m}=O(\sqrt{n})$ and hence $|N_{1}-\mathbb{E}[N_{1}]|=|\hat{N}_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$ for large $M<\infty$. In addition, if $|\hat{N}_{2}-\mathbb{E}_{m}[\hat{N}_{2}]|\leq\sqrt{m}$, then $|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$ for large $M<\infty$.
Then,
$\mathbb{P}(A_{n})\geq$ $\mathbb{P}(||G_{n}|-\mathbb{E}[|G_{n}|]|\leq\sqrt{n})$
$\times$ $\mathbb{P}\left(|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\bigg|||G_{n}|-\mathbb{E}[|G_{n}|]|\leq\sqrt{n}\right)$
$\geq$ $c\min_{|m-\mathbb{E}[|G_{n}|]|\leq\sqrt{n}}\mathbb{P}\left(|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\bigg||G_{n}|=m\right)$
$\geq$ $c\min_{|m-\mathbb{E}[|G_{n}|]|\leq\sqrt{n}}\mathbb{P}_{m}(\hat{A}_{m})\geq c$
and hence we obtain (3.1).
[…]
Hence, Stirling’s formula yields (3.2). Similarly,
$\frac{\mathbb{P}^{n}(N_{1}=a,N_{2}=b+\sqrt{n})}{\mathbb{P}^{n-\sqrt{n}}(N_{1}=a+\sqrt{n},N_{2}=b)}=\frac{(a+b+\sqrt{n})!}{a!(b+\sqrt{n})!}p^{a}p^{2(b+\sqrt{n})}\times\bigg(\frac{(a+b+\sqrt{n})!}{(a+\sqrt{n})!b!}p^{a+\sqrt{n}}p^{2b}\bigg)^{-1}$
and Stirling’s formula yields (3.3). ∎
Lemma 3.4.
It holds that $\alpha_{d}>0$.
Proof.
Set $D:=\{i\in[0,n/2]:S_{2i}=S_{2i+2}=S_{2i+1}+e\}$, where $e$ is some unit vector, $Y_{i}:=1\{i\in D\}$ and $T_{i}=\{s>T_{i-1}:Y_{s}\neq 1\}$. Let
$N_{1}$ $:=|\{i\in[1,n/2]:i-1,i+1\not\in D,i\in D\}|,$
$N_{2}$ $:=|\{i\in[1,n/2]:i-1,i+2\not\in D,i,i+1\in D\}|,$
and $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge(n/2)})_{l=1}^{\infty}$. By (3.1), there exists $c_{1}>0$ such that
$A_{n}:=\{|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\quad|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\},$
$\mathbb{P}(A_{n})>c_{1}>0.$
In addition, by (3.3), for any $d>0$,
$\mathbb{P}^{n}(X_{n}=d,A_{n})$
$=$ $\sum_{\begin{subarray}{c}|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\\ |b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\end{subarray}}\mathbb{P}^{n}(X_{n}=d,N_{1}=a,N_{2}=b+\sqrt{n})$
$\leq$ $c_{2}\sum_{\begin{subarray}{c}|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\\ |b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\end{subarray}}\mathbb{P}^{n-\sqrt{n}}(X_{n-\sqrt{n}}=d,N_{1}=a+\sqrt{n},N_{2}=b)$
$\leq$ $c_{2}\mathbb{P}(X_{n-\sqrt{n}}=d).$
Now, assume for contradiction that for any $\epsilon>0$, as $n\to\infty$,
(3.4) $\mathbb{P}(|\overline{X}_{n}|\leq\epsilon\sqrt{n})\to 1.$
Since $\mathbb{E}X_{n}/n\to c_{0}$ by Remark 2.3, we have $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}\geq c_{3}$ for some $c_{3}>0$. Indeed, if $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}<c_{3}$ and set $a_{m}=\mathbb{E}X_{m^{2}}$, then $(a_{m}-a_{m-1})/m\leq c_{3}$ for large $m$ and hence $a_{m}\leq c_{0}m^{2}/2$ for large $m$ if we pick small $c_{3}$. We now pick a subsequence $n$ such that $(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})\geq c_{3}\sqrt{n}$ and choose $\epsilon<c_{3}/2$. Then,
$c_{1}/2\leq$ $\mathbb{P}(\{|\overline{X}_{n}|\leq\epsilon\sqrt{n}\}\cap A_{n})$
$\leq$ $c_{2}\mathbb{P}(X_{n-\sqrt{n}}\geq\mathbb{E}X_{n}-\epsilon\sqrt{n})$
$\leq$ $c_{2}\mathbb{P}(X_{n-\sqrt{n}}\geq\mathbb{E}X_{n-\sqrt{n}}+c_{3}\sqrt{n}/2).$
Hence, we obtain
$\mathbb{P}(X_{n-\sqrt{n}}\geq\mathbb{E}X_{n-\sqrt{n}}+c_{3}\sqrt{n}/2)>c$
and this contradicts (3.4). Therefore, we obtain $\liminf_{n\to\infty}\mathbb{P}(|\overline{X}_{n}|\leq\epsilon\sqrt{n})<1$ and hence $\alpha_{d}>0$. ∎
3.2. Proof of CLT
Lemma 3.5.
For $d\geq 7$ and for all $n\in\mathbb{N}$, we have
$\mathbb{E}[(\overline{X}_{n})^{2+\epsilon}]\lesssim n^{1+\epsilon/2}.$
Proof.
Define
$a_{k}=\sup\left\{\|\overline{X}_{n}\|_{2+\epsilon}:2^{k}\leq n<2^{k+1}\right\}.$
Fix $k\geq 2$, and let $n$ satisfy $2^{k}\leq n<2^{k+1}$. Then, by the triangle inequality,
$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}+\|E_{n}\|_{2+\epsilon},$
Using the inequality
$\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}$
$\leq$ $\big(\mathbb{E}[|\overline{X}_{n}|^{2+\epsilon}]+\mathbb{E}[|\overline{X}[n,2n]|^{2+\epsilon}]+C\mathbb{E}[|\overline{X}_{n}|^{1+\epsilon}]\mathbb{E}[|\overline{X}[n,2n]|]$
$\hskip 28.45274pt+C\mathbb{E}[|\overline{X}_{n}|]\mathbb{E}[|\overline{X}[n,2n]|^{1+\epsilon}]\big)^{1/(2+\epsilon)},$
and (2.1) and Lemma 3.2, which gives $\mathbb{E}[|\overline{X}_{n}|^{1+\epsilon}]\mathbb{E}[|\overline{X}[n,2n]|]\lesssim n^{1+\epsilon/2}$, we obtain
$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\left(\mathbb{E}[\overline{X}_{n}^{2+\epsilon}]+\mathbb{E}[\overline{X}[n,2n]^{2+\epsilon}]\right)^{1/(2+\epsilon)}+Cn^{1/2}.$
From the definition of $a_{k}$, it follows that
$a_{k}\leq 2^{1/(2+\epsilon)}a_{k-1}+c_{2}\cdot(2^{k})^{1/2},$
for some constant $c_{2}>0$. Define $b_{k}=2^{-k/(2+\epsilon)}a_{k}$. Then we get
$b_{k}\leq b_{k-1}+c_{2}2^{-k/(2+\epsilon)}(2^{k})^{1/2},$
which implies that $(b_{k})\lesssim 2^{-k/(2+\epsilon)}(2^{k})^{1/2}$. Thus, $a_{k}\lesssim 2^{k/2}$, and hence for all $n$,
$\left(\mathbb{E}[\overline{X}_{n}^{2+\epsilon}]\right)^{1/(2+\epsilon)}\lesssim\sqrt{n}.$
This completes the proof. ∎
Proof of Theorem 1.1.
We decompose $X_{n}$ as follows:
$X_{n}=\sum_{i=1}^{2^{L}}\tilde{X}_{(L,i)}+\mathcal{E},$
where $\tilde{X}_{(L,i)}:=X[(i-1)n2^{-L},in2^{-L}]$ and $\mathcal{E}:=-\sum_{k=0}^{L}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$. Set $L=\lfloor\frac{1}{4}\log_{2}n\rfloor$ so that $2^{L}=n^{1/4}$. Then,
(3.5) $\frac{\mathbb{E}[|\mathcal{E}|]}{\sqrt{n}}\lesssim\frac{n^{1/4}\log n}{\sqrt{n}}\to 0\quad\text{as }n\to\infty.$
Hence, it suffices to prove that
$\frac{1}{\sqrt{n}}\sum_{i=1}^{2^{L}}\left(\tilde{X}_{(L,i)}-\mathbb{E}[\tilde{X}_{(L,i)}]\right)\xrightarrow{d}\mathcal{N}(0,\alpha_{d}).$
We now apply the Lindeberg–Feller central limit theorem. We write $Z_{L,i}:=\tilde{X}_{(L,i)}-\mathbb{E}[\tilde{X}_{(L,i)}]$.
We verify the Lindeberg–Feller conditions:
(i) Variance Condition. By Lemma 3.3, we have
$\sum_{i=1}^{2^{L}}\operatorname{Var}(\tilde{X}_{(L,i)})\sim 2^{L}\cdot\alpha_{d}\cdot\frac{n}{2^{L}}=\alpha_{d}n,$
hence
$\frac{1}{n}\sum_{i=1}^{2^{L}}\operatorname{Var}(\tilde{X}_{(L,i)})\to\alpha_{d}>0.$
(ii) Lindeberg Condition. For any $\epsilon>0$, we must show that as $n\to\infty$,
$\sum_{i=1}^{2^{L}}\frac{1}{n}\mathbb{E}\left[|Z_{L,i}|^{2}\cdot\mathbf{1}_{\{|Z_{L,i}|>\epsilon\sqrt{n}\}}\right]\to 0.$
Using Hölder and Chebyshev’s inequality:
$\mathbb{E}\left[|Z_{L,i}|^{2}\cdot\mathbf{1}_{\{|Z_{L,i}|>\epsilon\sqrt{n}\}}\right]\leq\mathbb{E}[|Z_{L,i}|^{2+\delta}]^{2/(2+\delta)}\mathbb{P}(|Z_{L,i}|\geq\epsilon\sqrt{n})^{\delta/(2+\delta)}.$
By Lemma 3.5,
$\mathbb{E}[|Z_{L,i}|^{2+\delta}]^{2/(2+\delta)}\lesssim\frac{n}{2^{L}}.$
Furthermore, Chebyshev’s inequality should give,
$\mathbb{P}(|Z_{L,i}|>\epsilon\sqrt{n})\leq\frac{\mathrm{Var}(\tilde{X}_{L,i})}{\epsilon^{2}n}\lesssim\frac{n}{\epsilon^{2}n(2^{L})}=\frac{1}{\epsilon^{2}(2^{L})}.$
Thus, we have,
$\mathbb{E}\left[|Z_{L,i}|^{2}\cdot\mathbf{1}_{\{|Z_{L,i}|>\epsilon\sqrt{n}\}}\right]\leq\frac{n}{(\epsilon^{2})^{\delta/(2+\delta)}(2^{L})^{1+\delta/(2+\delta)}}.$
Finally, we can sum up all of the appropriate terms and see that,
$\sum_{i=1}^{2^{L}}\frac{1}{n}\mathbb{E}\left[|Z_{L,i}|^{2}\cdot\mathbf{1}_{\{|Z_{L,i}|>\epsilon\sqrt{n}\}}\right]\lesssim 2^{L}\frac{1}{n}\frac{n}{(\epsilon^{2})^{\delta/(2+\delta)}(2^{L})^{1+\frac{\delta}{(2+\delta)}}}\to 0,$
as $n\to\infty$ and $2^{L}\to\infty$.
Therefore, both Lindeberg–Feller conditions are satisfied, and we conclude
$\frac{\sum_{i=1}^{2^{L}}(\tilde{X}_{(L,i)}-\mathbb{E}[\tilde{X}_{(L,i)}])}{\sqrt{n}}\xrightarrow{d}\mathcal{N}(0,\alpha_{d})$
as $n\to\infty$ and $2^{L}\to\infty$ and hence combining this with (3.5), we see that
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
as $n\to\infty$. ∎
4. 6 Dimensions
In this section, we estimate the variance and the Limit in distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ in $d=6$.
4.1. Lower and Upper Bounds on the Variance
4.1.1. Graph distance and cut points
Now, we produce an upper bound to the variance of the graph distance.
Lemma 4.1.
We have the following upper bound on the variance of $X_{n}$:
$\mathrm{Var}(X_{n})\lesssim n(\log n).$
Proof.
We now perform the full binary decomposition as,
$X_{n}=\sum_{k=0}^{\log_{2}n}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{k=0}^{n}X[k,k+1].$
We define the error term
$\mathcal{E}_{k}:=\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}.$
By independence of the components $E_{n}^{(k,l)}$, we have that,
$\mathrm{Var}(\mathcal{E}_{k})=\sum_{l=1}^{2^{k}}\mathrm{Var}(E_{n}^{(k,l)})\lesssim\sum_{l=1}^{2^{k}}\frac{n}{2^{k}}=n.$
Thus we have, by expanding the variance and applying the Cauchy-Schwarz inequality, that
(4.1) $\mathrm{Var}\left(\sum_{k=0}^{\log_{2}n}\mathcal{E}_{k}\right)\leq\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}n}\sum_{k_{2}=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k_{1}})^{1/2}\mathrm{Var}(\mathcal{E}_{k_{2}})^{1/2}\lesssim n(\log n)^{2}+n.$
In addition, since each $X[k,k+1]$ is deterministically 1, we have that $\mathrm{Var}(X[k,k+1])=0$. Then, we have
$\mathrm{Var}(X_{n})\lesssim n(\log n)^{2}.$
By our binary decomposition, we find that, for $a_{n}=n/(\log n)^{3}$,
$X_{n}=\sum_{l=1}^{a_{n}}X[(l-1)\frac{n}{a_{n}},l\frac{n}{a_{n}}]+\sum_{k=0}^{\log_{2}(a_{n})}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}).$
Now, we claim that
$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim(\log n)^{2}.$
Let $\tilde{S}$ be the reversal of the first part, that is $\tilde{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}-i}-S_{l\frac{n}{2^{k}}}$ . $\hat{S}$ will be the second part, that is $\hat{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}+i}-S_{l\frac{n}{2^{k}}}$. Indeed, if $E_{n}^{(k,l)}\geq a$, $\Omega(a,k,l)$ holds, where
$\Omega(a,k,l)$= $\{\hat{S}^{(k,l)}[a/2,\frac{n}{2^{k}}]\cap\tilde{S}^{(k,l)}[0,\frac{n}{2^{k}}]\neq\emptyset\}(=:A_{1}^{(a,k,l)})$
$\bigcup\{\tilde{S}^{(k,l)}[a/2,\frac{n}{2^{k}}]\cap\hat{S}^{(k,l)}[0,\frac{n}{2^{k}}]\neq\emptyset\}(=:A_{2}^{(a,k,l)}).$
Then,
$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]=\sum_{a,b}\mathbb{E}[\mathbbm{1}[E_{n}^{(k_{1},l_{1})}\geq a]\mathbbm{1}[E_{n}^{(k_{2},l_{2})}\geq b]]\leq\sum_{a,b}\sum_{i,j=1}^{2}\mathbb{P}(A_{i}^{(a,k_{1},l_{1})}\cap A_{j}^{(b,k_{2},l_{2})}).$
In addition,
$\mathbb{P}(A_{1}^{(a,k_{1},l_{1})}\cap A_{1}^{(b,k_{2},l_{2})})$
$\leq$ $\mathbb{P}\bigg(\{\hat{S}^{(k_{1},l_{1})}[\frac{a}{2},\frac{n}{2^{k_{1}}}]\cap\tilde{S}^{(k_{1},l_{1})}[0,\frac{n}{2^{k_{1}}}]\neq\emptyset\}\cap\{\hat{S}^{(k_{2},l_{2})}[\frac{b}{2},\frac{n}{2^{k_{2}}}]\cap\tilde{S}^{(k_{2},l_{2})}[0,\frac{n}{2^{k_{2}}}]\neq\emptyset\}\bigg)$
$\leq$ $\sum_{u_{1}=\frac{a}{2}+l_{1}\frac{n}{2^{k_{1}}}}^{(l_{1}+1)\frac{n}{2^{k_{1}}}}\sum_{u_{2}=(l_{1}-1)\frac{n}{2^{k_{1}}}}^{l_{1}\frac{n}{2^{k_{1}}}}\sum_{u_{3}=\frac{b}{2}+l_{2}\frac{n}{2^{k_{2}}}}^{(l_{2}+1)\frac{n}{2^{k_{2}}}}\sum_{u_{4}=(l_{2}-1)\frac{n}{2^{k_{2}}}}^{l_{2}\frac{n}{2^{k_{2}}}}\sum_{x}\sum_{y}$
$\mathbb{P}(S_{u_{1}}=S_{u_{2}}=x,S_{u_{3}}=S_{u_{4}}=y).$
By [15, Theorem 1.2.1], letting $p(x,n):=\frac{1}{(2d\pi n)^{d/2}}\exp(-\frac{|x|^{2}}{2dn})$ for $d=6$, we have that if $|\frac{2l_{1}-1}{2^{k_{1}+1}}n-\frac{2l_{2}-1}{2^{k_{2}+1}}n|>\frac{n}{2^{k_{2}+1}}$ for $k_{1}<k_{2}$,
$\sum_{a,b}\mathbb{P}(A_{1}^{(a,k_{1},l_{1})}\cap A_{1}^{(b,k_{2},l_{2})})$
$\leq$ $\sum_{a,b}\sum_{u_{1}=a/2+l_{1}\frac{n}{2^{k_{1}}}}^{(l_{1}+1)\frac{n}{2^{k_{1}}}}\sum_{u_{2}=(l_{1}-1)\frac{n}{2^{k_{1}}}}^{l_{1}\frac{n}{2^{k_{1}}}}\sum_{u_{3}=b/2+l_{2}\frac{n}{2^{k_{2}}}}^{(l_{2}+1)\frac{n}{2^{k_{2}}}}\sum_{u_{4}=(l_{2}-1)\frac{n}{2^{k_{2}}}}^{l_{2}\frac{n}{2^{k_{2}}}}\sum_{x}\sum_{y}$
$1_{\{u_{1}\leq u_{2}\leq u_{3}\leq u_{4}\}}Cp(x,u_{1})p(0,u_{2}-u_{1})p(y-x,u_{3}-u_{2})p(0,u_{4}-u_{3})$
$+$ $1_{\{u_{1}\leq u_{3}\leq u_{2}\leq u_{4}\}}Cp(x,u_{1})p(y-x,u_{3}-u_{1})p(x-y,u_{2}-u_{3})p(y-x,u_{4}-u_{2})\ldots$
$\lesssim$ $(\log n)^{2},$
where $\ldots$ represents other possible orderings of the variables $u_{1},u_{2},u_{3},u_{4}$. In addition, if $|\frac{2l_{1}-1}{2^{k_{1}+1}}n-\frac{2l_{2}-1}{2^{k_{2}+1}}n|=\frac{n}{2^{k_{2}+1}}$, by Lemma 2.4,
$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\leq\mathbb{E}[(E_{n}^{(k_{1},l_{1})})^{2}]^{1/2}\mathbb{E}[(E_{n}^{(k_{2},l_{2})})^{2}]^{1/2}\lesssim\frac{n}{2^{k_{2}/2}2^{k_{1}/2}}.$
We can obtain the same result for $\mathbb{P}(A_{i}^{(a,k_{1},l_{1})}\cap A_{j}^{(b,k_{2},l_{2})})$ by a similar computation for other cases. Hence,
$\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]$
$\leq$ $\sum_{k_{1}=0}^{\log_{2}(a_{n})}2^{k_{1}}[a_{n}/2^{k_{1}}](\log n)^{2}\leq a_{n}(\log n)^{3}\ll n(\log n).$
Therefore,
$\mathrm{Var}\left(\sum_{k=0}^{a_{n}}\mathcal{E}_{k}\right)\leq\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim n(\log n)$
and
$\mathrm{Var}(MT)\leq Ca_{n}[n/a_{n}(\log(n/a_{n}))^{2}]\leq Cn(\log\log n)^{2}.$
Finally, we can apply the Cauchy-Schwarz inequality to say that,
$\mathrm{Var}(X_{n})\leq\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})+2\mathrm{Var}(MT)^{1/2}\mathrm{Var}(\mathcal{E})^{1/2}\lesssim n(\log n).$
This completes the proof of the assertion. ∎
We start with the following lower bound.
Lemma 4.2.
There exists $0<c<\infty$ such that
$\mathrm{Var}(X_{n})\geq cn\log n.$
Proof.
Fix $\delta>0$ very close to 1. We will see where this condition plays a role close to the end of the proof.
By our binary decomposition, we find that,
(4.2) $X_{n}=\sum_{l=1}^{n^{\delta}}X[(l-1)n^{1-\delta},ln^{1-\delta}]+\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}).$
We will write
$MT:=\sum_{l=1}^{n^{\delta}}X[(l-1)n^{1-\delta},ln^{1-\delta}],\quad\mathcal{E}:=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}).$
We have that
$\mathrm{Var}(X_{n})=\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})+2\text{Cov}(MT,\mathcal{E})$
and
$\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
In addition, we obtain that, for sufficiently large $n$,
$\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})\geq\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}c\frac{n}{2^{k}}\geq c\delta n\log n.$
By Lemma 2.4, we have
$\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})})=$ $-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]+\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]$
$\geq$ $-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\geq-(\log n)^{2}.$
Furthermore, we remark that if the intervals $n/2^{k_{1}+1}(2l_{1}-2),n/2^{k_{1}+1}(2l_{1})]$ (this is the interval corresponding to the cross term $E^{(k_{1},l_{1})}$) and $n/2^{k_{2}+1}(2l_{2}-2),n/2^{k_{2}+1}(2l_{2})]$(this is the interval corresponding to $E^{(k_{2},l_{2})}$) are completely disjoint from each other, then $E^{(k_{1},l_{1})}$ and $E^{(k_{2},l_{2})}$ are independent of each other and the covariance is 0. Using these considerations, we see that a lower bound for the covariances is the following:
$\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(n^{\delta})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}-(\log n)^{2}\geq-\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}2^{k_{1}}[n^{\delta}/2^{k_{1}}](\log n)^{2}$
$=-n^{\delta}(\log n)^{3}\gg-n(\log n),\text{ if }\delta<1.$
Furthermore, notice that by Lemma 4.1, we have the upper bound $\mathrm{Var}(X_{n})\leq Cn\log n$. Thus, we can show that
$\mathrm{Var}(MT)\leq Cn^{\delta}[n^{1-\delta}\log(n^{1-\delta})]\leq C(1-\delta)n\log n,$
by setting $\delta$ sufficiently close to $1$, this can be made very small.
Finally, we can apply the Cauchy-Schwarz inequality to say that,
$\mathrm{Var}(X_{n})\geq\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})-2\mathrm{Var}(MT)^{1/2}\mathrm{Var}(\mathcal{E})^{1/2}.$
Then for $\delta$ close to but not equal to $1$, we have $\mathrm{Var}(\mathcal{E})\geq cn\log n$ and $\mathrm{Var}(MT)\leq C(1-\delta)n\log n$. Thus, $\mathrm{Var}(X_{n})\gtrsim n\log n$, as desired. ∎
4.1.2. Effective resistance
We start with the following lower bound.
Lemma 4.3.
There exist a constant $\epsilon>0$ and a subsequence $(n_{k})$ such that $\mathrm{Var}(X_{n_{k}})\geq\epsilon n_{k}\log n_{k}$. In other words,
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n\log n}>0.$
Proof.
Assume for contradiction that this is not true. Then, it must be the case that for every $\kappa>0$, there exists $N_{\kappa}$ such that for all $n\geq N_{\kappa}$, we have,
$\mathrm{Var}(X_{n})\leq\kappa n\log n.$
Fix $\delta>0$. For $\delta>0$, we will choose a sufficiently small $\kappa>0$ and derive a contradiction later. Recall our binary decomposition in (4.2). As long as $n^{1-\delta}\geq N_{\kappa}$, we have that
(4.3) $\mathrm{Var}(MT)\leq n^{\delta}[\kappa n^{1-\delta}\log n^{1-\delta}]=\kappa(1-\delta)n\log n$
and
(4.4) $\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
We have that, for sufficiently large $n$,
$\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})\geq\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}c\frac{n}{2^{k}}\geq c\delta n\log n.$
Notice that
$\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})})=\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$
$\geq-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\gtrsim-(\log n)^{2}.$
Here, we only used that $E_{n}\geq 0$ to get the second inequality.
Putting these estimates back into equation (4.4), we observe that
(4.5) $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n-O((\log n)^{2}(n^{\delta}\log(n^{\delta}))^{2})$
This last term will be greater than $\frac{c\delta}{2}n\log n$ for $n$ sufficiently large.
Finally, we can apply the Cauchy-Schwarz inequality to say that,
(4.6) $\mathrm{Var}(X_{n})\geq\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})-2\mathrm{Var}(MT)^{1/2}\mathrm{Var}(\mathcal{E})^{1/2}.$
Notice that as a function of the value of $\mathrm{Var}(\mathcal{E})$, the right-hand side is increasing when $\mathrm{Var}(\mathcal{E})\geq\mathrm{Var}(MT).$ We have already obtained a lower bound on $\mathrm{Var}(\mathcal{E})$ from equation (4.5) that is larger than the upper bound we have on $\mathrm{Var}(MT)$ from equation (4.3).
Indeed, notice that if $\mathrm{Var}(MT)<\frac{c\delta}{8}n\log n$, then we would have that the right hand side of (4.6) will be greater than,
$(\mathrm{Var}(\mathcal{E})^{1/2}-\mathrm{Var}(MT)^{1/2})^{2}\geq\left(\frac{\sqrt{c\delta}}{\sqrt{2}}\sqrt{n\log n}-\frac{\sqrt{c\delta}}{2\sqrt{2}}\sqrt{n\log n}\right)^{2}\geq\frac{c\delta}{8}n\log n.$
Then, we have that by setting $\kappa$ sufficiently small that and applying the previous inequality that $\mathrm{Var}(X_{n})\geq\frac{c\delta}{8}n\log n$. This is a contradiction to the assumption that we cannot produce a lower bound to the $\limsup$ of $\frac{\mathrm{Var}(X_{n})}{n\log n}$. ∎
Now, we produce an upper bound to the variance of the effective resistance. Since the proof is the same as that of (4.1), we omit it.
Lemma 4.4.
We have the following upper bound of the variance of $X_{n}$:
$\mathrm{Var}(X_{n})\lesssim n(\log n)^{2}.$
4.2. The Limit in Distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$
4.2.1. Graph distance and cut points
Proof of Theorem 1.2 for $d=6$.
Part 1: A CLT contribution
Define the value $n_{\delta}:=\exp\left[\frac{1}{2}\log n+(\sqrt{\log n})^{1-\delta}\right]$. Now,
$\mathcal{G}:=\sum_{i=1}^{n_{\delta}}X[(i-1)n/n_{\delta},in/n_{\delta}]-\mathbb{E}[X[(i-1)n/{n_{\delta}},in/n_{\delta}]]$
will be a normal random variable of variance of order $n\log n$. Define
$Z_{n_{\delta},i}:=\frac{X[(i-1)n/n_{\delta},in/n_{\delta}]-\mathbb{E}[X[(i-1)n/n_{\delta},in/n_{\delta}]]}{\sqrt{\mathrm{Var}(X_{n/n_{\delta}})}}.$
The variables $Z_{n_{\delta},i}$ are all independent of each other and one
$\sum_{i=1}^{n_{\delta}}\mathbb{E}\left[Z_{n_{\delta},i}^{2}\right]=1.$
Furthermore, we have that,
$\sum_{i=1}^{n_{\delta}}\mathbb{E}[Z_{n_{\delta},i}^{2}\mathbbm{1}[|Z_{n_{\delta},i}|\geq\epsilon]]=0,$
since deterministically $X[(i-1)n/n_{n_{\delta}},in/n_{\delta}]\leq n/n_{\delta}\ll\sqrt{n(\log n/n_{\delta})}$ holds.
Part 2: Controlling the error term at mesoscopic scales
Define $n_{\delta,-}:=\exp\left[\frac{1}{2}\log n-(\sqrt{\log n})^{1-\delta}\right]$. Let us control
$\mathcal{M}:=\sum_{k=\log_{2}n_{\delta,-}}^{\log_{2}n_{\delta}}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}+\mathbb{E}[E_{n}^{(k,l)}]).$
Since at a given level $k$, all of the $E_{n}^{(k,l)}$ are independent of each other, we can compute
$\mathbb{E}\left[\left(\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}+\mathbb{E}[E_{n}^{(k,l)}])\right)^{2}\right]=\sum_{l=1}^{2^{k}}\mathbb{E}[(-E_{n}^{(k,l)}+\mathbb{E}[E_{n}^{(k,l)}])^{2}]\lesssim\frac{n}{2^{k}}2^{k}=n.$
Thus, we see that $\mathbb{E}[\mathcal{M}^{2}]\leq 4n(\sqrt{\log n})^{2-2\delta}$. Hence, by Markov’s inequality, we have that
(4.7) $\mathbb{P}(\mathcal{M}\geq\sqrt{n(\log n)^{1-\delta/2}})\to 0$
as $n\to\infty$ for any $\epsilon>0$.
Part 3: Controlling the error terms at the smallest scales
We let $\Omega_{k,l}$ be the event that,
$\Omega_{k,l}:=\left\{|E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]|\ \leq\left(\frac{n}{2^{k}}\right)^{1-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}\right\}.$
Notice that for fixed $k$, the events $\Omega_{k,l}$ are all independent of each other. Define $\Omega_{k}:=\bigcap_{l=1}^{2^{k}}\Omega_{k,l}$ Thus, we see that when $2^{k}\leq n_{\delta,-}\leq n^{1/2}$, by (2.5),
$\mathbb{E}\left[\left(\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]\right)^{2}\mathbbm{1}[\Omega_{k}]\right]=\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}])^{2}\mathbbm{1}[\Omega_{k,l}]]$
$+\sum_{l_{1},l_{2}=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l_{1})}-\mathbb{E}[E_{n}^{(k,l_{1})}])\mathbbm{1}[\Omega_{k,l_{1}}]]\mathbb{E}[(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}])\mathbbm{1}[\Omega_{k,l_{2}}]]$
$\leq 2^{k}\left(\frac{n}{2^{k}}\right)^{1-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}+(2^{k})^{2}(\log n)^{2}=n^{1-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}(2^{k})^{\frac{1}{(\sqrt{\log n})^{1+2\delta}}}+n_{\delta,-}^{2}(\log n)^{2}$
$\leq n^{1-\frac{1}{2(\sqrt{\log n})^{1+2\delta}}}+\exp[\log n-\sqrt{\log n}^{1-\delta}+2\log\log n].$
In addition, we also have that when $2^{k}\leq n_{\delta,-}$,
$\mathbb{P}(\Omega_{k}^{c})\leq\sum_{l=1}^{2^{k}}\mathbb{P}(\Omega_{k,l}^{c})\leq\frac{2^{k}}{\left(\frac{n}{2^{k}}\right)^{1-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}}$
$=$ $\frac{(2^{k})^{2-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}}{n^{1-\frac{1}{(\sqrt{\log n})^{1+2\delta}}}}\leq\frac{\exp\left[\left(\frac{\log n}{2}-\sqrt{\log n}^{1-\delta}\right)\left(2-\frac{1}{(\sqrt{\log n})^{1+2\delta}}\right)\right]}{\exp[\log n-(\sqrt{\log n})^{1-2\delta}]}\leq\exp[-(\sqrt{\log n})^{1-\delta}],$
for $n$ sufficiently large.
If we now define $\Omega=\bigcap_{k=1}^{\log_{2}n_{\delta,-}}\Omega_{k}$, then we see that,
$\mathbb{P}(\Omega^{c})\leq\sum_{k=1}^{\log_{2}n_{\delta,-}}\mathbb{P}(\Omega_{k}^{c})\leq\log n\exp[-(\sqrt{\log n})^{1-\delta}]\ll 1.$
Furthermore, if we define
$\mathcal{L}:=\sum_{k=0}^{\log_{2}n_{\delta,-}}\sum_{l=1}^{2^{k}}(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]),$
we have
$\mathbb{E}\left[\mathcal{L}^{2}\mathbbm{1}[\Omega]\right]$
$\leq\sum_{k_{1},k_{2}=0}^{\log_{2}n_{\delta,-}}\mathbb{E}\left[\left(\sum_{l=1}^{2^{k_{1}}}E_{n}^{(k_{1},l)}-\mathbb{E}[E_{n}^{(k_{1},l)}]\right)\left(\sum_{l=1}^{2^{k_{2}}}E_{n}^{(k_{2},l)}-\mathbb{E}[E_{n}^{(k_{2},l)}]\right)\mathbbm{1}[\Omega]\right]$
$\leq\sum_{k_{1},k_{2}=0}^{\log_{2}n_{\delta,-}}\mathbb{E}\left[\left(\sum_{l=1}^{2^{k_{1}}}E_{n}^{(k_{1},l)}-\mathbb{E}[E_{n}^{(k_{1},l)}]\right)^{2}\mathbbm{1}[\Omega_{k_{1}}]\right]^{1/2}\mathbb{E}\left[\left(\sum_{l=1}^{2^{k_{2}}}E_{n}^{(k_{2},l)}-\mathbb{E}[E_{n}^{(k_{2},l)}]\right)^{2}\mathbbm{1}[\Omega_{k_{2}}]\right]^{1/2}$
$\leq(\log n)^{2}n^{1-\frac{1}{2(\sqrt{\log n})^{1+2\delta}}}+(\log n)^{2}\exp[\log n-\sqrt{\log n}^{1-\delta}+4\log\log n]\ll n.$
By using Markov’s inequality, we see that,
(4.8) $\mathbb{P}\left(\sum_{k=0}^{\log_{2}n_{\delta,-}}\sum_{l=1}^{2^{k}}(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]>\sqrt{n}\right)$
$\leq\mathbb{P}(\Omega^{c})+n^{-1}\mathbb{E}\left[\left(\sum_{k=0}^{\log_{2}n_{\delta,-}}\sum_{l=1}^{2^{k}}(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]\right)^{2}\mathbbm{1}[\Omega]\right]\to 0,$
as $n\to\infty$.
Recall that $\mathcal{G}$ from Part 1 was a Gaussian random variable and the order of the variance of $\mathcal{G}$ was $n\log n$, which is much larger than the error terms of Parts 2 and 3. Then, combining Part 1, 2, 3, we can obtain the desired result using the Lindeberg–Feller central limit theorem.
4.2.2. Effective resistance
Next, we will give the proof of Theorem 1.4 for $d=6$. We present only the necessary adjustments from the proof of Theorem 1.2. This involves changing Parts 1 and 4, while keeping Parts 2 and 3. We also remark that the results of this section only depend on the tail behavior of the cross terms $E_{n}$ behaving like $\mathbb{P}(E_{n}\geq l)\asymp\frac{1}{l}$. In particular, the results are general enough to not require any estimate on the correlation between the cross terms. Thus, the methods of this section are of general interest and can potentially be applied to other heavy tailed statistics of the random walk range.
We now define the function $f(n)$ as,
$\mathrm{Var}(X_{n})=n(\log n)^{f(n)}.$
Part 1: A CLT contribution
Define the value $n_{\delta}:=\exp\left[\frac{1}{2}\log n+(\sqrt{\log n})^{1-\delta}\right]$. Now,
$\mathcal{G}:=\sum_{i=1}^{n_{\delta}}X[(i-1)n/n_{\delta},in/n_{\delta}]-\mathbb{E}[X[(i-1)n/{n_{\delta}},in/n_{\delta}]]$
will be a normal random variable of variance $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$, provided $f(n/n_{\delta})\geq 0.$ Define
$Z_{n_{\delta},i}:=\frac{X[(i-1)n/n_{\delta},in/n_{\delta}]-\mathbb{E}[X[(i-1)n/n_{\delta},in/n_{\delta}]]}{\sqrt{n(\log n/n_{\delta})^{f(n/n_{\delta})}}}.$
The variables $Z_{n_{\delta},i}$ are all independent of each other and one
$\sum_{i=1}^{n_{\delta}}\mathbb{E}\left[Z_{n_{\delta},i}^{2}\right]=1.$
Furthermore, we have that,
$\sum_{i=1}^{n_{\delta}}\mathbb{E}[Z_{n_{\delta},i}^{2}\mathbbm{1}[|Z_{n_{\delta},i}|\geq\epsilon]]=0,$
since deterministically, we have that $X[(i-1)n/n_{n_{\delta}},in/n_{\delta}]\leq n/n_{\delta}\ll\sqrt{n(\log n/n_{\delta})^{f(n/n_{\delta})}}$, provided $f(n^{1/2-\delta})\geq 0$.
Part 4: Nontriviality of the Limit
We will consider those values of $n$ such that $f(n/n_{\delta})\geq 1-\delta/4$ (by Lemma 4.2, this should happen infinitely often.) We will divide the analysis into two cases.
Case 1: $f(n/n_{\delta})\geq 1-\delta/4$ and $f(n)\leq f(n/n_{\delta})+\frac{1}{\log\log n}$ happen infinitely often together.
Then, we have that
$\mathrm{Var}(X_{n})=n(\log n)^{f(n/n_{\delta})}(\log n)^{f(n)-f(n/n_{\delta})}\leq n(\log(n/n_{\delta})^{3})^{f(n/n_{\delta})}\log n^{\frac{1}{\log\log n}}$
[…]
Since $\mathcal{G}$ is Gaussian with variance $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$, with constant probability, it will be greater than $\sqrt{n(\log(n/n_{\delta}))^{f(n/n_{\delta})}}$. Equations (4.8) and (4.7) show that the remaining error terms will not be larger than $\epsilon\sqrt{n(\log n)^{1-\delta/2}}$ with high probability. Thus, on a set of non-vanishing probability, we must have that as $n\to\infty$,
$\mathbb{P}\bigg(\frac{\mathcal{M}+\mathcal{L}}{\sqrt{n(\log(n/n_{\delta}))^{f(n/n_{\delta})}}}\geq\epsilon\bigg)\to 0$
for any $\epsilon>0$. Thus, $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ converges to $\mathcal{N}(0,a)$.
Case 2: $f(n/n_{\delta})\geq 1-\delta/4$ and $f(n)\leq f(n/n_{\delta})+\frac{1}{\log\log n}$ do not happen infinitely often together.
Then there exists a sequence $m_{1},m_{2},\ldots,m_{k},\ldots$ such that $\frac{m_{k+1}}{(m_{k+1})_{\delta}}=m_{k},\forall k$ and $f(m_{k+1})-f(m_{k})\geq\frac{1}{\log\log m_{k+1}}$. (Notice that by Lemma 4.2, we will always be able to find elements such that $f(n/n_{\delta})\geq 1-\delta/4$, so we do not need to worry about the possibility that $f(n/n_{\delta})\leq 1-\delta/4$ for all large $n$.)
The relationship $m_{k}=\frac{m_{k+1}}{(m_{k+1})_{\delta}}$ will imply that $m_{k+1}\leq(m_{k})^{3}$. Thus, $m_{k+1}\leq(m_{1})^{3^{k}}$. Therefore, we have that
$f(m_{k+1})-f(m_{k})\geq\frac{1}{\log\log m_{k+1}}\geq\frac{1}{(k+1)\log 4},$
for all $k$ large enough. This will imply that $\lim_{k\to\infty}f(m_{k})=\infty$. This contradicts the upper bound on $f$ from Lemma 4.1.
Thus, we see that this proof establishes that there is a subsequence where $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}},$ does not converge to 0. Moreover, along this subsequence $X_{n}-\mathbb{E}[X_{n}]$ will be distributed (after appropriate normalization) as a normal random variable whose variance is given by $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$. ∎
5. The fluctuations converge to 0 in dimension $d=5$
In this section, we will give the proof of Propositions 1.1 and 1.3 and Theorems 1.2 and 1.4 in $d=5$. Since we do not show the correlation estimates (5.1) for the effective resistance, we present the weaker estimates of its variance. Since the proof is similar, we will omit the proof of Proposition 1.3 in $d=5$.
Lemma 5.1.
We have
$\mathrm{Var}(X_{n})\asymp n^{3/2}.$
Proof.
Fix some small $\alpha>0$. We consider the variance of $\mathcal{E}=-\sum_{k=0}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$. For $0\leq k_{1},k_{2}\leq(1-\alpha)\log_{2}n$, $1\leq l_{1}\leq 2^{k_{1}}$, $1\leq l_{2}\leq 2^{k_{2}}$, let
$\Lambda_{1}:=\{((k_{1},l_{1}),(k_{2},l_{2})):[2^{-k_{1}}(l_{1}-1)n,2^{-k_{1}}l_{1}n]\cap[2^{-k_{2}}l_{2}n,2^{-k_{2}}(l_{2}+1)n]\neq\emptyset\},$
$\Lambda_{2}:=\{((k_{1},l_{1}),(k_{2},l_{2})):[2^{-k_{1})}(l_{1}-1)n,2^{k_{1}}l_{1}n]\cap[2^{-k_{2}}(l_{2}-1)n,2^{-k_{2}}l_{2}n]=\emptyset\}.$
Since
$\sum_{((k_{1},l_{1}),(k_{2},l_{2}))\in\Lambda_{2}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]=\sum_{((k_{1},l_{1}),(k_{2},l_{2}))\in\Lambda_{2}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}],$
we have
$\mathrm{Var}(\mathcal{E})\geq\sum_{k=0}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]-\sum_{((k_{1},l_{1}),(k_{2},l_{2}))\in\Lambda_{1}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}].$
By (2.2), we have $\sum_{k=0}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\geq cn^{3/2}$. Now we claim that
$\sum_{((k_{1},l_{1}),(k_{2},l_{2}))\in\Lambda_{1}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\leq n^{3/2-\alpha/2+o(1)}.$
Indeed,
$\sum_{((k_{1},l_{1}),(k_{2},l_{2}))\in\Lambda_{1}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$
$\lesssim$ $(\log n)^{2}\sup_{\alpha\leq\beta_{1}\leq\beta_{2}\leq 1}\sum_{l_{1},l_{2}:((1-\beta_{1})\log_{2}n,l_{1})((1-\beta_{2})\log_{2}n,l_{2})\in\Lambda_{1}}\mathbb{E}[E_{n}^{((1-\beta_{1})\log_{2}n,l_{1})}]\mathbb{E}[E_{n}^{{(1-\beta_{2})}\log_{2}n,l_{2})}]$
$\lesssim$ $(\log n)^{2}\sup_{\alpha\leq\beta_{1}\leq\beta_{2}\leq 1}n^{\beta_{2}-\beta_{1}}n^{1-\beta_{2}}n^{\beta_{2}/2}n^{\beta_{1}/2}\leq n^{3/2-\alpha/2+o(1)}.$
Hence, we obtain $\mathrm{Var}(\mathcal{E})\gtrsim n^{3/2}$.
In addition, (2.2) yields $\sum_{k=1}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\leq cn^{3/2}$. Now by the same argument as that of Lemma 4.1, we have
(5.1) $\sum_{k_{1}=0}^{(1-\alpha)\log_{2}n}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{(1-\alpha)\log_{2}n}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\leq n^{3/2-\alpha/2+o(1)}.$
Hence, we obtain $\mathrm{Var}(\mathcal{E})\lesssim n^{3/2}$.
If we let $MT=\sum_{l=1}^{n^{1-\alpha}}X[(l-1)n^{\alpha},ln^{\alpha}]$, $\mathrm{Var}(MT)\leq\sum_{l=1}^{n^{1-\alpha}}\mathbb{E}[X[(l-1)n^{\alpha},ln^{\alpha}]^{2}]\lesssim n^{1+\alpha}$ for sufficiently small $\alpha>0$. Then, $\mathrm{Var}(X_{n})\geq(\mathrm{Var}(\mathcal{E})^{1/2}-\mathrm{Var}(MT)^{1/2})^{2}\geq cn^{3/2}$ and $\mathrm{Var}(X_{n})\leq(\mathrm{Var}(\mathcal{E})^{1/2}+\mathrm{Var}(MT)^{1/2})^{2}\lesssim n^{3/2}$. ∎
Proof of Theorems 1.2 and 1.4 in $d=5$.
For small $\delta>0$, we use Markov’s inequality in the computation of the variance, to derive,
$\sup_{\alpha\leq\beta\leq 1-\delta}\mathbb{P}\bigg(\frac{\sum_{l=1}^{n^{1-\beta}}(E_{n}^{((1-\beta)\log n,l)}-\mathbb{E}[E_{n}^{((1-\beta)\log n,l)}])}{\sqrt{\mathrm{Var}(X_{n})}}\geq\frac{\epsilon}{\log n}\bigg)$
$\leq$ $\sup_{\alpha\leq\beta\leq 1-\delta}(\log n)^{2}\epsilon^{-2}\frac{\sum_{l=1}^{n^{1-\beta}}\mathrm{Var}(E_{n}^{((1-\beta)\log n,l)})}{\mathrm{Var}(X_{n})}$
$\lesssim$ $\sup_{\alpha\leq\beta\leq 1-\delta}(\log n)^{2}\frac{n^{1-\beta}n^{3\beta/2}}{n^{3/2}}\lesssim(\log n)^{2}n^{-\delta/2}.$
Now, we use Markov’s inequality in the computation of the absolute value combined with the triangle inequality, to get,
$\sup_{1-\delta\leq\beta\leq 1}\mathbb{P}\bigg(\frac{\sum_{l=1}^{n^{1-\beta}}(E_{n}^{((1-\beta)\log n,l)}-\mathbb{E}[E_{n}^{((1-\beta)\log n,l)}])}{\sqrt{\mathrm{Var}(X_{n})}}\geq\frac{\epsilon}{\log n}\bigg)$
$\leq$ $\sup_{1-\delta\leq\beta\leq 1}(\log n)\epsilon^{-1}\frac{\sum_{l=1}^{n^{1-\beta}}\mathbb{E}[E_{n}^{((1-\beta)\log n,l)}]}{\sqrt{\mathrm{Var}(X_{n})}}$
$\lesssim$ $\sup_{1-\delta\leq\beta\leq 1}(\log n)\frac{n^{1-\beta}n^{\beta/2}}{n^{3/4}}\lesssim(\log n)n^{-1/4+\delta/2}.$
Therefore,
$\mathbb{P}\bigg(\frac{|X_{n}-\mathbb{E}[X_{n}]|}{\sqrt{\mathrm{Var}(X_{n})}}\geq\epsilon\bigg)$
$\leq$ $\mathbb{P}\bigg(\frac{|MT-\mathbb{E}[MT]|}{\sqrt{\mathrm{Var}(X_{n})}}\geq\epsilon/2\bigg)$
$+C(\log n)\sup_{\alpha\leq\beta\leq 1}\mathbb{P}\bigg(\frac{\sum_{l=1}^{n^{1-\beta}}(E_{n}^{(\beta\log n,l)}-\mathbb{E}[E_{n}^{(\beta\log n,l)}])}{\sqrt{\mathrm{Var}(X_{n})}}\geq\epsilon/(2\log n)\bigg)\lesssim n^{-\alpha/4}$
as desired. ∎
6. The fluctuations converge to 0 in dimension $d=4$
In this section, we estimate the variance and the Limit in distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ in $d=4$.
In what follows, we will simplify notation by reducing to a discussion of only the cross term $E_{n}$ since we will only apply a dyadic decomposition and the only other cross term that could appear, $X_{2n+1}-X_{n}-X[n,2n+1]$, will have a similar analysis.
In subsection 6.1, we will give the lower bounds for the cross term, which holds for the graph distance, cut points and the effective resistance. In subsections 6.2-6.4, we will show the upper bounds for the cross term and main results in $d=4$. In these subsections, we are considering only the graph distance, cut points since Lemma 6.8 does not work for the effective resistance. We will give our main result in $d=4$ for the effective resistance in subsection 6.5.
6.1. Lower bounds for the Cross Term
We will first start by obtaining a lower bound of $\mathbb{E}[E_{n}].$ Our heuristic argument is the following: if there is an intersection between $S[3n/2,2n]$ and $S[0,n]$, then this will form a shortcut in the graph distance between the points $S_{2n}$ and $S_{0}$. The amount that the graph distance would reduce via this shortcut will be related to the graph distance between $S_{n}$ and $S_{2n}$, which by a law of large numbers should be $O\left(\frac{n}{(\log n)^{1/2}}\right)$. Furthermore, the probability of the intersection of $S[3n/2,2n]$ and $S[0,n]$ is approximately $O((\log n)^{-1})$. Thus, one would heuristically argue that $\mathbb{E}[E_{n}]=\frac{n}{(\log n)^{3/2\pm o(1)}}$.
However, the exact value of the reduction of the graph distance is a complicated function of the loop intersection structure of the two walks. Thus, knowing that there is a long-range intersection may not be enough the deduce the reduction of the graph distance. Furthermore, to get the exact order $\frac{n}{(\log n)^{3/2}}$, that we desire, then we would need to argue that there is exact independence between the event that there is an intersection between $S[3n/2,2n]$ and $S[0,n]$ and the event that the graph distance between $S_{n}$ and $S_{2n}$ is $O\left(\frac{n}{(\log n)^{1/2}}\right)$. The following lemma will address these issues by finding appropriate events that will allow us to constrain the geometry of the random walk range.
Lemma 6.1.
Let $E_{n}$ denote the cross term that appears when considering the number of cut points, the effective resistance or the graph distance. We have $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$.
Remark 6.2.
In [21, Section 4], he has already shown $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$ for the effective resistance. We obtain the same result by a slightly simpler proof which derives intersection estimates by using the capacity of the random walk.
Proof.
We split the walk $S[0,2n]$ into two parts. $\tilde{S}$ will be the reversal of the first part, that is, $\tilde{S}_{i}=S_{n-i}-S_{n}$ . $\hat{S}$ will be the second part, that is $\hat{S}_{i}=S_{n+i}-S_{n}$.
We now define the following events that impose good properties on the first walk $\tilde{S}_{n}$. Pick small $\epsilon>0$. We let $\tilde{T}_{i}$ be the sequence of cut points along the walk $\tilde{S}_{n}$, and
$A_{1}=\{\max_{\tilde{T}_{i}\leq n}(\tilde{T}_{i+1}-\tilde{T}_{i})\leq\frac{n}{(\log n)^{7/12}}\},\quad A_{2}=\{\tilde{X}_{n/2}\geq\frac{1}{1+\epsilon}\mathbb{E}[X_{n/2}]\},$
$A_{3}=\{\tilde{X}[n-n/(\log n)^{\epsilon},n]\leq n/(\log n)^{1/2+\epsilon/2}\},$
$A_{4}=\{{\rm Cap}(\tilde{S}(n-n/(\log n)^{\epsilon},n))\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}\},$
$A_{5}=\{\tilde{S}_{i}\leq C_{1}\sqrt{n},\quad\forall i\in[n-n/(\log n)^{\epsilon},n]\}$
noting $\mathbb{E}[{\rm Cap}(\tilde{S}[0,n])]\sim\frac{\pi^{2}}{8}\frac{n}{\log n}$ (e.g., see [6]). In the above equation, we use the notation $\tilde{X}[a,b]$ to denote the graph distance between the points $\tilde{S}_{a}$ and $\tilde{S}_{b}$ along the graph $\tilde{S}[a,b]$. By [11, Lemma 2.3], $\mathbb{P}(A_{1})=1-o(1)$. Similarly, the law of large numbers ensures that $\mathbb{P}(A_{2})$, $\mathbb{P}(A_{3})=1-o(1)$ noting [11, Theorem 1.2] and $\mathbb{P}(A_{4})=1-o(1)$ (see [6, Theorem 1.1]). By the local central limit theorem, we have that $\mathbb{P}(A_{5})=1-\epsilon(C_{1})$, where $\epsilon(C_{1})\to 0$ as $C_{1}\to\infty$. Let $\mathcal{Q}:=A_{1}\cap A_{2}\cap A_{3}\cap A_{4}\cap A_{5}$. From our previous estimates, we can take a union bound to deduce that $\mathbb{P}(\mathcal{Q})\geq 1-2\epsilon(C_{1})$, for $n$ sufficiently large.
Now, we need to define an event on the intersection of $\hat{S}$ with $\tilde{S}$. Let $\tau$ be defined to be the first time of intersection between $\hat{S}$ and $\tilde{S}[n-n/(\log n)^{\epsilon},n]$ after time $n-n/(\log n)^{\epsilon}$. That is, $\tau:=\inf\{i>n-n/(\log n)^{\epsilon}:\hat{S}_{i}\in\tilde{S}[n-n/(\log n)^{\epsilon},n]\}$. We define,
$B_{1}=\{\hat{X}[\tau,n]\leq n/(\log n)^{1/2+\epsilon/2}\},\quad B_{2}=\{\hat{S}_{n-n/(\log n)^{\epsilon}}\leq C_{1}\sqrt{n}\},$
$B_{3}=\{\hat{S}[n-n/(\log n)^{\epsilon},n]\cap\tilde{S}[n-n/(\log n)^{\epsilon},n]\neq\emptyset\}.$
As before, we let $\hat{X}[a,b]$ to denote the graph distance between the points $\hat{S}_{a}$ and $\hat{S}_{b}$ on the graph connecting $\hat{S}[a,b]$. Note that event $B_{1}$ is independent of $B_{2}\cap B_{3}$ by the Strong Markov Property, even if you have conditioned on the walk $\tilde{S}[0,n]$. By the law of large numbers, we have that $\mathbb{P}(B_{1})=1-o(1)$ and
(6.1) $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q})$ $=\mathbb{P}(B_{1}|B_{2}\cap B_{3}\cap\mathcal{Q})\mathbb{P}(B_{2}\cap B_{3}\cap\mathcal{Q})$
$=\mathbb{P}(B_{1})\mathbb{P}(B_{3}|B_{2}\cap\mathcal{Q})\mathbb{P}(B_{2}\cap\mathcal{Q}).$
We first claim that we have $\mathbb{P}(B_{3}|B_{2}\cap\mathcal{Q})\geq\frac{c}{(\log n)^{1+2\epsilon}}$ for some constant $c>0$. We use $\mathcal{S}$ as a shorthand for the set $\tilde{S}[n-n/(\log n)^{\epsilon},n]$. Let $\hat{S}_{n-n/(\log n)^{\epsilon}}\leq C_{1}n^{1/2}$, as is necessary to satisfy the event $B_{2}$. By the last hitting time decomposition, we have that on $\mathcal{Q}$
$\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\tau<n/(\log n)^{\epsilon})$
$=$ $\sum_{i=1}^{n/(\log n)^{\epsilon}}\sum_{a\in\mathcal{S}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=a)\mathbb{P}^{a}(\tau>n/(\log n)^{\epsilon}-i)$
$\geq$ $\sum_{i=1}^{n/(\log n)^{\epsilon}}\sum_{a\in\mathcal{S}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=a)\mathbb{P}^{a}(\tau=\infty)$
$\geq$ $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right)\sum_{a\in\mathcal{S}}\mathbb{P}^{a}(\tau=\infty)$
$=$ $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right){\rm Cap}(\mathcal{S}).$
On the union $A_{5}\cap B_{2}$, we can apply the local central limit theorem to deduce that $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right)\geq cn^{-1}(\log n)^{-\epsilon}$, for some constant $c>0$ on $\mathcal{Q}$. Furthermore, the event $A_{5}$ ensures that ${\rm Cap}(\mathcal{S})\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}$ on $\mathcal{Q}$. Multiplying these together, we see that there is some $c>0$ such that $\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\tau<n/(\log n)^{\epsilon})>c/(\log n)^{1+2\epsilon}$ on $\mathcal{Q}$. This means that $\mathbb{P}(B_{3}|B_{2}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$. Returning to equation (6.1), we see that we see that we must have $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$.
Now, on the event $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$, we claim we have $E_{n}>n/(\log n)^{1/2+\epsilon/2}$. Due to the presence of the cut points from the event $A_{1}$ and the lower bound on the graph distance $\tilde{X}_{n/2}$ from $A_{2}$, we have that $\tilde{X}_{n}\geq\frac{1}{1+\epsilon}\mathbb{E}[X_{n/2}]-\frac{n}{(\log n)^{7/12}}$. By events $A_{3}$, $A_{1}$, and $B_{1}$, the graph distance between the points $\hat{S}_{n}$ and $\tilde{S}_{n}$, is at most $2n/(\log n)^{1/2+\epsilon/2}+2\frac{n}{(\log n)^{7/12}}$. Computing the difference, we now see that $E_{n}>n/(\log n)^{1/2+\epsilon/2}$ on $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$ as desired and this event happens with probability $c/(\log n)^{1+2\epsilon}$. By multiplying these two terms together, we see that $\mathbb{E}[E_{n}]>n/(\log n)^{3/2+\epsilon/2+2\epsilon}$. This proof also works even for the effective resistance or the number of cut points. ∎
Lemma 6.3.
There exists a set $\Omega$ with probability $\text{O}\left(\frac{\log\log n}{\log n}\right)$ such that we have
$E_{n}\mathbbm{1}[\Omega^{c}]\ll\mathbb{E}[E_{n}].$
Proof.
Recall the notation $\hat{S}$ and $\tilde{S}$ from the proof of Lemma 6.1. Consider the event
$\Omega$= $\{\hat{S}[n/(\log n)^{4},n]\cap\tilde{S}[0,n]\neq\emptyset\}\bigcup\{\tilde{S}[n/(\log n)^{4},n]\cap\hat{S}[0,n]\neq\emptyset\}$
$\bigcup\left\{\sum_{i\in\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]}1\{\tilde{S}[0,i-1]\cap\tilde{S}[i,n]=\emptyset\}=0\right\}$
$\bigcup\left\{\sum_{i\in\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]}1\{\hat{S}[0,i-1]\cap\hat{S}[i,n]=\emptyset\}=0\right\}.$
The first two events that compose $\Omega$ ensure that there exists some long-range intersection between $\tilde{S}$ and $\hat{S}$. The third and fourth events force there to be no cut points in $\tilde{S}$ between $n/(\log n)^{4}$ and $n/(\log n)^{2}$.
We claim both that $\Omega$ holds with probability $O\left(\frac{\log\log n}{\log n}\right)$ and that on the event $\Omega^{c}$, we must have that the cross term $E_{n}$ for number of cut points, the effective resistance, and the graph distance must by less than $\frac{n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$ by Lemma 6.1.
By [15, Theorem 4.3.6 part(ii)], we must have that $\mathbb{P}(\hat{S}[n/(\log n)^{4},n]\cap\tilde{S}[0,n]\neq\emptyset)=\text{O}\left(\frac{\log\log n}{\log n}\right)$. Similarly, by [11, equation (7)],
$\mathbb{P}\left(\sum_{i\in\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]}1\{\tilde{S}[0,i-1]\cap\tilde{S}[i,n]=\emptyset\}\right)=\text{O}\left(\frac{\log\log n}{\log n}\right).$
Combining these bounds gives us that $\mathbb{P}(\Omega)\leq\text{O}\left(\frac{\log\log n}{\log n}\right)$, as desired.
Now, we will check that on $\Omega^{c}$, our cross term can be no more than $\frac{n}{(\log n)^{2}}$. Notice that on $\Omega^{c}$, there must be some cut point in the interval $\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]$ for both $\tilde{S}$ and $\hat{S}$. Let $\mathcal{C}_{\tilde{S}}$ denote the time instance of the cut point of $\tilde{S}$ (cut time) in this interval and let $\mathcal{C}_{\hat{S}}$ denote the time instance of the cut point of $\hat{S}$ in this interval. Furthermore, since there is no intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$, $\mathcal{C}_{\hat{S}}$ must still be a cut point in the full graph $\hat{S}\cup\tilde{S}$.
Notice that we have $\hat{X}_{n}=\hat{X}[0,\mathcal{C}_{\hat{S}}]+\hat{X}[\mathcal{C}_{\hat{S}},n]\leq\hat{X}[\mathcal{C}_{\hat{S}},n]+n/(\log n)^{2}.$ Similarly, we have $\tilde{X}_{n}\leq\tilde{X}[\mathcal{C}_{\tilde{S}},n]+n/(\log n)^{2}.$ Lastly, we see that $X_{2n}\geq\hat{X}[\mathcal{C}_{\hat{S}},n]+\tilde{X}[\mathcal{C}_{\tilde{S}},n]$. Altogether, we see that on $\Omega^{c}$, we have $E_{n}=\tilde{X}_{n}+\hat{X}_{n}-X_{2n}\leq\frac{2n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$. ∎
Remark 6.4.
If one is interested in the graph distance or the number of cut points, then we can proceed to more simply define $\Omega$ by considering only its first line from its definition in Lemma 6.3. We will use this fact later in the proof of Lemma 6.8.
By applying the Cauchy-Schwarz inequality, we obtain the following lower bound on $\mathbb{E}[E_{n}^{2}]$ as a corollary.
Corollary 6.5.
Let $E_{n}$ denote the cross term that appears when computing the number of cut points, the effective resistance, or the graph distance. Then, we have,
$\mathbb{E}[E_{n}^{2}]\geq\frac{n^{2}}{(\log n)^{2+o(1)}}.$
Proof.
Let $\Omega$ be the set from Lemma 6.3. Note that by the Cauchy-Schwarz inequality, we have that,
$\mathbb{E}[E_{n}^{2}]\mathbb{P}(\Omega)\geq\mathbb{E}[E_{n}\mathbbm{1}[\Omega]]^{2}\geq(1-o(1))\mathbb{E}[E_{n}]^{2}.$
Since $\mathbb{P}(\Omega)=O\left(\frac{\log\log n}{\log n}\right)$, we can take ratios and deduce that $\mathbb{E}[E_{n}^{2}]\geq\frac{n^{2}}{(\log n)^{2+o(1)}}$, as desired. ∎
6.2. Upper bounds for the Cross terms
Our proof of the upper bound for $\mathbb{E}[E_{n}^{2}]$ requires more steps. Indeed, unlike our proof of the lower bound, we cannot establish our upper bound directly. However, we will first show that if we have the inequality $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$, this will be enough to show that $\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2-o(1)}}$ and ultimately that $\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{2-o(1)}}$.
6.2.1. Our first reduction
The goal of this section is to establish our reduction that showing $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$ is enough to establish $\mathbb{E}[E_{n}^{2}],\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$.
Lemma 6.6.
Let $X_{n}$ denote the number of cut points or the graph distance. Assume that there is some $\epsilon>0$ such that for all $n$, we have
(6.2) $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n}).$
Then, we must have that,
$\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2-o(1)}},\quad\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{2-o(1)}}.$
Proof.
Fix any $\kappa>0$, we will first show that for $n$ large enough, we must have that,
$\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2-\kappa}}.$
Let us write
$\mathrm{Var}(X_{n})=\frac{n^{2}}{(\log n)^{f(n)}}.$
We will use the dyadic decomposition to find an equation for $f(n)$. We see that,
$\mathrm{Var}(X_{2n})=\mathrm{Var}(X_{n})+\mathrm{Var}(X_{[n,2n]})+\mathrm{Var}(\mathcal{E}_{n})+\text{Cov}(\mathcal{E}_{n},X_{n}+X_{[n,2n]}).$
We then have that using the previous expansion on the variance and equation (6.2) that,
$\frac{(2n)^{2}}{(\log(2n))^{f(2n)}}$ $\leq\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\frac{n^{2}}{(\log n)^{f(n)}}$
$+\sqrt{\frac{(\log n)^{-\epsilon}n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}}\sqrt{\frac{2n^{2}}{(\log n)^{f(n)}}}.$
We can bound the right-hand side as,
$\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+\frac{n^{2}}{(\log n)^{f(n)+\epsilon}}+C\frac{n^{2}}{(\log n)^{f(n)+\epsilon/2}}+C\frac{n^{2}\sqrt{\log\log n}}{(\log n)^{f(n)/2+1}}$
$\leq\frac{(2+\delta)n^{2}}{(\log n)^{f(n)}}+C\frac{n^{2}\log\log n}{(\log n)^{2}},$
where $\delta>0$ can be set to be any small constant and $C$ is some universal constant (that could depend on $\delta$) and could change from line to line.
We consider two cases:
Case 1: $f(n)\geq 2-\kappa$
In this case, note that for $n$ large enough, we must have that $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{2-\kappa}}$. Thus, we obtain the inequality,
(6.3) $\frac{4n^{2}}{(\log 2n)^{f(2n)}}\leq\frac{(2+2\delta)n^{2}}{(\log n)^{2-\kappa}},$
$\log\frac{4}{2+2\delta}+(2-\kappa)\log\log n\leq f(2n)\log\log 2n,$
$f(2n)\geq(2-\kappa)\frac{\log\log n}{\log\log n+\log\left[1+\frac{\log 2}{\log n}\right]}+\frac{1}{\log\log 2n}\log\frac{4}{2+2\delta}$
$\geq(2-\kappa)-\frac{2-\kappa}{2\log n\log\log n}+\frac{1}{\log\log 2n}\log\frac{4}{2+2\delta}\geq(2-\kappa),$
when $n$ is large enough.
Case 2: $f(n)\leq 2-\kappa$
In this case, when $n$ is large enough, we can still obtain the inequality for any small $\delta>0$, $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{f(n)}}$. Thus, we obtain the inequality,
(6.4) $\frac{4n^{2}}{(\log 2n)^{f(2n)}}\leq\frac{(2+2\delta)n^{2}}{(\log n)^{f(n)}},$
$(\log 2n)^{f(2n)}\geq\frac{4}{2+2\delta}(\log n)^{f(n)},$
$f(2n)\log\log 2n\geq\log\frac{4}{2+2\delta}+f(n)\log\log n,$
$f(2n)\geq f(n)\frac{\log\log n}{\log\log 2n}+\frac{1}{\log\log 2n}\log\frac{4}{2+2\delta}$
$\geq f(n)-\frac{f(n)}{2\log n\log\log n}+\frac{1}{\log\log(2n)}\log\frac{4}{2+2\delta}\geq f(n)+\frac{c}{\log\log n},$
where $c>0$ will be some universal constant, and the least line will hold provided $n$ is sufficiently large.
Now, let $n_{0}$ be large enough integer such that both equations (6.3) and (6.4) hold. Furthermore, consider $k$ large enough so that
$\sum_{j=0}^{k-1}\frac{c}{\log\log(2^{j+1}n_{0})}\geq 2-\kappa.$
We claim that for any $n\geq 2^{k}n_{0}$, we must have that $f(n)\geq 2-\kappa$. First, consider the sequence $f(n),f(n/2),f(n/4),\ldots,f(n/2^{L})$, where $L$ is the largest integer such that $n/2^{L}\geq n_{0}$. If there existed some $0\leq i\leq L$ such that $f(n/2^{i})\geq 2-\kappa$, then by repeatedly applying equation (6.3), we must have that $f(n)\geq 2-\kappa$.
Otherwise, (6.4) must hold for all $0\leq i\leq L$, and we see that $f(n)\geq f(n/2^{L})+\sum_{j=0}^{L}\frac{c}{\log\log(n/2^{j})}\geq\sum_{j=0}^{k-1}\frac{c}{\log\log 2^{j+1}n_{0}}\geq 2-\kappa$. Here, we remark that we used the fact that $f(n)\geq 0$ is a trivial lower bound for all $n$.
By substituting our bound on $\mathrm{Var}(X_{n})$ into equation (6.2), we derive an upper bound on $\mathbb{E}[E_{n}^{2}]$. ∎
Assuming the consequence of the above result, we also have the following corollary.
Corollary 6.7.
Assume that we know $\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$, then we have $\mathbb{E}[E_{n}]\leq\frac{n}{(\log n)^{3/2+o(1)}}$.
Proof.
Recall the set $\Omega$ from Lemma 6.3. By the Cauchy-Schwarz inequality and the definition of $\Omega$, we have
$\mathbb{E}[E_{n}^{2}]\mathbb{P}(\Omega)\geq\mathbb{E}[E_{n}\mathbbm{1}[\Omega]]^{2}\geq(1-o(1))\mathbb{E}[E_{n}]^{2}.$
Using the fact that $\mathbb{P}(\Omega)=O(\log\log n/\log n)$ gives the result. ∎
6.2.2. Establishing equation (6.2)
The goal of this section is to establish equation (6.2). We will first show that this equation holds provided that one derives an appropriate conditional statement for the probability that two random walks intersect. Before we start, let us recall the notation that we used in the proof of Lemma 6.1. Let us call the reversal of the walk $S[0,n]$ from time 0 to $n$ $\hat{S}$. That is $\hat{S}_{i}=S_{n}-S_{n-i}$. Let us call $S[n,2n]$ as $\tilde{S}$. That is $\tilde{S}_{i}=S_{n+i}-S_{n}$.
In this subsection, we will give the estimate only for cut points and the graph distance.
Lemma 6.8.
Let $X_{n}$ denote the number of cut points, or the graph distance. Assume that there exists some $\epsilon>0$ such that with probability significantly less than $(\log n)^{-2}$, we have that,
(6.5) $\mathbb{P}(\tilde{S}[n/(\log n)^{4},n]\cap\hat{S}\neq\emptyset|\hat{S})\geq(\log n)^{-\epsilon}.$
Then, equation (6.2) holds.
Proof.
  1. (1)
    Notice that we have $E_{n}>l$ implies that $\tilde{X}_{n}\geq\frac{l}{2}$ or $\hat{X}_{n}\geq\frac{l}{2}$. Furthermore, $\Omega$ also holds by Lemma 6.3. Then,
    $\{E_{n}>l\}\subset(\{\tilde{X}_{n}\geq l/2\}\cap\Omega)\cup(\{\hat{X}_{n}\geq l/2\}\cap\Omega).$
  2. (2)
    Furthermore, (again still considering that $E_{n}>l$) in the case $\tilde{X}_{n}\geq\frac{l}{2}$, it must be the case that there is an intersection between $\tilde{S}{[n/(\log n)^{4},n]}$ and $\hat{S}$ by the proof of Lemma 6.3. (Recall, we needed the event $\Omega$ to hold in order to ensure that there is a large reduction in the graph distance.)
  3. (3)
    Vice versa, in the case $\hat{X}_{n}\geq\frac{l}{2}$, it must be the case that there is an intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$.
Thus, we have that,
[…]
Furthermore, observe that,
[…]
We used $\mathbb{E}[\tilde{X}_{n}]\leq\frac{n}{(\log n)^{1/2+o(1)}}$ to show the final inequality. (Note that in [11, Lemma 2.2], they stated that it should be obtained for the graph distance if one follows the proof of [21].) Now, assume that equation (6.5) holds with probability significantly less than $(\log n)^{-2}$. Let $\Omega_{B}$ be the set of walks $\tilde{S}$ such that equation (6.5) holds. We have that,
$\sum_{l>4\mathbb{E}[\tilde{X}_{n}]}l\mathbb{P}(\tilde{X}_{n}\geq l/2,\tilde{S}[n/(\log n)^{4},n]\cap\hat{S}\neq\emptyset,\tilde{S}\not\in\Omega_{B})$
$+\sum_{l>4\mathbb{E}[\tilde{X}_{n}]}l\mathbb{P}(\tilde{X}_{n}\geq l/2,\tilde{S}[n/(\log n)^{4},n]\cap\hat{S}\neq\emptyset,\tilde{S}\in\Omega_{B})$
$\leq n\sum_{l>4\mathbb{E}[\tilde{X}_{n}]}\mathbb{P}(\tilde{S}\not\in\Omega_{B})$
$+\sum_{l>4\mathbb{E}[\tilde{X}_{n}]}l\mathbb{P}(\hat{S}\cap\tilde{S}[n/(\log n)^{4},n]\neq\emptyset|\tilde{S}\in\Omega_{B},\tilde{X}_{n}>l/2)\mathbb{P}(\tilde{S}\in\Omega_{B},\tilde{X}_{n}\geq l/2)$
$\leq o(n^{2}(\log n)^{-2})+\sum_{l>4\mathbb{E}[\tilde{X}_{n}]}l(\log n)^{-\epsilon}\mathbb{P}(\tilde{X}_{n}\geq l/2)\leq o(n^{2}(\log n)^{-2})+2\mathrm{Var}(X_{n})(\log n)^{-\epsilon}.$
Combining the previous equation with equation (6.6) gives the desired result. ∎
6.3. Establishing equation (6.5)
Lemma 6.9.
With probability greater than $1-\frac{1}{(\log n)^{2}}$ on walks $\tilde{S}$, there exists some $c>0$ such that we have,
$\mathbb{P}(\hat{S}\cap\tilde{S}[\frac{n}{(\log n)^{4}},n]\neq\emptyset|\tilde{S})\leq(\log n)^{-c}.$
We start with the following observation. We define a sequence of scales $0=K_{0}<K_{1}<K_{2}<K_{3}<\ldots<K_{f}$ and consider the following bound. Also, let $A(R_{o},R_{i})$ denote the annulus around the origin with outer radius $R_{o}$ and inner radius $R_{i}$. Let $\mathcal{A}_{i,n}:=A[\sqrt{n/K_{i}},\sqrt{n/K_{i+1}}]$.
By considering the last time of intersection between $\hat{S}$ and the set $\tilde{S}[\frac{n}{(\log n)^{4}},n]$, we can derive the following bound on the probability of intersection.
[…]
(6.7) $\leq$ $\sum_{i=0}^{10}\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\max_{z\in\mathcal{A}_{i,n}}G(z)$
(6.8) $+$ $\sum_{i=11}^{f-1}\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\max_{z\in\mathcal{A}_{i,n}}G(z).$
Our heuristics for (6.8) are the following:
  1. (1)
    The capacity of $\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})$ can be bounded by $|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|$. Namely, the capacity will be bounded by the volume.
  2. (2)
    The volume can be bounded very closely by $\frac{n}{K_{i}^{1+\delta}}$, which is related to the expected local time of the random walk $\tilde{S}$ in the annulus $\mathcal{A}_{i,n}$.
  3. (3)
    For (6.7), we will use improved estimates for the capacity of random walks.
Eventually, we will choose $K_{i}=\frac{1}{(\log n)^{i(1-\epsilon)}}$ for $i\leq 11$, and after this we choose $K_{i}=2K_{i-1}$ for $i\geq 12$.
Furthermore, we can choose $K_{f}=n^{1/10}$. Note that by the local central limit theorem, the probability a point $\tilde{S}_{k}$ for $k\in[n/(\log n)^{4},n]$ lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{n^{2/10}}{\frac{n^{2}}{(\log n)^{8}}}$. Taking a union bound, the probability that there is a point in $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ that lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{nn^{1/5}(\log n)^{8}}{n^{2}}\ll(\log n)^{-2}$.
6.3.1. The large scales
We will first begin our analysis of equation (6.8) by considering the largest scales after $K_{11}$.
We need the following lemma.
Lemma 6.10.
Let $S[0,n]$ be a random walk starting at the origin in $d=4$. Let $\mathcal{B}_{r}$ be a ball of radius $r$ centered at the origin. Let $L(\mathcal{B}_{r})$ be defined as,
$L(\mathcal{B}_{r}):=\sum_{t=\frac{n}{(\log n)^{4}}}^{n}\mathbbm{1}[S_{t}\in\mathcal{B}_{r}].$
Namely, it is the amount of times that $S_{t}$ lies in the ball $\mathcal{B}_{r}$.
Then, we have that
(6.9) $\mathbb{E}[L(\mathcal{B}_{r})]\leq C\frac{r^{4}(\log n)^{4}}{n}.$
Furthermore, we would also have,
(6.10) $\mathbb{E}[L(\mathcal{B}_{r})^{k}]\leq C^{k}\frac{r^{4}(r^{2})^{k-1}(\log n)^{4}}{n}.$
Proof.
By using the local central limit theorem, one observes that the probability of a point $S_{t}$ being in the ball $\mathcal{B}_{r}$ is of order $\frac{r^{4}}{t^{2}}$. If we sum up this probability over all times $t$ from $n/(\log n)^{4}$ and $n$, we see that the expectation of the local time within this ball is $\frac{r^{4}(\log n)^{4}}{n}$.
Now, if we want to consider the expectation of the $k$th power of the local time. We need to consider the probability that there are k distinct times $t_{1}<t_{2}<\ldots<t_{k}$ such that $S_{t_{i}}$ are all in the ball with radius $r$. First, we remark that given that $S_{t_{1}}$ is in $\mathcal{B}_{r}$, the probability that $S_{t_{2}}\in\mathcal{B}_{r}$ is at most $\text{O}\left(\frac{r^{4}}{(t_{2}-t_{1})^{2}}\right)$ if $t_{2}-t_{1}\geq r^{2}$ or $1$ if $t_{2}-t_{1}\leq r^{2}$. Thus, we see a bound on $\mathbb{E}[L(\mathcal{B}_{r})^{k}]$ would be the following.
$\sum_{t_{1}=n/(\log n)^{4}}^{n}\sum_{t_{2}-t_{1}=1}^{n}\ldots\sum_{t_{k}-t_{k-1}=1}^{n}\frac{Cr^{4}(\log n)^{4}}{n}\prod_{i=1}^{k-1}\frac{Cr^{4}}{\max(r^{4},(t_{i+1}-t_{i})^{2})}\leq C^{k}\frac{r^{4}(r^{2})^{k-1}(\log n)^{4}}{n}.$
Here, we used the fact that $\sum_{t=1}^{n}\frac{Cr^{4}}{\max(r^{4},t^{2})}\leq Cr^{2}$. ∎
Proof of Lemma 6.9: Estimate for (6.8).
Define the event
$\mathcal{O}_{i}:=\{L(\mathcal{B}_{\sqrt{n/K_{i}}})\leq\frac{n}{(K_{i})^{1+\delta}}\},$
for some small $\delta>0$. By using Markov’s inequality and equation (6.9), we see that we have
$\mathbb{P}(\mathcal{O}_{i}^{c})\leq\frac{(\log n)^{4}}{K_{i}^{1-\delta}}.$
Thus, we see that,
$\mathbb{P}((\bigcap_{k=11}^{f}\mathcal{O}_{k})^{c})\leq\sum_{k=11}^{f}\frac{(\log n)^{4}}{K_{k}^{1-\delta}}\leq C\frac{(\log n)^{4}}{K_{11}^{1-\delta}}\ll\frac{1}{(\log n)^{2}}.$
In what follows, we will restrict our study to the event
(6.11) $\tilde{\Omega}:=\bigcap_{k=11}^{f}\mathcal{O}_{k}.$
On this event, we will analyze (6.8). Notice that we must have $|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})|\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}\ll 1$. By using subadditivity of the capacity and the fact that the capacity must be less than the volume, we must have that
$\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\leq|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|$
$\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}.$
Now, since $\max_{z\in\mathcal{A}_{i,n}}G(z)\leq C\frac{K_{i+1}}{n}$, then we have,
$\sum_{i=11}^{f-1}\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\max_{z\in\mathcal{A}_{i,n}}G(z)\leq\sum_{i=11}^{f-1}\frac{n}{(K_{i})^{1+\delta}}\frac{K_{i+1}}{n}\leq CK_{11}^{-\delta}.$
6.3.2. The smallest scales
We now continue our analysis of equation (6.8) by considering the contributions from the large scales of $K_{i}$ with $i\leq 11$. We start by giving some lemmas that will be useful for our analysis. In the following, it is assumed that we consider an index with $i\leq 11$.
Notice that by the moment estimate for the capacity of the random walk in four dimensions, we have the following lemma.
Lemma 6.11.
Let $S[0,n]$ be a random walk starting at the origin. For any integer $k>0$ and $\chi>0$, we have that,
(6.12) $\mathbb{P}\left({\rm Cap}(S[0,n])\geq\frac{n}{(\log n)^{1-\chi}}\right)\leq\frac{1}{(\log n)^{k}},$
for any $k\geq 0$.
Proof.
In [12, Lemma 4.6], we have that these are $c>0$ and $C<\infty$ such that
$\mathbb{E}\left[\exp\left(c\frac{(\log n)^{2}}{n}({\rm Cap}(S[0,n])-\mathbb{E}[{\rm Cap}(S[0,n])])\right)\right]\leq C$
and $\mathbb{E}[{\rm Cap}(S[0,n])]\sim\frac{\pi^{2}}{8}\frac{n}{(\log n)}$ (e.g., see [6, Theorem 1.1]). Hence, we can derive (6.12) by applying Markov’s inequality. ∎
In order to discuss the properties of the parts of the random walk inside each annulus, we need to consider the following stopping times.
Define
$T^{i}_{k}=\inf\{t\geq\tau^{i}_{k-1}:S_{t}\leq\sqrt{n/K_{i}}\},$
$\tau^{i}_{k}=\inf\{t\geq T^{i}_{k}:S_{t}\geq 2\sqrt{n/K_{i}}\}.$
These are a sequence of times indicating when the random walk hits that ball of radius $\sqrt{n/K_{i}}$ and later details when it escapes said ball.
The following lemma will ensure that with high probability that there are not too many times in which the random walk will make an excursion into the ball of radius $\sqrt{n/K_{i}}$. Furthermore, each excursion into this ball will lie very close to its expected time.
Lemma 6.12.
Let $M^{i}=\sup\{k:\tau^{i}_{k}\leq n\}$.
Let $\mathcal{H}^{i}$ be the event that
$\mathcal{H}^{i}:=\{M^{i}\leq(\log n)^{C}\}\bigcap_{k=1}^{(\log n)^{C}}\{|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1-\kappa}}\}\bigcap_{k=1}^{(\log n)^{C}}\{|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{n}{K_{i}^{1+\kappa}}\}.$
We claim that there is some $C$ such that $\mathbb{P}((\mathcal{H}^{i})^{c})\ll\frac{1}{(\log n)^{2}}$.
Proof.
Notice that $M^{i}$ can be bounded from above by the number of spheres of radius $\sqrt{n/K_{i}}$ that are necessary in order to cover the full random walk $S[1,n]$. Considering a sufficiently large $C<\infty$, bounding $M^{i}$ with high probability is the estimate of [12, Lemma 3.4].
Note that the probability that $|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{n}{K_{i}^{1-\kappa}}$ is bounded by the probability that a Brownian motion stays within a ball of radius $2\sqrt{n/K_{i}}$ for time $\frac{n}{K_{i}^{1-\kappa}}$. But, this happens with probability less than $\exp[-cK_{i}^{\kappa}]$. By taking a union bound, this is less than $(\log n)^{C}\exp[-cK_{i}^{\kappa}]\ll(\log n)^{-2}$.
Furthermore, the probability that $|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1+\kappa}}$ can be bounded by $\exp[-cK_{i}^{\kappa}]$ by the reflection principle. Again, by taking a union bound, this is less than $(\log n)^{C}\exp[-cK_{i}^{\kappa}]\ll(\log n)^{-2}$. ∎
Proof of Lemma 6.9: Estimate for (6.7).
We can also define the event,
$\mathcal{G}^{i}:=\bigcap_{i=1}^{(\log n)^{C}}\left\{{\rm Cap}\left(S\left[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}\right]\right)\leq\frac{n}{(\log n)^{1-\chi}K_{i}^{1-\kappa}}\right\}.$
By equation (6.12), we see that $\mathbb{P}((\mathcal{G}^{i})^{c})\ll(\log n)^{-2}$.
Finally, we will need the following event,
$\mathcal{S}^{i}:=\left\{L(\mathcal{B}_{2\sqrt{n/K_{i}}})\leq\frac{n}{(K_{i})^{1-\delta}}\right\},$
where we recall the notation $L(\mathcal{B}_{r})$ from Lemma 6.10.
By using equation (6.10) and Markov’s inequality with $k$ large enough, we would have that
$\mathbb{P}((\mathcal{S}^{i})^{c})\ll(\log n)^{-2}.$
Now we consider the intersection of events,
$\Omega^{i}=\mathcal{S}^{i}\cap\mathcal{G}^{i}\cap\mathcal{H}^{i}.$
By combining the probability bounds from earlier, we have $\mathbb{P}((\Omega^{i})^{c})\ll(\log n)^{-2}.$
Notice that, on the event $\Omega^{i}$, we can show that,
(6.13) ${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},\tau^{i}_{k}])\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}])$
$\leq\sum_{k=1}^{M^{i}}\frac{n}{(\log n)^{1-\chi}K_{i}^{1-\kappa}}\leq K_{i}^{2\kappa}\sum_{k=1}^{M^{i}}\frac{n}{(\log n)^{1-\chi}K_{i}^{1+\kappa}}\leq\frac{K_{i}^{2\kappa}}{(\log n)^{1-\chi}}\sum_{k=1}^{M^{i}}|\tau^{i}_{k}-T^{i}_{k}|$
$\leq\frac{K_{i}^{2\kappa}}{(\log n)^{1-\chi}}L(\mathcal{B}_{2\sqrt{n/K_{i}}})\leq\frac{K_{i}^{2\kappa}}{(\log n)^{1-\chi}}\frac{n}{(K_{i})^{1-\delta}}.$
The second inequality used the fact that $|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. The third inequality used the bound on the capacity from the event $\mathcal{G}^{i}$. The fifth inequality used that $|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{1}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. Note that the sum of these differences is a lower bound on the amount of times in the intersection between $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ and $\mathcal{B}_{2\sqrt{n/K_{i}}}$. The final inequality used the event $\mathcal{S}^{i}$.
Now, we see that
${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\max_{z\in\mathcal{A}_{i,n}}G(z)$
$\leq\frac{K_{i}^{2\kappa}n}{(\log n)^{1-\chi}(K_{i})^{1-\delta}}\frac{K_{i+1}}{n}\leq(\log n)^{3(2\kappa+\delta)+\chi}(\log n)^{-\epsilon}.$
By setting $3(2\kappa)+\delta+\chi\leq\epsilon/2$, we see that the line above will be less than $(\log n)^{-\epsilon/2}$. We now consider $\hat{\Omega}:=\bigcap_{i=1}^{11}\Omega^{i}$, we must have $\mathbb{P}(\hat{\Omega}^{c})\ll(\log n)^{-2}$. Furthermore, we must also have
$\sum_{i=1}^{11}{\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\max_{z\in\mathcal{A}_{i,n}}G(z)\leq(\log n)^{-\epsilon/2}.$
Now, consider the intersection $\tilde{\Omega}\cap\hat{\Omega}$, where we recall $\tilde{\Omega}$ from equation (6.11). We must have that $\mathbb{P}((\tilde{\Omega}\cap\hat{\Omega})^{c})\ll(\log n)^{-2}$ and, in addition, the right-hand side of equation (6.8) can be bounded by $(\log n)^{-c}$ for some $c>0$. This completes the proof of Lemma 6.9. ∎
6.4. An analysis of $X_{n}-\mathbb{E}[X_{n}]$
In this section, we will prove that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\to 0$ in distribution. We will first start by proving upper and lower bounds on $\mathrm{Var}(X_{n})$. Note that an upper bound on $\mathrm{Var}(X_{n})$ was already derived in Lemma 6.6. We now derive the corresponding lower bound for $\mathrm{Var}(X_{n})$. Is this subsection, we will give the estimate only for cut points and graph distance.
Lemma 6.13.
Let $X_{n}$ denote the number of cut points, or the graph distance. We have that
$\mathrm{Var}(X_{n})\geq\frac{n^{2}}{(\log n)^{2+o(1)}}.$
Proof.
We perform a dyadic decomposition of $X_{n}$ up to $(\log n)^{\epsilon}$ levels for some $\epsilon>0$. We thus write
$X_{n}=\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}].$
Notice, now that since $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]$ is a sum of i.i.d. terms where each individual term has variance at most $\frac{n^{2}}{(\log n)^{2-o(1)}}$, we can compute its variance as
$(\log n)^{\epsilon}\frac{n^{2}}{(\log n)^{2\epsilon}(\log n)^{2-o(1)}}\leq\frac{n^{2}}{(\log n)^{2+\epsilon-o(1)}}.$
This will not be the major contributor to the variance.
We will now consider the contribution to the variance from $\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}$. Notice, we have,
$\mathrm{Var}\left(\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}\right)$
$=\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1},k_{2},l_{1},l_{2}:(k_{1},l_{1})\neq(k_{2},l_{2})}\mathbb{E}\left[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\right]$
$\geq\mathrm{Var}(E_{n}^{(0,1)})-\sum_{k_{1},k_{2},l_{1},l_{2}:(k_{1},l_{1})\neq(k_{2},l_{2})}\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$
$\geq\frac{n}{(\log n)^{2+o(1)}}-(\log n)^{2\epsilon}\frac{n^{2}}{(\log n)^{3+o(1)}}\gg\frac{n}{(\log n)^{2+o(1)}}.$
We used the fact that $E_{n}^{(k_{1},l_{1})}$ are all positive terms, so the only possible negative contribution to the covariance comes from the product of expectations. However, we know from Corollary 6.7 that the product of expectations of these expectations is bounded from above by $\frac{n^{2}}{(\log n)^{3+o(1)}}$. This completes the proof of our assertion. ∎
With a small modification to the proof above, we can derive the following main result.
Theorem 6.14.
Let $X_{n}$ denote the number of cut points, or the graph distance. We have that as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\to 0,$
in distribution.
Proof.
We proceed with the same dyadic decomposition as we have done in the previous proof.
$X_{n}=\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}].$
Let $\Omega^{(k,l)}$ be the set from Lemma 6.3 (constructed for the appropriate cross term $E_{n}^{(k,l)}$) and consider $\Omega_{A}:=\bigcap_{(k,l)}\Omega^{(k,l)}$. By Lemma 6.3, we can ensure that $(\Omega^{(k,l)})^{c}$ holds with probability no larger than $\text{O}\left(\frac{\log\log n}{\log n}\right)$. By taking a union bound, we see that $\mathbb{P}(\Omega_{A}^{c})=\text{O}\left(\frac{\log\log n}{(\log n)^{1-\epsilon}}\right)\ll 1.$ On the event, $\Omega^{(k,l)}$, we see that
$|-E_{n}^{(k,l)}+\mathbb{E}[E_{n}^{(k,l)}]|\leq\frac{n}{(\log n)^{3/2+o(1)}}.$
Thus, on $\Omega_{A}$, we must have
$\left|\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\mathbb{E}[E_{n}^{(k,l)}]\right|\leq(\log n)^{\epsilon}\frac{n}{(\log n)^{3/2-o(1)}}.$
Furthermore, notice that by Lemma 6.6,
$\mathrm{Var}\left(\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]\right)\leq\frac{n^{2}}{(\log n)^{2+\epsilon+o(1)}}.$
Let $\Omega_{B}$ be the event that $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]-\mathbb{E}[X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]]$ is smaller than $\frac{n}{(\log n)^{1+\epsilon/4}}$. By Markov’s inequality $\mathbb{P}(\Omega_{B}^{c})\leq(\log n)^{-\epsilon/2}$.
Now, on $\Omega_{A}\cap\Omega_{B}$, we have
$X_{n}-\mathbb{E}[X_{n}]\leq\frac{n}{(\log n)^{1+\epsilon/4}}+(\log n)^{\epsilon}\frac{n}{(\log n)^{3/2-o(1)}}\ll\frac{n}{(\log n)^{1+o(1)}}=\sqrt{\mathrm{Var}(X_{n})}.$
Furthermore, $\mathbb{P}(\Omega_{A}\cap\Omega_{B})=1-o(1)$ by a union bound. This completes the proof. ∎
6.5. An analysis of the effective resistance
In this subsection, we present, for independent interest, a similar analysis done in the previous sections when we impose weaker conditions on the bounds we obtain for the variance and the expectation for the cross terms under consideration in $d=4$. Namely, we will assume that the cross-terms are very heavy tailed and that we have some almost optimal upper and lower bounds for the variance up to powers of $\log n$. In particular, we will give the proof of the desired result for the effective resistance in $d=4$ in Subsection 6.5.2. The following assumption will detail what we assume about the cross terms under consideration.
Before this, in Subsection 6.5.1, we will give an estimate for the effective resistance, which assures that Assumption 6.17 parts (2) (3) will hold.
6.5.1. Upper bound for the cross term of the effective resistance
We will give the upper bound for the second moment of the cross term of the effective resistance $E[E_{n}^{2}]$.
Lemma 6.15.
Consider $d=4$. There exists an event $\Omega$ of probability $O(\frac{\log\log n}{\log n})$ such that
(6.14) $E_{n}\mathbbm{1}[\Omega^{c}]\leq\mathbb{E}[E_{n}].$
We have that
(6.15) $\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{1-o(1)}}.$
Remark 6.16.
The constant that appears on the right-hand side of equation (6.15) is not the best possible (rather it should be improved to 1o(1)1-o(1)). However, obtaining the optimal constant is not necessary for the estimates that follow.
Proof.
Recall $\Omega$ in Lemma 6.3. On $\Omega^{c}$, we have that,
$E_{n}\leq 2\frac{n}{(\log n)^{b/2}}.$
Hence, we obtain (6.14). Notice, also that since $E_{n}$ is deterministically less than $n$, we must have that,
$\mathbb{E}[E_{n}^{2}]=\mathbb{E}[E_{n}^{2}\mathbbm{1}[\Omega]]+\mathbb{E}[E_{n}^{2}\mathbbm{1}[\Omega^{c}]]\leq n^{2}\mathbb{P}(\Omega)+4\frac{n^{2}}{(\log n)^{b}}\leq\frac{n^{2}}{(\log n)^{1-o(1)}}.$
Hence, we obtain (6.15). ∎
Proof of Proposition 1.3.
By Lemmas 6.1 and 6.15, we obtain the result, which corresponds to (2) (3) for Assumption 6.17 for $b_{1}=2+o(1)$ and $b_{2}=1-o(1)$. ∎
6.5.2. General arguments
Assumption 6.17.
We let $X[a,b]$ for $a,b\in\mathbb{Z}^{+},b>a$ be a collection of random variables satisfying the following properties. First, define $E_{n}:=X_{n}+X[n,2n]-X_{2n}$
  1. (1)
    The law of $X[a,b]$ depends only on the value of |ba||b-a|.
  2. (2)
    If $a,b]\cap[c,d]=\emptyset$ independent, then $X[a,b]$ is independent of $X[c,d]$.
  3. (3)
    There exist constants $b_{1},b_{2}$ such that
    $\frac{n^{2}}{(\log n)^{b_{1}}}\leq\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{b_{2}}}.$
  4. (4)
    There is a set $\Omega$ with probability $O(\frac{\log\log n}{\log n})$ such that
    $E_{n}\mathbbm{1}[\Omega^{c}]\leq\mathbb{E}[E_{n}].$
We first define the function $f(n)$ as,
$\mathrm{Var}(X_{n})=\frac{n^{2}}{(\log n)^{f(n)}}.$
In order to characterize this, we need to first define the associated function $g(n)$ as as follows:
$\mathbb{E}[E_{n}^{2}]:=\frac{n^{2}}{(\log n)^{g(n)}}.$
We remark that the third part of our Assumption 6.17 shows that we have $b_{1}+o(1)\leq g(n)\leq b_{2}+o(1)$, but we do not get a uniform control of $g(n)$. In addition, as a consequence of the fourth part of Assumption 6.17, we must also have:
(6.16) $\mathbb{E}[E_{n}^{2}]O\left(\frac{\log\log n}{\log n}\right)\geq\mathbb{E}[E_{n}]^{2}.$
To try to get a uniform control of $g(n)$ on a fixed number of scales, we fix a constant $K$ and additionally define the function $g(n,K)$ as:
$g(n,K):=\min_{1\leq j\leq K}g(n2^{-j}).$
This function will be used to characterize the size of the expectation of the square of the graph distance, as detailed in the following lemma.
Lemma 6.18.
Fix some $K>0$. Assume that there exists $n$ such that $f(n2^{-K})\geq g(n,K)+\frac{1}{4}$. Then, we will have that $f(n)\leq g(n,K)+o(1)$, where $o(1)$ is a term that will decay to 0 as $n$ is taken to $\infty$.
Proof.
We start with a dyadic decomposition of the random variables. As a consequence, we observe that,
$X_{n}-\mathbb{E}[X_{n}]=\sum_{j=1}^{2^{K}}(\tilde{X}_{K,j}-\mathbb{E}[\tilde{X}_{K,j}])+\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]).$
We first consider the variance of the term,
$\overline{\mathcal{E}}:=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}(E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]).$
We remark that when computing the second moment $\mathcal{E}^{2}$, the only negative contribution can come from terms of the form,
(6.17) $N_{(k_{1},l_{1}),(k_{2},l_{2})}=-\mathbb{E}[(E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}])\mathbbm{1}[E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]>0]$
$\times(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]$
$\leq\mathbb{E}[(E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}])^{2}\mathbbm{1}[E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]>0]^{1/2}$
$\times\mathbb{E}[(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])^{2}\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]^{1/2}$
$\leq\mathbb{E}[(E_{n}^{(k_{1},l_{1})})^{2}]^{1/2}\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\leq\frac{n2^{-k_{1}}}{(\log(n2^{-k_{1}}))^{g(n2^{-k_{1}})/2}}\frac{n2^{-k_{2}}}{(\log n2^{-k_{2}})^{g(n2^{-{k_{2}}})/2+1/2-o(1)}}$
$\leq\frac{n^{2}}{(\log n)^{g(n,K)+1/2+o(1)}}.$
To get the third inequality, we used the fact that $E_{n}^{(k_{2},l_{2})}$ must be positive. Thus, under the condition that $E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$ is negative, the expectation of $\mathbb{E}[(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])^{2}\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]^{1/2}\leq\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$. The second to last inequality used the definition of $g$ along with equation (6.16) for $n$ large enough. Finally, the last inequality used the fact that $\frac{n}{(\log n)^{k}}$ is increasing for $n$ large enough as well as the definition of $g(n,K)$.
This should contrast with the positive contribution coming from,
$\mathbb{E}[(E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}])^{2}]=\frac{(n2^{-k_{1}})^{2}}{(\log n2^{-k_{1}})^{g(n2^{-k_{1}})}}.$
There will be at least one term $(k_{1},l_{1})$ where $g(n2^{-k_{1}})$ attains its maximum value, and for this specific value $k^{\max}$, we have that,
$\frac{(n2^{-k^{\max}})^{2}}{(\log n2^{-k^{\max}})^{g(n2^{-k^{\max}})}}=\frac{n^{2}}{(\log n)^{g(n,K)+o(1)}}.$
The factor of $o(1)$ is to compensate for possible changes of the constant factor in front. We remark that due to these computations, there is a gap of of a factor of $(\log n)^{1/2}$ between the positive term above and the negative contribution from equation (6.17). This will ensure a sufficiently good lower bound for the expectation of $\overline{\mathcal{E}}$. Indeed, we have that,
(6.18) $\mathbb{E}[\mathcal{E}^{2}]\geq\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},l_{1})}])^{2}]+\sum_{k_{1}=1}^{K}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=1}^{K}\sum_{l_{1}=1}^{2^{k_{2}}}N_{(k_{1},l_{1}),(k_{2},l_{2})}$
$\geq\frac{n^{2}}{(\log n)^{g(n,K)+o(1)}}-K^{2}4^{K}\frac{n^{2}}{(\log n)^{g(n,K)+\frac{1}{2}+o(1)}}\geq\frac{n^{2}}{(\log n)^{g(n,K)+o(1)}}.$
Furthermore, if we define
$\mathcal{I}:=\sum_{j=1}^{2^{K}}(\tilde{X}_{(K,j)}-\mathbb{E}[\tilde{X}_{(K,j)}]),$
we notice that we have,
(6.19) $\mathbb{E}[\mathcal{I}^{2}]\leq 2^{K}\frac{(n2^{-K})^{2}}{(\log n2^{-K})^{f(n2^{-K})}}\leq 2^{K}\frac{(n2^{-K})^{2}}{(\log n2^{-K})^{g(n,K)+\frac{1}{4}}}\leq\frac{n^{2}}{(\log n)^{g(n,K)+\frac{1}{4}}}.$
In the first inequality, we used the fact that $\mathcal{I}$ is a sum of independent random variables. Since $g(n,K)$ lies between $b_{2}+o(1)$ and $b_{1}+o(1)$, the final inequality must be true for large enough $n$.
Notice now, that we have, by the Cauchy-Schwarz inequality,
(6.20) $\mathbb{E}[(\overline{\mathcal{E}}+\mathcal{I})^{2}]\geq(\mathbb{E}[\overline{\mathcal{E}}^{2}]^{1/2}-\mathbb{E}[\mathcal{I}^{2}]^{1/2})^{2}\geq\frac{n^{2}}{(\log n)^{g(n,K)+o(1)}}.$
Remark 6.19.
By (6.18), (6.19) and (6.20), we have $\epsilon>0$,
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{b_{2}+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{b_{1}-\epsilon}}<\infty$
noting $(\mathbb{E}[\overline{\mathcal{E}}^{2}]^{1/2}-\mathbb{E}[\mathcal{I}^{2}]^{1/2})^{2}\leq\mathbb{E}[(\overline{\mathcal{E}}+\mathcal{I})^{2}]\leq(\mathbb{E}[\overline{\mathcal{E}}^{2}]^{1/2}+\mathbb{E}[\mathcal{I}^{2}]^{1/2})^{2}$.
We now present the heuristic argument we will use to analyze the behavior of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$.
  • We start by taking the high probability event $\Omega$ to be the event that all $E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]\leq\mathbb{E}[E_{n}^{(k,l)}]$.
  • If there exists an infinite sequence of $m$ such that we have $f(n2^{-K})$ is greater than $g(n,K)+\frac{1}{4}$, then we have $f(n)\leq g(n,K)-o(1)$. For such $n$, we will be able to show that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ will go to 0.
  • In the case that the above does not happen, it must be the case that eventually all $f(n2^{-K})\leq g(n,K)+\frac{1}{4}<g(n,K)+\frac{1}{2}$. (The latter condition will ensure that the contribution to the variance of the sum of the graph distance terms will be greater than the contribution of the cross terms.) Then, if we find an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$ infinitely often, then we will be able to show for these $n$ that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ will go to 0.
  • Finally, we will show that there must be an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$ infinitely often by contradiction.
We can now prove our main result
Theorem 6.20.
Let $X_{n}$ be a series of random variables that satisfies the assumptions of 6.17. We have that there exists a subsequence such that as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\to}}0.$
Proof.
Initial Estimates : Recall that the function $f(n)$ is defined as $\mathrm{Var}(X_{n})=\frac{n^{2}}{(\log n)^{f(n)}}$. Now, we start by performing the dyadic decomposition to $K$ levels. Thus, we have,
$X_{n}=\sum_{l=1}^{2^{K}}\tilde{X}_{(K,l)}-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}.$
Here, $\tilde{X}_{(K,l)}=X[(l-1)n2^{-K},ln2^{-K}]$ is a collection of independent copies with the same distribution as $X_{n2^{-K}}$ and for each given $K$, $E_{n}^{(k,l)}$ is a collection of independent random variables with the same distribution as $E_{n2^{-k}}$.
We now define the event,
$\Omega_{(k,l)}=\left\{|E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]|\leq\mathbb{E}[E_{n}^{(k,l)}]\right\},\quad\Omega=\bigcap_{k,l}\Omega_{(k,l)}.$
By the third Assumption 6.17, we see that $\Omega$ occurs with probability 1o(1)1-o(1). We define $\mathcal{E}=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$ and observe that,
(6.21) $\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]$
$\leq$ $\sum_{(k_{1},l_{1},k_{2},l_{2})}\mathbb{E}[(E_{n}^{(k_{1},l_{1})}-\mathbb{E}[E_{n}^{(k_{1},j_{1})}])^{2}\mathbbm{1}[\Omega_{(k_{1},l_{1})}]]^{1/2}\mathbb{E}[(E_{n}^{(k_{2},j_{2})}-\mathbb{E}[E_{n}^{(k_{2},j_{2})}])^{2}\mathbbm{1}[\Omega_{(k_{2},j_{2})}]]^{1/2}$
$\leq$ $2^{2K+2}4\max_{(k,l)}(\mathbb{E}[E_{n}^{(k,l)}]^{2})\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}.$
In the last equation, we used equation (6.16) to bound $\mathbb{E}[E_{n}^{(k,l)}]^{2}$ relative to $\mathbb{E}[(E_{n}^{(k,l)})^{2}]$ and we placed irrelevant constant factors into an $o(1)$ exponent of $\log n$ in the denominator.
In addition, we have that,
(6.22) $\mathbb{E}\left[\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)^{2}\mathbbm{1}[\Omega]\right]$
$\leq$ $\mathbb{E}\left[\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)^{2}\right]=\sum_{l=1}^{2^{K}}\mathbb{E}[(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])^{2}]$
$=$ $2^{K}\frac{(n2^{-K})^{2}}{(\log n2^{-K})^{f(n2^{-K})}}=\frac{n^{2}}{(\log n)^{f(n)}}2^{-K}\frac{(\log n)^{f(n)}}{(\log n2^{-K})^{f(n2^{-K}})}.$
Analysis of properties of $f$
Now, we seek to analyze various cases.
Case 1: There exists an infinite sequence of $n$ such that $f(n2^{-K})\geq g(n,K)+\frac{1}{4}$
In this case, we observe that Lemma 6.18 would imply that $f(n)\leq g(n,K)+o(1)$ along this infinite sequence of $n$. Furthermore, we would have that,
$\frac{n^{2}}{(\log n)^{f(n)}}2^{-K}\frac{(\log n)^{f(n)}}{(\log n2^{-K})^{f(n2^{-K}})}\leq C\frac{n^{2}}{(\log n)^{f(n)}}\frac{1}{(\log n)^{1/4+o(1)}}\ll\frac{n^{2}}{(\log n)^{f(n)}},$
when $n$ is sufficiently large. Indeed, we used that for $n$ large and fixed $K$, we can bound $\frac{1}{(\log n2^{-K})^{f(n2^{-K})}}\leq\frac{1}{(\log n)^{g(n,K)+\frac{1}{4}+o(1)}}$ where the $o(1)$ term in the exponent is introduced to from $\log n2^{-K}\to\log n$ in the denominator. Recall, we assume $f(n2^{-K})\geq g(n,K)+\frac{1}{4}.$ In the numerator, we bounded $(\log n)^{f(n)}$ by $(\log n)^{g(n,K)+o(1)}$.
Combining this estimate of equation (6.22), we see that
$\mathbb{E}\left[\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)^{2}\mathbbm{1}[\Omega]\right]\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
Thus, there will be some set $\Omega^{\prime}$, which will still hold with probability 1o(1)1-o(1) such that,
$\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)<\epsilon\sqrt{\mathrm{Var}(X_{m})}.$
Furthermore, recall from equation (6.21) that the sums of the cross terms $\mathcal{E}$ satisfies,
$\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
Again, there will be a further subset $\Omega^{\prime\prime}\subset\Omega^{\prime}$ that will still hold with probability 1o(1)1-o(1) on which we have additionally that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|<\epsilon\sqrt{\mathrm{Var}(X_{n})}$. Thus, on $\Omega^{\prime\prime}$, we have $|X_{n}-\mathbb{E}[X_{n}]|<2\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
By our assumption, we can do this for an infinite sequence of $n$.
Case 2: Eventually, all $f(n2^{-K})\leq g(n,K)+\frac{1}{4}$
This can be divided into two more subcases.
Case 2a: There does not exists an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
In this case, we must be able to find a sequence $f(a2^{mK})$ such that
$f(a2^{mK})-f(a2^{(m-1)K})\geq\frac{1}{\log\log(a2^{mK})}.$
Now, since $\sum_{m=1}^{\infty}\frac{1}{\log\log(a2^{mK})}$ diverges, this must mean that $f(a2^{mK})$ is unbounded from above. This contradicts our assumption that $f(n)\leq g(n,K)+\frac{1}{4}$ for $x$ large enough. Thus, we do not have to worry about this case.
The only remaining case is the following.
Case 2b: There exists an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
In what follows, we will perform computations on values $n$ that satisfy $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$. In this case, we notice that,
$\frac{(\log n)^{f(n)}}{(\log n2^{-K})^{f(n2^{-K}})}\leq(\log n)^{f(n)-f(n2^{-K})}\left(\frac{\log n}{\log n-K\log 2}\right)^{f(n2^{-K})}$
$\leq$ $(\log n)^{\frac{1}{\log\log n}}\left(\frac{\log n}{\log n-K\log 2}\right)^{2+\frac{1}{4}}\leq 2e,$
where the final inequality will hold for large enough $m$. Notice that we also used that $f(n2^{-K})\leq g(n,K)+\frac{1}{4}\leq\frac{3}{2}+\frac{1}{4}+o(1)$, by our assumption that eventually all $f(n)\leq g(n,K)+\frac{1}{4}$ and our upper bound on $g(n,K)$.
Returning to equation (6.22), we can bound the expectation of the square of $\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])$ by $2e2^{-K}\mathrm{Var}(X_{n})$ for large enough $n$. Thus, by Markov’s inequality, we have that,
$|\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})},$
with probability at most $2e2^{-K}\epsilon^{-2}$.
Furthermore, recall again from equation (6.21) that we have,
$\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
Again, here we used that first we are considering an $m$ such that first we have $f(n)\leq f(n2^{-K})+\frac{1}{\log\log n}$. Secondly, under the general assumption of case 2, we have $f(n2^{-K})+\frac{1}{\log\log n}\leq g(n,K)+\frac{1}{4}+\frac{1}{\log\log n}\leq g(n,K)+1+o(1)$ for $n$ large enough. Thus, we have $f(n)\leq g(n,K)+1+o(1)$ for some infinite sequence of $n$. Thus (again by Markov’s inequality on the set $\Omega$), with probability $o(1)$, we have that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
In total, we see that with probability at most $o(1)+2e2^{-K}\epsilon^{-2}$, we have that
$|X_{n}-\mathbb{E}[X_{n}]|\geq 2\epsilon\sqrt{\mathrm{Var}(X_{n})}.$
Finishing the Argument
Notice that in all cases, for a fixed $K$, we can find a subsequence $X_{n_{1}},X_{n_{2}},\ldots$ such that with probability at least $1-4e2^{K(1-q)}\epsilon^{-2}$, we must have that
$|X_{n_{i}}-\mathbb{E}[X_{n_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{n_{i}})}$
as $K\to\infty$, we see that $1-4e2^{-K}\epsilon^{-2}\to 0$. By a diagonalization argument, we can find a subsequence $\tilde{n}_{i}$ such that $|X_{\tilde{n}_{i}}-\mathbb{E}[X_{\tilde{n}_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{\tilde{n}_{i}})}$ with probability going to 1 for each $\epsilon$. This subsequence converges in distribution to 0. ∎
Reference
  • [1] A. Adhikari and I. Okada. Moderate Deviations for the Capacity of the Random Walk range in dimension four, to appear in Ann. Probab.
  • [2] A. Adhikari and J. Park. Capacity of the range of random walk: Moderate deviations in dimensions 4 and 5. (https://arxiv.org/abs/2507.05585)
  • [3] M. T. Barlow. Random walks on supercritical percolation clusters. Ann. Probab. 32, 3024–3084, 2004.
  • [4] E. Bolthausen, A.-S. Sznitman, and O. Zeitouni. Cut points and diffusive random walks in random environment, Ann. Inst. H. Poincaré Probab. Stat. 39, no. 3, 527–555, 2003.
  • [5] K. Burdzy and G. F. Lawler. Rigorous exponent inequalities for random walks. J. Phys. A Math. Gen. 23(1), L23–L28 1990.
  • [6] Y. Chang. Two observations on the capacity of the range of simple random walk on Z3 and Z4. Electron. Com. Probab. 22 (2017), No. 25, 9 pp.
  • [7] D. A. Croydon. Random walk on the range of random walk. J. Stat. Phys. 136(2):349–372, 2009.
  • [8] D. A. Croydon. Scaling limits of stochastic processes associated with resistance forms. Ann. Inst. H. Poincaré Probab. Stat. 54(4):1939–1968, 2018.
  • [9] D. A. Croydon and B. M. Hambly. Local limit theorems for sequences of simple random walks on graphs. Potential Anal. 29(4):351–389, 2008.
  • [10] D. A. Croydon, B. M. Hambly and T. Kumagai. Time-changes of stochastic processes associated with resistance forms. Electron. J. Probab. 22:82, 2017.
  • [11] D. A. Croydon and D. Shiraishi. Scaling limit for random walk on the range of random walk in four dimensions. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 59 (1):166–184, 2023.
  • [12] A. Dembo and I. Okada. Capacity of the range of random walk: The law of the iterated logarithm. Ann. Probab. 52, 5, 1954-1991, 2024.
  • [13] N. C., Jain and S. Orey. On the range of random walk. Israel J. Math. 6 , 373-380, 1968.
  • [14] T. Kumagai. Random walks on disordered media and their scaling limits. Lecture notes from the 40th Probability Summer School held in Saint-Flour, 2010.
  • [15] G. F. Lawler. Intersections of Random Walks. Birkhäuser, 1991.
  • [16] G. F. Lawler. Cut times for simple random walk. Electron. J. Probab. 1:13, 1996.
  • [17] G. F. Lawler and V. Limic. Random Walk: A Modern Introduction. Cambridge Univ. Press, 2010.
  • [18] J.-F. Le Gall. Propriétés d’intersection des marches aléatoires. I. Convergence vers le temps local d’intersection. Comm. Math. Phys. 104, no. 3, 471–507, 1986.
  • [19] H. Kesten. The incipient infinite cluster in two-dimensional percolation. Probab. Theory Related Fields 73, 369-394, 1986.
  • [20] D. Shiraishi. Heat kernel for random walk trace on $\mathbb{Z}^{3}$ and $\mathbb{Z}^{4}$. Ann. Inst. H. Poincaré Probab. Stat. 46(4):1001–1024, 2010.
  • [21] D. Shiraishi. Exact value of the resistance exponent for four dimensional random walk trace. Probab. Theory Relat. Fields 153(1–2):191–232, 2012.
  • [22] D. Shiraishi. Growth exponent for loop-erased random walk in three dimensions. Ann. Probab. 46(2): 687-774, 2018.

### v2

@@
 We investigate fluctuation phenomena for the graph distance and the number of cut points associated with random media arising from the range of a random walk.
-Our results demonstrate a sequence of dimension-dependent phase transitions in the scaling behavior of these fluctuations, leading to qualitatively different regimes across dimensions lower than six, equal to six, and higher than six.
-In particular, in six dimensions, we show that the convergence to the normal distribution occurs under a scaling different from that of the standard central limit theorem.
-We also give the related results for the effective resistance of the random walk range.
+Our results demonstrate a sequence of dimension-dependent phase transitions in the scaling behavior of these fluctuations, leading to qualitatively different regimes, with a distinct phase transition in dimension 6.
+In particular, we remark that convergence in dimension 6 occurs with a non-standard rescaling.
 Research supported in part by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists (No. JP24K16931) (I.O.).
-Research partially supported by NSF grant DMS 2102842
 1. Random walk on random walk range
 Consider the simple random walk $(S_{n})_{n=0}^{\infty}$ on ${\mathbb{Z}}^{d}$ for $d\geq 4$. We define the graph of the random walk range as follows:
-$V({\mathcal{G}}):=\{S_{m}:m\geq 0\},$
-$E({\mathcal{G}}):=\{\{S_{m},S_{m+1}\}:m\geq 0\}.$
-Originally, Lawler and Burdzy [5] discussed the exponents for the fluctuation of the random walk range and proposed many open problems.
-Croydon and Shiraishi [11, 20, 21] estimated the graph distance, the effective resistance of $(V({\mathcal{G}}),E({\mathcal{G}}))$ and the number of cut points of the random walk range between $S_{0}$ and $S_{n}$ in $d=4$, and showed that their typical order is $n/(\log n)^{1/2+o(1)}$.
-In addition, Croydon [7] did similar work in $d\geq 5$, and showed that their typical order is linear.
-In [22], he discussed the loop erased random walk and other related concepts for the graph of the random walk range in $d=2,3$.
+$V({\mathcal{G}}):=\{S_{m}:m\geq 0\},\quad E({\mathcal{G}}):=\{\{S_{m},S_{m+1}\}:m\geq 0\}.$
+Originally, Lawler and Burdzy [5] discussed the exponents for the fluctuations of the random walk range and proposed many open problems.
+Croydon and Shiraishi [11, 20, 21] estimated the graph distance, the effective resistance of $(V({\mathcal{G}}),E({\mathcal{G}}))$ and the number of cut points of the random walk range between $S_{0}$ and $S_{n}$ in $d=4$, and showed that their typical orders are $n/(\log n)^{1/2+o(1)}$.
+In addition, Croydon [7] did similar work in $d\geq 5$ and showed that their typical order is linear.
+In [22], he discussed the loop-erased random walk and other related concepts for the graph of the random walk range in $d=2,3$.
 As our motivation, random walks in random media have attracted considerable interest as mathematical models for transport phenomena in disordered environments.
@@
 We refer the reader to [14] for further references on these topics.
-An important tool for analyzing random walks in random media is the effective resistance or graph distance associated with the underlying random media.
-In particular, the behavior of the effective resistance is closely connected to the spectral dimension of the random subset, which characterizes the long-time decay of return probabilities and provides a key indicator of whether the associated random walk exhibits diffusive or anomalous scaling.
 Independent of the study of random walks on random media, understanding the graph distance in random media is of intrinsic interest.
@@
 In particular, in light of recent developments, it is conjectured that graph distances in several models of random media exhibit non-trivial scaling behavior, potentially related to general universality phenomena.
-These considerations provide strong motivation for a detailed analysis of graph distances in random media, which is a focus of the present work.
-Motivated by these considerations, we aim to understand the fluctuation behavior of the effective resistance and the graph distance in random media.
+These considerations provide strong motivation for a detailed analysis of graph distances in random media, which is the focus of the present work.
+Motivated by these considerations, we aim to understand the fluctuation behavior of the graph distance in random media.
 In this paper, we take the range of a random walk as a concrete example of a random medium.
@@
 In particular, we find distinct regimes in dimensions four, five, six, and seven and higher.
-A major phase transition occurs at dimension six: for $d\geq 6$, the fluctuations, when normalized by the square root of the variance, converge in distribution to a Gaussian random variable, whereas in dimensions four and five, the same normalization leads to convergence to zero.
+A major phase transition occurs at dimension six: for $d\geq 6$, the fluctuations, when normalized by the square root of the variance, converge in distribution to a Gaussian random variable, whereas in dimensions four and five, the same normalization leads to convergence of the fluctuation to zero.
 We also note that the variance of the normal random variable in $d=6$ that appears in the central limit theorem is non-standard in the sense that it is actually not equal to the variance of the original random variable.
-Moreover, our proofs suggest that in dimensions four and five, if the fluctuations are rescaled by the appropriate order, they are expected to converge to non-Gaussian random variables.
 We will give the definitions necessary to state the main results. We consider the graph ${\mathcal{G}}[a,b]$ as the set of edges constructed by the random walk trace:
@@
 $\sum_{i=2}^{n}1\{S[1,i-1]\cap S[i,n]=\emptyset\}.$
-We define a quadratic form $\mathcal{E}$ on functions $f,g:V(G)\to\mathbb{R}$ by
-$\mathcal{E}(f,g)=\frac{1}{2}\sum_{\begin{subarray}{c}x,y\in V\\ \{x,y\}\in E\end{subarray}}(f(x)-f(y))(g(x)-g(y)).$
-This corresponds to the energy dissipation of the resistance electrical network on $G$. We also introduce the space
-$H^{2}:=\{\,f\in\mathbb{R}^{V}:\mathcal{E}(f,f)<\infty\,\},$
-which serves as the natural domain of the energy form. The effective resistance between $A$ and $B$ is defined via the energy minimization
-$R_{G}(A,B)^{-1}\;=\;\inf\Biggl\{\mathcal{E}(f,f)\;\Big|\;\quad f|_{A}=1,\ f|_{B}=0,f\in H^{2}\Biggr\}.$
-This quantity represents the voltage difference required to send the current from $A$ to $B$ and therefore corresponds to the usual notion of effective resistance in an electrical network.
-For vertices $x,y\in V$, we write $R_{G}(x,y)$ to denote $R_{G}(\{x\},\{y\})$.
-We will now define the notation used to represent the graph distance, the effective resistance, and the number of cut points.
-Let $X^{1}_{[a,b]}(x,y):=d_{\mathcal{G}_{[a,b]}}(x,y)$, $X^{2}_{[a,b]}(i,j)$ be the number of cut points on $\mathcal{G}_{[a,b]}$ between the time $i$ and $j$, and $X^{3}_{[a,b]}(x,y):=R_{\mathcal{G}_{[a,b]}}(x,y)$.
-A trivial relationship among these variables is $X^{3}_{[a,b]}(x,y)\leq X^{1}_{[a,b]}(x,y)$ and $X^{2}_{[a,b]}(i,j)\leq X^{3}_{[a,b]}(S_{i},S_{j})$.
-We sometimes write $X^{i}[a,b]$ as $X^{i}_{[a,b]}(S_{a},S_{b})$ and $X^{i}_{n}=X^{i}_{[0,n]}(S_{0},S_{n})$ for $i=1,3$ and $X^{2}[a,b]$ as $X^{2}_{[a,b]}(a,b)$ and $X^{2}_{n}=X^{2}_{[1,n]}(1,n)$.
+We will now define the notation used to represent the graph distance and the number of cut points.
+Let $X^{1}_{[a,b]}(x,y):=d_{\mathcal{G}_{[a,b]}}(x,y)$ and $X^{2}_{[a,b]}(i,j)$ be the number of cut points on $\mathcal{G}_{[a,b]}$ between the time $i$ and $j$.
+A trivial relationship among these variables is $X^{1}_{[a,b]}(S_{i},S_{j})\geq X^{2}_{[a,b]}(i,j)$.
+We sometimes write $X^{i}[a,b]$ as $X^{i}_{[a,b]}(S_{a},S_{b})$ and $X^{1}_{n}=X^{1}_{[0,n]}(S_{0},S_{n})$ and $X^{2}[a,b]$ as $X^{2}_{[a,b]}(a,b)$ and $X^{2}_{n}=X^{2}_{[1,n]}(1,n)$.
 Since the following proof is the same, we omit the symbol “ii”, wherever it appears.
@@
 We write $f(n)\lesssim g(n)$ if there exists a (deterministic) constant $c>0$ such that $f(n)\leq cg(n)$ for all $n$, and $f(n)\gtrsim g(n)$ if $g(n)\lesssim f(n)$.
-First, we will give our main results for the graph distance and the number of cut points:
+We now present our main results on the fluctuation of the graph distance and the number of cut points.
 Proposition 1.1.
-It holds that, as $n\to\infty$,
-$\mathrm{Var}(X_{n})\quad\begin{cases}\sim\alpha^{i}_{d}n&\text{for }d\geq 7,\\ \asymp n\log n&\text{for }d=6,\\ \asymp n^{3/2}&\text{for }d=5,\\ =\frac{n^{2}}{(\log n)^{2\pm o(1)}}&\text{for }d=4,\end{cases}$
-for some constant $\alpha^{i}_{d}>0$ for $i=1,2$.
+Consider $i=1,2$. It holds that, as $n\to\infty$,
+$\mathrm{Var}(X_{n}^{i})\quad\begin{cases}\sim\alpha^{i}_{d}n&\text{for }d\geq 7,\\ \asymp n\log n&\text{for }d=6,\\ \asymp n^{3/2}&\text{for }d=5,\\ =\frac{n^{2}}{(\log n)^{2\pm o(1)}}&\text{for }d=4,\end{cases}$
+for some constant $\alpha^{i}_{d}>0$ with $i=1,2$.
 Theorem 1.2.
-For $d\geq 7$, as $n\to\infty$,
-$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
+Consider $i=1,2$. For $d\geq 7$, as $n\to\infty$,
+$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\mathrm{Var}(X_{n}^{i})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
 and for $d=6$, there is a deterministic function $\phi(n)\asymp n\log n$ such that as $n\to\infty$,
-$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\phi(n)}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
+$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\phi(n)}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
 For $d=4,5$, as $n\to\infty$,
-$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
-Next, we will give the result for the effective resistance.
-Proposition 1.3.
-For $d\geq 7$ and $\alpha^{3}_{d}>0$, as $n\to\infty$,
-$\mathrm{Var}(X_{n})\sim\alpha^{3}_{d}n.$
-For $d=6$,
-$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n\log n}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n(\log n)^{2}}<\infty.$
-For $d=5$,
-$\mathrm{Var}(X_{n})\approx n^{3/2}.$
-For $d=4$ and $\epsilon>0$,
-$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{2+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{1-\epsilon}}<\infty.$
-Theorem 1.4.
-For $d\geq 7$, as $n\to\infty$,
-$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
-for $d=6$, there is a subsequence $(n_{k})$ and $\phi(n)\geq cn(\log n)$ for some $c>0$ such that as $n\to\infty$,
-$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\phi(n_{k})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
-For $d=5$, as $n\to\infty$,
-$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0,$
-and for $d=4$, there is a subsequence $(n_{k})$ such that as $n\to\infty$,
-$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\mathrm{Var}(X_{n_{k}})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
-Open problem 1.5.
+$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\mathrm{Var}(X_{n}^{i})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
+Open problem 1.3.
 Does a strong law of large numbers hold for $X_{n}$?
@@
 Even if this obstacle could be overcome, especially in the case $d=4$, the variance is not much smaller than that of $\mathbb{E}[X_{n}]^{2}$ (they differ only by a factor of $\log n$" in $d=4$), so it seems unlikely that the result can be proved using standard concentration inequalities.
-Remark 1.6.
-We believe that the stronger results of Propositions 1.1 and 1.2 are true for the effective resistance as well, but we avoid writing the extra technical details for simplicity.
-Instead, we present a weaker version that holds under weaker assumptions, such as without any correlation assumptions on the cross terms; we believe that the analysis performed in this section is of independent interest and can be readily applicable to any random variable on the random walk range that is governed by heavy tailed behavior.
-As a key ingredient of our approach to the analysis of fluctuations, we employ a decomposition of the random walk path.
-More precisely, we split the trajectory $S[0,n]$ into two subpaths $S[0,n/2]$ and $S[n/2,n]$, and compare the corresponding effective resistances or graph distances.
-The quantity
-X[0,n/2]+X[n/2,n]X[0,n]X[0,n/2]+X[n/2,n]-X[0,n]
-naturally arises as error terms.
-By iterating this decomposition procedure, we obtain a representation consisting of a sum of finely decomposed random variables, which behave approximately as independent and identically distributed terms, together with accumulated error terms.
-A key observation is that a phase transition occurs depending on the dimension: while in dimensions $d\geq 6$ the sum of the i.i.d.-type terms dominates the fluctuations, in lower dimensions the error terms become dominant and significantly affect the scaling behavior.
-Consequently, it is expected that for $d\geq 6$ the fluctuations converge in distribution to a Gaussian random variable, whereas such Gaussian behavior does not persist in lower dimensions.
-We outline the main technical difficulties in the proof of our results.
-For the graph distance and the number of cut points, establishing the validity or failure of the central limit theorem, which constitutes our main theorem, requires a precise and delicate analysis of the variance.
-In particular, this analysis necessitates a careful control of the correlations among the error terms arising from the decomposition employed in the proof.
-As a first step, it is necessary to improve the result of [4, Lemma 1.1] by identifying the optimal strategy for events in the large deviation regime that give rise to large error terms.
-We show that such events correspond to situations in which the random walk forms a large loop.
-The second step consists in estimating the correlations of these error terms.
-More precisely, one must show that these correlations are sufficiently small when the time intervals associated with the decomposition that generates the error terms are well separated.
-A technical difficulty to show the validity or failure of the central limit theorem is the fact that the large deviation probabilities governing these error terms exhibit heavy-tailed behavior.
-In fact, a related phenomenon for the decomposition was already observed in earlier work of the authors [1, 2, 12] concerning the capacity of the range of a random walk.
-In that setting, when analyzing fluctuations via a similar decomposition scheme, a phase transition occurred between dimensions four and five, depending on which term became dominant in the decomposition.
-While this qualitative behavior is closely related to the phenomena studied in the present paper, the analysis in our current setting requires substantially different techniques.
-More specifically, in the previous work the error terms arising from the decomposition became negligible in dimensions $d\geq 5$, whereas in the present case the error terms remain of polynomial order even for $d\geq 6$, necessitating a much more delicate analysis.
-Moreover, although in dimension four the error terms are leading in both settings, their control in the present work requires global information on the entire path, leading to more intricate estimates.
-We also obtained partial results for the effective resistance; however, the resulting bounds are weaker than those for the other quantities.
-This is due to technical difficulties arising in the estimation of the correlations of the error terms.
-Nevertheless, even without a detailed analysis of these correlations, once an estimate of the error terms corresponding to an improvement of [4, Lemma 1.1] is available, we are able to obtain an analogy that yields a phase transition for the central limit theorem of the effective resistance under weaker assumptions on the error terms.
-We believe that this analogy is useful for a broad class of random variables for which a decomposition of the type considered in this work can be carried out.
+We emphasize that the theorem above on the fluctuations of the graph distance and the number of cut points demonstrates a key phase transition in dimension 6.
+While in dimensions strictly greater than 6, one has the standard central limit theorem, this fails in dimensions less than 6.
+The standard scaling for the fluctuation leads to a trivial random variable in dimensions 4 and 5.
+In addition, dimension 6 presents a subtle, but deep change in the behavior of the fluctuation; instead of fluctuations on the scale of $\sqrt{n}$, as in higher dimensions, the fluctuation will be of the scale $\sqrt{n\log n}$.
+Moreover, the precise order of the fluctuation in dimension 6 is not given by the square root of the variance, but differs from the variance by a constant.
+Indeed, by looking at the proof, one can intuit that the main contribution to the fluctuation in dimensions higher than seven arises primarily from local fluctuations in the random walk range.
+In dimension 6, the proof demonstrates that the primary source of the fluctuation arises from interactions at intermediate scales of order $\sqrt{n}$.
+As such, in dimension 6, we observe the first demonstration of the effect of long-range interactions for the graph distance of the random walk range.
+We remark that the source of this new phase transition is due to the effect of the heavy tailed nature of the long range interactions of the random walk that can cause significant effects on the graph distance.
+One of the key ingredients of our proof is to quantify the effect of the long-range interactions of the random walk range by a decomposition.
+On the highest level, one can measure the effect of the interaction between the first and second halves of the random walk by splitting the trajectory $S[0,n]$ into two subpaths $S[0,n/2]$ and $S[n/2,n]$.
+Once this is done, one can compare the corresponding graph distances by studying the variable
+X[0,n/2]+X[n/2,n]X[0,n],X[0,n/2]+X[n/2,n]-X[0,n],
+that naturally arises as an error term.
+By iteratively applying this decomposition procedure, we obtain a representation consisting of a sum of finely decomposed random variables that behave approximately as independent and identically distributed terms, together with accumulated error terms.
+The key question here is to determine which of the terms contributes the most to the variance, and on which scale they operate.
+In dimensions $d\geq 7$, one can very finely decompose the random variables and show that local contributions will contribute the most to the fluctuations; since the local contributions are independent of each other, this will lead to a Gaussian fluctuation.
+In dimensions 4 and 5, one will observe that the variance almost entirely comes from the error term at the largest scale; namely, it comes from considering X[0,n/2]+X[n/2,n]X[0,n]X[0,n/2]+X[n/2,n]-X[0,n].
+In dimension 6, there becomes a key interplay between the contributions of the error terms and the sum of the independent terms.
+Indeed, fully decomposing the random walk will not lead to a correct understanding of the fluctuation; rather, the most important contributions to the fluctuation in dimension 6 come from intermediate scales of order $\sqrt{n}$.
+Due to the heavy tails, the probability that the error term X[0,n/2]+X[n/2]X[0,n]X[0,n/2]+X[n/2]-X[0,n] will make significant contributions to the fluctuation is around $n^{-1}$.
+At the intermediate scale of size $\sqrt{n}$, we will have $\sqrt{n}$ terms that have a probability $(\sqrt{n})^{-1}$ of contributing non-trivially.
+Thus, the intermediate scales of size $\sqrt{n}$ are the first scales that start to contribute with high probability to the fluctuation.
+The total fluctuation of $n\log n$ comes from accumulating all the interactions from scale $\sqrt{n}$ or lower and carefully taking into account their interactions.
+This is in contrast to the naive computation of the variance, which accumulates some level of contributions from all scales, even those that are not seen in the fluctuation.
+Many of the key challenges in determining the fluctuations in the graph distance or the number of cut points require a deep understanding of the nature of the error terms that arise in our analysis.
+In contrast to the study of many other variables for the random walk, we remark that the number of cut points or the graph distance is a purely discrete quantity.
+That is, one cannot understand the graph distance by considering the scaling limit of the random walk to the Brownian motion, and then analyzing the appropriate statistics of the Brownian motion.
+As a result, there is a lack of closed formulas that can be used to analyze the error terms.
+Another aspect of the analysis of the error terms is that they are very heavy tailed.
+This occurs because large values of this error term are governed by events that, while rather unlikely, involve very long-range dependencies in the random walk.
+This behavior of the error terms causes two fundamental problems for our analysis.
+First, long-range dependencies cause the error terms at different scales to become correlated.
+Secondly, one now has to take into account a balance between the rarity of a given interaction and the effect of a given error term to determine whether it contributes to the scale of our fluctuations.
+Solving these problems requires a very precise analysis of the covariance structure of the error terms.
+In particular, many previous estimates we found in the literature were insufficient for our purposes.
+In higher dimensions, we sharpen a result of [4, Lemma 1.1] to identify the optimal order of fluctuation of the given error terms and remove an important $O(\log n)$ factor.
+A delicate analysis on the correlation of cut times and long-range intersections of the random walk allows us to determine the covariance structures of the cross terms in dimensions $d\geq 5$ (observe the proofs of Lemma 4.1 and Lemma 5.1).
+In dimension 4, we used a link between intersection properties and the capacity of the random walk, combined with a multiscale analysis in a series of lemmas, in order to decouple the interaction between two subpaths of a random walk and the graph distance on either subpath and obtain optimal estimates on the variance of the cross terms in Section 6.
+Finally, in the proof of Theorem 1.2, we performed the careful balancing act that allowed us to exactly determine the correct order of fluctuation; this involved showing that, due to the heavy tails, the highest scales occur too rarely, and then carefully analyzing the correlations between the cross terms at intermediate scales.
+Our proof reveals that the intermediate scales of order $\sqrt{n}$ determined the main contribution to the fluctuation in six dimensions.
+In addition to our analysis of the graph distance, we present, for independent interest, an analysis of quantities for the random walk range via our decomposition strategy, whose error terms exhibit the same heavy tailed behavior as the graph distance or the number of cut points.
+In this more general analysis, we do not make any assumptions about the correlations of the cross term.
+As a result, it will not be possible to obtain the same sharp results as one could obtain for the graph distance; however, our analysis demonstrates that heavy tailed behavior, by itself, is enough to demonstrate a type of phase transition similar to those shown for the graph distance.
+Indeed, we believe that these more general results would be helpful for analyzing other random variables of interest, and, furthermore, demonstrate the key role of the size of the heavy tails in the cross terms in creating a phase transition for the fluctuations.
+Proposition 1.4.
+Let $X^{3}_{[a,b]}$ be any collection positive random variables that can be written as a measurable function, $\mathcal{F}(S[a,b])$, with respect to the random walk range $S[a,b]$ (with $X^{3}_{n}:=X^{3}_{[0,n]}$).
+Here, we treat $S[a,b]$ as a subgraph of $\mathbb{Z}^{d}$.
+We further assume that $\mathcal{F}$ is translation invariant and can be bounded from above by the volume.
+Define $E_{n}:=X^{3}_{n}+X^{3}_{[n,2n]}-X^{3}_{[0,2n]}$ and assume further that $E_{n}$ satisfies the bounds
+$\mathbb{P}(E_{n}\geq l)\asymp l^{-(d/2-2)},\quad d\geq 5,$
+and $E_{n}$ is always positive, that is, $\mathcal{F}$ is subadditive.
+For $d=4$, the right-hand side above does not meaningfully characterize heavy tailed behavior, so the natural analogs of the heavy tailed assumption are those found in Assumption 6.16.
+Then, for $d\geq 7$ and $\alpha^{3}_{d}>0$, as $n\to\infty$,
+$\mathrm{Var}(X_{n}^{3})\sim\alpha^{3}_{d}n.$
+For $d=6$,
+$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n\log n}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n(\log n)^{2}}<\infty.$
+For $d=5$,
+$\mathrm{Var}(X_{n}^{3})\approx n^{3/2}.$
+For $d=4$ and $\epsilon>0$,
+$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n^{2}/(\log n)^{2+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n^{2}/(\log n)^{1-\epsilon}}<\infty.$
+Theorem 1.5.
+Let $X_{n}^{3}$ satisfy the assumptions outlined in the previous lemma. For $d\geq 7$, as $n\to\infty$,
+$\frac{X_{n}^{3}-\mathbb{E}X_{n}^{3}}{\sqrt{\mathrm{Var}(X_{n}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
+for $d=6$, there is a subsequence $(n_{k})$ and $\phi(n)\geq cn(\log n)$ for some $c>0$ such that as $n\to\infty$,
+$\frac{X_{n_{k}}^{3}-\mathbb{E}X_{n_{k}}^{3}}{\sqrt{\phi(n_{k})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
+For $d=5$, as $n\to\infty$,
+$\frac{X_{n}^{3}-\mathbb{E}X_{n}^{3}}{\sqrt{\mathrm{Var}(X_{n}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0,$
+and for $d=4$, there is a subsequence $(n_{k})$ such that as $n\to\infty$,
+$\frac{X_{n_{k}}^{3}-\mathbb{E}X_{n_{k}}^{3}}{\sqrt{\mathrm{Var}(X_{n_{k}}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
 2. Preparations in $d\geq 5$
@@
 $N(I,[a,b]):=\sum_{n\in I}1\{S[a,n-1]\cap S[n,b]=\emptyset\}$
-and $T^{1}$ (resp. $T^{-1}$) be the first global cut point after (resp. before) the origin (see the details in [4, Lemma 1.1]). Note that $T^{1}$ is the same distribution as that of $T^{-1}$ and
+and $T^{1}$ (resp.
+$T^{-1}$) be the first global cut point after (resp.
+before) the origin (see the details in [4, Lemma 1.1]).
+Note that $T^{1}$ has the same distribution (up to a sign change) as $T^{-1}$ and
 $\mathbb{P}(E_{n}\geq l)\leq$ $\mathbb{P}(X_{[1,2n]}(n+1,T^{1}(\theta^{n}\omega))\geq l/2-1)+\mathbb{P}(X_{[1,2n]}(n,T^{-1}(\theta^{n-1}\omega))\geq l/2-1)$
@@
 In the equation above, we used the notation of [4], where $\omega$ is a variable in the state space that represents the sequence of steps taken by the random walk.
-$\theta^{n}\omega$ is the translation of the walk by $n$ time steps; namely, the new walk $\hat{S}$ given by $\theta^{n}\omega$ satisfies $\hat{S}_{i}=S_{i-n}$.
-We now fix a parameter $J$, whose value will be set sufficiently large later. We then define $L=n/(3J)$ and set $k_{j}:=Lj$ for $0\leq j\leq 3J$.
-If we let
+$\theta^{n}\omega$ is the translation of the walk by $n$ time steps; namely, the new walk $\hat{S}$ given by $\theta^{n}\omega$ satisfies $\hat{S}_{i}=S_{i-l}$.
+We now fix a parameter $J$, whose value will be set sufficiently large later. We then define $L=l/(3J)$ and set $k_{j}:=Lj$ for $0\leq j\leq 3J$. If we let
 $a_{1}:=\sum_{0\leq j\leq 3J-1}\mathbb{P}(S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)\neq\emptyset),$
-$a_{2}:=\mathbb{P}(\cap_{j=0}^{3J-1}\{N([1,n])=0,S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)=\emptyset\}),$
+$a_{2}:=\mathbb{P}(\cap_{j=0}^{3J-1}\{N([1,l])=0,S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)=\emptyset\}),$
 we have
-(2.3) $\mathbb{P}(T^{1}>n)\leq a_{1}+a_{2}.$
+(2.3) $\mathbb{P}(T^{1}>l)\leq a_{1}+a_{2}.$
 By the same proof in [4, Lemma 1.1] (which essentially computes the probability of having a large loop), we can obtain
-(2.4) $a_{1}\leq C(J)n^{-\frac{d}{2}+2}.$
-Note that if $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$, $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$ and $N([1,n])=0$, then $N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0$.
-To see the reason for this, note that $N([1,n])=0$ means that no point in the interval $k_{3j+1},k_{3j+2}-1]$ can be a cut point; thus, for $t\in[k_{3j+1},k_{3j+2}-1]$, we must have that $S(-\infty,t-1]\cap S[t,\infty)\neq\emptyset$.
+(2.4) $a_{1}\leq C(J)l^{-\frac{d}{2}+2}.$
+Note that if $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$, $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$ and $N([1,l])=0$, then $N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0$.
+To see the reason for this, note that $N([1,l])=0$ means that no point in the interval $k_{3j+1},k_{3j+2}-1]$ can be a cut point; thus, for $t\in[k_{3j+1},k_{3j+2}-1]$, we must have that $S(-\infty,t-1]\cap S[t,\infty)\neq\emptyset$.
 However, since $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$ and $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$, $S[k_{3j},t-1]\cap S[t,k_{3j+3}-1]\neq\emptyset$ holds instead.
@@
 For $0\leq j<J$ and $0<\alpha<1$, we have:
-$A_{1}^{j,\alpha}=$ $\Big\{S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+n^{\alpha},k_{3j+3}-1]\neq\emptyset\Big\}$
-$\cup\Big\{S[k_{3j},\,k_{3j+2}-n^{\alpha}]\cap S[k_{3j+2},\,k_{3j+3}-1]\neq\emptyset\Big\}.$
+$A_{1}^{j,\alpha}=$ $\Big\{S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+l^{\alpha},k_{3j+3}-1]\neq\emptyset\Big\}$
+$\cup\Big\{S[k_{3j},\,k_{3j+2}-l^{\alpha}]\cap S[k_{3j+2},\,k_{3j+3}-1]\neq\emptyset\Big\}.$
 Then,
@@
 $\leq\prod_{j=1}^{J}\mathbb{P}(A_{1}^{j,\alpha}\cup((A_{1}^{j,\alpha})^{c}\cap\{N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\})).$
-Note that
-$\mathbb{P}(A_{1}^{j,\alpha})\leq C(n^{\alpha})^{-\frac{d}{2}+2}.$
+By the same proof as that of (2.4),
+$\mathbb{P}(A_{1}^{j,\alpha})\leq C(l^{\alpha})^{-\frac{d}{2}+2}.$
 Now, we can split $k_{3j+1},k_{3j+2}-1]$” like in (2.3).
-That is, let $L^{\prime}=n^{\alpha}$, $J^{\prime}=\frac{k_{3j+2}-k_{3j+1}}{2L^{\prime}}$, $k^{\prime}_{\ell}=\ell L^{\prime}$ for $1\leq\ell\leq 2J^{\prime}$.
+That is, let $L^{\prime}=l^{\alpha}$, $J^{\prime}=\frac{k_{3j+2}-k_{3j+1}}{2L^{\prime}}$, $k^{\prime}_{u}=uL^{\prime}$ for $1\leq u\leq 2J^{\prime}$.
 Then, we have
 $\mathbb{P}\Big(N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0,(A_{1}^{j,\alpha})^{c}\Big)$
-$\quad\leq\bigg\{\sum_{\ell=1}^{2J-1}\mathbb{P}\Big(S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell}-1]\cap S[k_{3j+1}+k^{\prime}_{\ell+1},\,k_{3j+2}-1]\neq\emptyset\Big)\bigg\}(:=\hat{a}_{1})$
-$\qquad+\mathbb{P}\Big(\text{for all }1\leq\ell\leq 2J^{\prime},\,S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+2}-1]=\emptyset,\,$
+$\quad\leq\bigg\{\sum_{u=1}^{2J-1}\mathbb{P}\Big(S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{u}-1]\cap S[k_{3j+1}+k^{\prime}_{u+1},\,k_{3j+2}-1]\neq\emptyset\Big)\bigg\}(:=\hat{a}_{1})$
+$\qquad+\mathbb{P}\Big(\text{for all }1\leq u\leq 2J^{\prime},\,S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{u-1}]\cap S[k_{3j+1}+k^{\prime}_{u},\,k_{3j+2}-1]=\emptyset,\,$
 $\quad\quad N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0,$
@@
 Note that
-$\hat{a}_{2}\leq$ $\mathbb{P}\Big(\text{for all }0\leq\ell\leq 2J^{\prime}+1,\,S(k_{3j},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+3})=\emptyset,\,$
+$\hat{a}_{2}\leq$ $\mathbb{P}\Big(\text{for all }0\leq u\leq 2J^{\prime}+1,\,S(k_{3j},\,k_{3j+1}+k^{\prime}_{u-1}]\cap S[k_{3j+1}+k^{\prime}_{u},\,k_{3j+3})=\emptyset,\,$
 $\quad\quad N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0\Big).$
 By the same proof as that of [4, Lemma 1.1], we have
-$\hat{a}_{1}\leq C\,n^{1-\alpha}\,(n^{\alpha})^{-\frac{d}{2}+2},\quad\hat{a}_{2}\leq C\,n^{-\alpha}.$
-Then if we choose $\alpha$ close to $1$, for some $\delta>0$ (it should not depend on $J$),
-$\hat{a}_{1}+\hat{a}_{2}\ll n^{-\delta}$
+$\hat{a}_{1}\leq C\,l^{1-\alpha}\,(n^{\alpha})^{-\frac{d}{2}+2},\quad\hat{a}_{2}\leq C\,l^{-\alpha}.$
+Then, if we choose $\alpha$ close to $1$, for some $\delta>0$ (note that the obtained value of $\delta$ does not depend on $J$),
+$\hat{a}_{1}+\hat{a}_{2}\ll l^{-\delta}$
 and hence
-$a_{2}\leq C(J)(n^{-\delta})^{J}.$
-Therefore, if we pick sufficiently large $J$, we obtain $a_{2}\leq Cn^{-\frac{d}{2}+2}$ and in addition the desired result follows by combining (2.3) and (2.4). ∎
+$a_{2}\leq C(J)(l^{-\delta})^{J}.$
+Therefore, if we pick sufficiently large $J$, we obtain $a_{2}\leq Cl^{-\frac{d}{2}+2}$ and, in addition, the desired result, equation (2.1), follows by combining (2.3) and (2.4). ∎
 Proof of (2.2).
 We start by defining a collection of sets which, when they hold together, will imply $E_{n}>cn$ for an appropriate constant $c$. Let
-$D_{1}:=\{\ X[n,2n]\geq c_{d}n\},$ $D_{2}:=\{{\rm Cap}(S[2n-sn,2n])\geq c^{\prime}_{d}sn\},$
-$D_{3}:=\{\max_{2n-sn\leq i\leq 2n}\|S_{i}-S_{n}\|\leq C_{1}n^{1/2}\},$ $D_{4}:=\{\|S_{sn}-S_{n}\|\leq C_{1}n^{1/2}\}.$
-Here, $c_{d}$ and $c_{d}^{\prime}$ are some dimension-dependent constants, (by the law of large numbers for the capacity of the random walk range, we choose $c_{d}^{\prime}$ to satisfy ${\rm Cap}(S[0,n])\sim 2c^{\prime}_{d}n$).
+$D_{1}:=\{\ X[n,2n]\geq c_{d}n/2\},$ $D_{2}:=\{{\rm Cap}(S[2n-sn,2n])\geq c^{\prime}_{d}sn\},$
+$D_{3}:=\{\max_{2n-sn\leq i\leq 2n}\|S_{i}-S_{n}\|\leq C_{1}n^{1/2}\},$ $D_{4}:=\{\|S_{sn}-S_{n}\|\leq C_{1}n^{1/2}\},$
+(see the definition of the capacity in [12] et al.).
+Here, $c_{d}$ and $c_{d}^{\prime}$ are some dimension-dependent constants (by the law of large numbers for the capacity of the random walk range, we choose $c_{d}^{\prime}$ to satisfy ${\rm Cap}(S[0,n])\sim 2c^{\prime}_{d}n$).
+In addition, we choose $0<s<c_{d}/8$.
 Recall the (slightly modified) definition of $T^{1}$ as the first cut point after time $n$.
@@
 $\mathbb{P}(D_{1})\geq$ $\mathbb{P}(T^{1}\leq n+f(n),X^{2}_{(-\infty,\infty)}[T^{1},2n]\geq c_{d}n/2)-\mathbb{P}(T^{1}\geq n+f(n))$
-$\geq$ $\mathbb{P}(X^{2}_{(-\infty,\infty)}(0,n-f(n))\geq c_{d}n/2|0\text{ is cut point})-\mathbb{P}(T^{1}\geq n+f(n))=1-o(1)$
+$\geq$ $\mathbb{P}(X^{2}_{(-\infty,\infty)}[0,n-f(n)]\geq c_{d}n/2|0\text{ is cut point})-\mathbb{P}(T^{1}\geq n+f(n))=1-o(1)$
 if $f(n)=o(n)$ and $f(n)\to\infty$ as $n\to\infty$ with the aid of the estimate in the proof of (2.1).
@@
 Let us define the set A:=S[2nsn,2n]A:=S[2n-sn,2n] and $z:=S_{sn}$.
-On the event $D_{1}\cap D_{2}$, we see that we have $\max_{x\in A}\|x-z\|\leq 2C_{1}n^{1/2}$.
+On the event $D_{3}\cap D_{4}$, we see that we have $\max_{x\in A}\|x-z\|\leq 2C_{1}n^{1/2}$.
 We now reverse the random walk $S_{0},\ldots,S_{sn}$, to create a new random walk $\tilde{S}_{0},\ldots,\tilde{S}_{sn}$, defined as,
@@
 $\mathbb{P}^{z}(T_{A}<sn)$ $=\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}>sn-i)$
-$\geq\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}=\infty)$
-$\geq\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a)\sum_{x\in A}\mathbb{P}^{x}(T_{A}=\infty)$
+$\geq\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}=\infty)\geq\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a)\sum_{x\in A}\mathbb{P}^{x}(T_{A}=\infty)$
 $=\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a){\rm Cap}(A)\geq c(s)n^{-d/2+1}{\rm Cap}(A).$
@@
 [15, Theorem 1.2.1]).
-On the event $D_{2}$, we assumed that the capacity of $A$ is greater than $sn$.
+On the event $D_{2}$, we assumed that the capacity of $A$ is greater than $c_{d}^{\prime}sn$.
 Thus, conditionally on the intersection of the events $D_{2}$, $D_{3}$, and $D_{4}$, we can derive a lower bound on the probability of intersection between $\tilde{S}$ (which is also $S[1,sn]$) and $A$ (which is S[2nsn,2n]S[2n-sn,2n]) as
@@
 We remark now that on this intersection event, we can derive the following lower bound on the cross term $E_{n}$. First, notice that,
-$X[n,2n]\geq c_{d}n,$
-$X_{n}\geq 0,$
-$X_{2n}\leq 2sn.$
-The first line comes from the event $D_{1}$ and the second line is trivial.
-To see the last line (for the graph distance), notice that if we let $S_{t_{1}}$ and $S_{t_{2}}$ be the time indices of the intersection $\{S_{1},\ldots,S_{sn}\}$ with $\{S_{2n-sn},\ldots,S_{2n}\}$ with $t_{1}\in[0,sn]$ and $t_{2}\in[2n-sn,2n]$, the shortest geodesic must be smaller than $t_{1}+(2n-t_{2})$.
-Namely, shortest geodesic must be shorter than the path $S_{0},\ldots,S_{t_{1}},S_{t_{2}+1},\ldots,S_{2n}$.
+$X[n,2n]\geq c_{d}n/2,\quad X_{n}\geq 0,\quad X_{2n}\leq 2sn.$
+The first line comes from the event $D_{1}$, and the second line is trivial.
+To see the last line (for the graph distance), notice that if we let $S_{t_{1}}$ and $S_{t_{2}}$ be the points of the intersection $\{S_{1},\ldots,S_{sn}\}$ and $\{S_{2n-sn},\ldots,S_{2n}\}$ with $t_{1}\in[0,sn]$ and $t_{2}\in[2n-sn,2n]$, the shortest geodesic must be smaller than $t_{1}+(2n-t_{2})$.
+Namely, the shortest geodesic must be shorter than the path $S_{0},\ldots,S_{t_{1}},S_{t_{2}+1},\ldots,S_{2n}$.
 Now, $t_{1}$ and $2n-t_{2}$ are both less than $sn$.
-The same argument applies for the number of cut points.
-To see how the proof holds for the effective resistance, notice that since the effective resistance decreases upon adding edges, we must have that $X_{2n}$ is less than the effective resistance of the graph consisting only of the points $\{S_{1},\ldots,S_{t_{1}}\}\cup\{S_{t_{2}},\ldots,S_{2n}\}$.
-The effective resistance of this graph is less than $t_{1}+(2n-t_{2})$, the number of points along this graph.
-Thus, we have that $E_{n}=X_{n}+X[n+1,2n]-X_{2n}\geq(c_{d}-2s)n$. Therefore,
-$\mathbb{P}(E_{n}\geq 2(c_{d}-s)n)\geq\mathbb{P}(\{S[1,sn]\cap S[2n-sn,2n]\neq\emptyset\}\cap(\cap_{i=1}^{4}A_{i}))\geq c(s)n^{-d/2+2}.$
+The same argument applies to the number of cut points.
+Thus, we have that $E_{n}=X_{n}+X[n+1,2n]-X_{2n}\geq(c_{d}/2-2s)n$. Therefore,
+$\mathbb{P}(E_{n}\geq 2(c_{d}/2-s)n)\geq\mathbb{P}(\{S[1,sn]\cap S[2n-sn,2n]\neq\emptyset\}\cap(\cap_{i=1}^{4}A_{i}))\geq c(s)n^{-d/2+2}.$
 Remark 2.3.
@@
 Lemma 2.4 ($d=6$).
-We have the following estimates for the error terms $E_{n}$.
+We have the following estimates for the error terms $E_{n}$:
 $\mathbb{E}[E_{n}]\lesssim(\log n),\quad\mathbb{E}[E_{n}^{2}]\asymp n.$
-As a consequence of the first and last inequality above, we have that $\mathrm{Var}(E_{n})\gtrsim n$.
-Let $\Omega_{n}$ be the event that $\mathbb{E}[E_{n}]-E_{n}\geq 0$, then, we have that
+As a consequence of the first and last inequality above, $\mathrm{Var}(E_{n})\gtrsim n$.
+Let $\Omega_{n}$ be the event that $\mathbb{E}[E_{n}]-E_{n}\geq 0$, then,
 (2.5) $\mathbb{E}[(E_{n}-\mathbb{E}[E_{n}])^{2}\mathbbm{1}[\Omega_{n}]]\lesssim(\log n)^{2}\ll n.$
@@
 $\mathbb{E}[E_{n}]=\sum_{l=1}^{n}\mathbb{P}(E_{n}\geq l)\lesssim\sum_{l=1}^{n}\frac{1}{l}\lesssim(\log n).$
-In addition, we have
+In addition, we obtain
 $\mathbb{E}[E_{n}^{2}]=\sum_{l=1}^{n}l\mathbb{P}(E_{n}\geq l)\lesssim\sum_{l=1}^{n}\frac{1}{l}l\lesssim\sum_{l=1}^{n}1=n.$
@@
 $\mathbb{E}[E_{n}^{2}]=\sum_{l=1}^{n}l\mathbb{P}(E_{n}\geq l)\gtrsim\sum_{l=an}^{bn}l\frac{1}{l}\gtrsim n.$
-Since $E_{n}\geq 0$, we have that $\mathbb{E}[E_{n}]-E_{n}\lesssim\log n$. This gives us the final estimate we desire. ∎
+Since $E_{n}\geq 0$, we obtain that $\mathbb{E}[E_{n}]-E_{n}\lesssim\log n$. This gives us the final estimate we desire. ∎
 3. Proof of CLT for $d\geq 7$
@@
 3.1. Proof of variance
-In this subsection, we will give the estimate of the variance.
+In this subsection, we will give an estimate for the variance.
 Lemma 3.1.
@@
 $\mathrm{Var}(X^{(i)}_{n})\sim\alpha_{d}^{i}n$
-with $\alpha_{d}^{i}>0$ for $i=1,2,3$.
+with $\alpha_{d}^{i}>0$ for $i=1,2$.
 To show it, we first prove the following lemma.
@@
 $\|\overline{X}_{n}+\overline{X}[n+1,2n]\|_{2}=\left(\|\overline{X}_{n}\|_{2}^{2}+\|\overline{X}[n,2n]\|_{2}^{2}\right)^{1/2}.$
-Using (2.1) and that $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq C(\log n)^{M}$ for $d\geq 8$ and $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq n^{1/4}$ for $d=7$, we find
+Using (2.1) and that $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq C(\log n)^{M}$ with $M<\infty$ for $d\geq 8$ and $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq n^{1/4}$ for $d=7$, we find
 $\|\overline{X}_{2n}\|_{2}\leq\left(\|\overline{X}_{n}\|_{2}^{2}+\|\overline{X}[n,2n]\|_{2}^{2}\right)^{1/2}+c_{1}n^{1/4}.$
@@
 $N_{2}$ $:=|\{i\in[1,n]:i-1,i+2\not\in D_{n},i,i+1\in D_{n}\}|.$
-Namely, $N_{1}$ denotes the number of events in which $Y_{i}=1$ occurs exactly once and $N_{2}$ the number of events in which $Y_{i}=1$ appears exactly twice in a row.
+Namely, $N_{1}$ denotes the number of events in which $Y_{i}=1$ occurs exactly once, and $N_{2}$ denotes the number of events in which $Y_{i}=1$ appears exactly twice in a row.
 $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge n})_{l=1}^{\infty}$.
-For large $M$, there exists $c>0$ such that
+Now, we state our desired conclusions. For large $M$, there exists $c>0$ such that
 $A_{n}:=\{|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\quad|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\},$
@@
 […]
-Eventually, for $|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$, $|b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$,
+As a consequence, for $|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$, $|b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$,
 (3.3) $\mathbb{P}^{n}(N_{1}=a,N_{2}=b+\sqrt{n})\leq C\mathbb{P}^{n-\sqrt{n}}(N_{1}=a+\sqrt{n},N_{2}=b).$
@@
 Lemma 3.4.
-It holds that $\alpha_{d}>0$.
+It holds that $\alpha_{d}>0$ for $d\geq 7$.
 Proof.
@@
 $N_{2}$ $:=|\{i\in[1,n/2]:i-1,i+2\not\in D,i,i+1\in D\}|,$
-and $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge(n/2)})_{l=1}^{\infty}$. By (3.1), there exists $c_{1}>0$ such that
+and $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge(n/2)})_{l=1}^{\infty}$.
+$N_{1}$ counts the number of backtracks, while $N_{2}$ counts the number of double backtracks.
+We remark that changing a single backtrack to a double backtrack will not change the graph distance of the walk.
+However, the act of changing all single backtracks to double backtracks and vice versa would substantially change the length of the random walk.
+The existence of a procedure to change the volume without changing the graph distance essentially allows us to argue that the variance of the graph distance must be non-trivial.
+By (3.1), there exists $c_{1}>0$ such that
 $A_{n}:=\{|N_{1}-\mathbb{E}[N_{1}]|\leq M\sqrt{n},\quad|N_{2}-\mathbb{E}[N_{2}]|\leq M\sqrt{n}\},$
@@
 Since $\mathbb{E}X_{n}/n\to c_{0}$ by Remark 2.3, we have $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}\geq c_{3}$ for some $c_{3}>0$.
-Indeed, if $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}<c_{3}$ and set $a_{m}=\mathbb{E}X_{m^{2}}$, then $(a_{m}-a_{m-1})/m\leq c_{3}$ for large $m$ and hence $a_{m}\leq c_{0}m^{2}/2$ for large $m$ if we pick small $c_{3}$.
+Indeed, if $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}<c_{3}$ and we set $a_{m}=\mathbb{E}X_{m^{2}}$, then $(a_{m}-a_{m-1})/m\leq c_{3}$ for large $m$; hence, $a_{m}\leq c_{0}m^{2}/2$ for large $m$ if we pick $c_{3}$ were chosen small enough.
 We now pick a subsequence $n$ such that $(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})\geq c_{3}\sqrt{n}$ and choose $\epsilon<c_{3}/2$.
@@
 Lemma 3.5.
-For $d\geq 7$ and for all $n\in\mathbb{N}$, we have
+For $d\geq 7$ and $n\in\mathbb{N}$, we have
 $\mathbb{E}[(\overline{X}_{n})^{2+\epsilon}]\lesssim n^{1+\epsilon/2}.$
@@
 Fix $k\geq 2$, and let $n$ satisfy $2^{k}\leq n<2^{k+1}$. Then, by the triangle inequality,
-$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}+\|E_{n}\|_{2+\epsilon},$
+$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}+\|E_{n}\|_{2+\epsilon}.$
 Using the inequality
@@
 as $n\to\infty$. ∎
+Remark 3.6.
+We remark here that all the proofs in this section will hold without change as long as we assume the tail statistic $\mathbb{P}(X_{n}+X_{[n,2n]}-X_{2n}\geq l)\leq Cl^{-d/2+2}$, for $d\geq 7$.
+This establishes Proposition 1.4 and Theorem 1.5 for $d\geq 7$.
 4. 6 Dimensions
@@
 $X_{n}=\sum_{k=0}^{\log_{2}n}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{k=0}^{n}X[k,k+1].$
-We define the error term
-$\mathcal{E}_{k}:=\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}.$
+We define
+$\mathcal{E}_{k}:=\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)},\quad\mathcal{E}:=\sum_{k=0}^{\log_{2}n}\mathcal{E}_{k},\quad MT:=\sum_{k=0}^{n}X[k,k+1].$
 By independence of the components $E_{n}^{(k,l)}$, we have that,
-$\mathrm{Var}(\mathcal{E}_{k})=\sum_{l=1}^{2^{k}}\mathrm{Var}(E_{n}^{(k,l)})\lesssim\sum_{l=1}^{2^{k}}\frac{n}{2^{k}}=n.$
-Thus we have, by expanding the variance and applying the Cauchy-Schwarz inequality, that
+(4.1) $\mathrm{Var}(\mathcal{E}_{k})=\sum_{l=1}^{2^{k}}\mathrm{Var}(E_{n}^{(k,l)})\lesssim\sum_{l=1}^{2^{k}}\frac{n}{2^{k}}=n.$
+By expanding the variance and applying the Cauchy-Schwarz inequality, we obtain
 (4.2) $\mathrm{Var}\left(\sum_{k=0}^{\log_{2}n}\mathcal{E}_{k}\right)\leq\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}n}\sum_{k_{2}=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k_{1}})^{1/2}\mathrm{Var}(\mathcal{E}_{k_{2}})^{1/2}\lesssim n(\log n)^{2}+n.$
-In addition, since each $X[k,k+1]$ is deterministically 1, we have that $\mathrm{Var}(X[k,k+1])=0$. Then, we have
+In addition, since each $X[k,k+1]$ is deterministically 1, we have that $\mathrm{Var}(X[k,k+1])=0$. Then,
 $\mathrm{Var}(X_{n})\lesssim n(\log n)^{2}.$
@@
 $X_{n}=\sum_{l=1}^{a_{n}}X[(l-1)\frac{n}{a_{n}},l\frac{n}{a_{n}}]+\sum_{k=0}^{\log_{2}(a_{n})}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}).$
+Furthermore, we remark that if the intervals $n/2^{k_{1}+1}(2l_{1}-2),n/2^{k_{1}+1}(2l_{1})]$ (, which is the interval corresponding to the cross term $E^{(k_{1},l_{1})}$) and $n/2^{k_{2}+1}(2l_{2}-2),n/2^{k_{2}+1}(2l_{2})]$(this is the interval corresponding to $E^{(k_{2},l_{2})}$) are completely disjoint, then $E^{(k_{1},l_{1})}$ and $E^{(k_{2},l_{2})}$ are independent and the covariance is 0.
+Thus, the only non-zero covariances $\text{Cov}(E^{(k_{1},l_{1})},E^{(k_{2},l_{2})})$ are those with $k_{2}\geq k_{1}$ and $2^{k_{2}-k_{1}}l_{1}\leq l_{2}\leq 2^{k_{2}-k_{1}}(l_{1}+1)$.
 Now, we claim that
-$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim(\log n)^{2}.$
-Let $\tilde{S}$ be the reversal of the first part, that is $\tilde{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}-i}-S_{l\frac{n}{2^{k}}}$ .
+$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim(\log n)^{2},$
+for most pairs of $(k_{1},l_{1})$ and $(k_{2},l_{2})$ with non-trivial intersection.
+Let $\tilde{S}$ be the reverse of the first part, that is $\tilde{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}-i}-S_{l\frac{n}{2^{k}}}$ .
 $\hat{S}$ will be the second part, that is $\hat{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}+i}-S_{l\frac{n}{2^{k}}}$.
@@
 $\mathbb{P}(A_{1}^{(a,k_{1},l_{1})}\cap A_{1}^{(b,k_{2},l_{2})})$
-$\leq$ $\mathbb{P}\bigg(\{\hat{S}^{(k_{1},l_{1})}[\frac{a}{2},\frac{n}{2^{k_{1}}}]\cap\tilde{S}^{(k_{1},l_{1})}[0,\frac{n}{2^{k_{1}}}]\neq\emptyset\}\cap\{\hat{S}^{(k_{2},l_{2})}[\frac{b}{2},\frac{n}{2^{k_{2}}}]\cap\tilde{S}^{(k_{2},l_{2})}[0,\frac{n}{2^{k_{2}}}]\neq\emptyset\}\bigg)$
+$\leq$ $\mathbb{P}\bigg(\{\hat{S}^{(k_{1},l_{1})}\left[\frac{a}{2},\frac{n}{2^{k_{1}}}\right]\cap\tilde{S}^{(k_{1},l_{1})}\left[0,\frac{n}{2^{k_{1}}}\right]\neq\emptyset\}\cap\{\hat{S}^{(k_{2},l_{2})}\left[\frac{b}{2},\frac{n}{2^{k_{2}}}\right]\cap\tilde{S}^{(k_{2},l_{2})}\left[0,\frac{n}{2^{k_{2}}}\right]\neq\emptyset\}\bigg)$
 $\leq$ $\sum_{u_{1}=\frac{a}{2}+l_{1}\frac{n}{2^{k_{1}}}}^{(l_{1}+1)\frac{n}{2^{k_{1}}}}\sum_{u_{2}=(l_{1}-1)\frac{n}{2^{k_{1}}}}^{l_{1}\frac{n}{2^{k_{1}}}}\sum_{u_{3}=\frac{b}{2}+l_{2}\frac{n}{2^{k_{2}}}}^{(l_{2}+1)\frac{n}{2^{k_{2}}}}\sum_{u_{4}=(l_{2}-1)\frac{n}{2^{k_{2}}}}^{l_{2}\frac{n}{2^{k_{2}}}}\sum_{x}\sum_{y}$
@@
 We can obtain the same result for $\mathbb{P}(A_{i}^{(a,k_{1},l_{1})}\cap A_{j}^{(b,k_{2},l_{2})})$ by a similar computation for other cases. Hence,
-$\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]$
+$\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]$
 $\leq$ $\sum_{k_{1}=0}^{\log_{2}(a_{n})}2^{k_{1}}[a_{n}/2^{k_{1}}](\log n)^{2}\leq a_{n}(\log n)^{3}\ll n(\log n).$
-Therefore,
-$\mathrm{Var}\left(\sum_{k=0}^{a_{n}}\mathcal{E}_{k}\right)\leq\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim n(\log n)$
+Therefore, with the aid of (4.1), we obtain,
+(4.3) $\mathrm{Var}\left(\sum_{k=0}^{a_{n}}\mathcal{E}_{k}\right)$
+$\leq$ $\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}\text{Cov}(E_{n}^{(k_{1},l_{1})},E_{n}^{(k_{2},l_{2})})$
+$\leq$ $\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim n(\log n)$
 and
@@
 Proof.
-Fix $\delta>0$ very close to 1. We will see where this condition plays a role close to the end of the proof.
-By our binary decomposition, we find that,
+Fix $\delta>0$ very close to 1. We will see where this condition comes into play near the end of the proof. By our binary decomposition, we find that,
 (4.4) $X_{n}=\sum_{l=1}^{n^{\delta}}X[(l-1)n^{1-\delta},ln^{1-\delta}]+\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}(-E_{n}^{(k,l)}).$
@@
 and
-$\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
+(4.5) $\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
 In addition, we obtain that, for sufficiently large $n$,
@@
 $\geq$ $-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\geq-(\log n)^{2}.$
-Furthermore, we remark that if the intervals $n/2^{k_{1}+1}(2l_{1}-2),n/2^{k_{1}+1}(2l_{1})]$ (this is the interval corresponding to the cross term $E^{(k_{1},l_{1})}$) and $n/2^{k_{2}+1}(2l_{2}-2),n/2^{k_{2}+1}(2l_{2})]$(this is the interval corresponding to $E^{(k_{2},l_{2})}$) are completely disjoint from each other, then $E^{(k_{1},l_{1})}$ and $E^{(k_{2},l_{2})}$ are independent of each other and the covariance is 0.
-Using these considerations, we see that a lower bound for the covariances is the following:
-$\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(n^{\delta})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}-(\log n)^{2}\geq-\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}2^{k_{1}}[n^{\delta}/2^{k_{1}}](\log n)^{2}$
+$\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(n^{\delta})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}-(\log n)^{2}\geq-\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}2^{k_{1}}[n^{\delta}/2^{k_{1}}](\log n)^{2}$
 $=-n^{\delta}(\log n)^{3}\gg-n(\log n),\text{ if }\delta<1.$
+By combining these estimates, we have,
+(4.6) $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n.$
 Furthermore, notice that by Lemma 4.1, we have the upper bound $\mathrm{Var}(X_{n})\leq Cn\log n$. Thus, we can show that
 $\mathrm{Var}(MT)\leq Cn^{\delta}[n^{1-\delta}\log(n^{1-\delta})]\leq C(1-\delta)n\log n,$
-by setting $\delta$ sufficiently close to $1$, this can be made very small.
+setting $\delta$ sufficiently close to $1$, this can be made very small.
 Finally, we can apply the Cauchy-Schwarz inequality to say that,
 $\mathrm{Var}(X_{n})\geq\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})-2\mathrm{Var}(MT)^{1/2}\mathrm{Var}(\mathcal{E})^{1/2}.$
-Then for $\delta$ close to but not equal to $1$, we have $\mathrm{Var}(\mathcal{E})\geq cn\log n$ and $\mathrm{Var}(MT)\leq C(1-\delta)n\log n$.
+Then for $\delta$ close to but not equal to $1$, we have $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n$ and $\mathrm{Var}(MT)\leq C(1-\delta)n\log n$.
 Thus, $\mathrm{Var}(X_{n})\gtrsim n\log n$, as desired.
 ∎
-4.1.2. Effective resistance
+4.1.2. General Analysis with heavy tailed Cross Terms
+In this section, we will just assume that we are considering a collection of variables $X_{[a,b]}$ that are measurable with respect to the walk $\{S_{a},\ldots,S_{b}\}$ and that the cross term $E_{n}:=X_{n}+X_{[0,n]}-X_{2n}$ satisfies the tail bounds $\mathbb{P}(E_{n}\geq l)\asymp\frac{1}{l}$.
 We start with the following lower bound.
@@
 $\mathrm{Var}(X_{n})\leq\kappa n\log n.$
-Fix $\delta>0$.
-For $\delta>0$, we will choose a sufficiently small $\kappa>0$ and derive a contradiction later.
-Recall our binary decomposition in (4.2).
+Fix $\delta>0$. For $\delta>0$, we will choose a sufficiently small $\kappa>0$ and derive a contradiction later. Recall our binary decomposition in (4.4).
 As long as $n^{1-\delta}\geq N_{\kappa}$, we have that
-(4.3) $\mathrm{Var}(MT)\leq n^{\delta}[\kappa n^{1-\delta}\log n^{1-\delta}]=\kappa(1-\delta)n\log n$
-and
-(4.4) $\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
-We have that, for sufficiently large $n$,
-$\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})\geq\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}c\frac{n}{2^{k}}\geq c\delta n\log n.$
-Notice that
-$\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})})=\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$
-$\geq-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\gtrsim-(\log n)^{2}.$
-Here, we only used that $E_{n}\geq 0$ to get the second inequality.
-Putting these estimates back into equation (4.4), we observe that
-(4.5) $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n-O((\log n)^{2}(n^{\delta}\log(n^{\delta}))^{2})$
-This last term will be greater than $\frac{c\delta}{2}n\log n$ for $n$ sufficiently large.
-Finally, we can apply the Cauchy-Schwarz inequality to say that,
+$\mathrm{Var}(MT)\leq n^{\delta}[\kappa n^{1-\delta}\log n^{1-\delta}]=\kappa(1-\delta)n\log n.$
+In addition, equation (4.6), will give us $\text{Var}(\mathcal{E})\geq c\delta n\log n.$ Finally, we can apply the Cauchy-Schwarz inequality to say that,
 (4.7) $\mathrm{Var}(X_{n})\geq\mathrm{Var}(MT)+\mathrm{Var}(\mathcal{E})-2\mathrm{Var}(MT)^{1/2}\mathrm{Var}(\mathcal{E})^{1/2}.$
-Notice that as a function of the value of $\mathrm{Var}(\mathcal{E})$, the right-hand side is increasing when $\mathrm{Var}(\mathcal{E})\geq\mathrm{Var}(MT).$ We have already obtained a lower bound on $\mathrm{Var}(\mathcal{E})$ from equation (4.5) that is larger than the upper bound we have on $\mathrm{Var}(MT)$ from equation (4.3).
-Indeed, notice that if $\mathrm{Var}(MT)<\frac{c\delta}{8}n\log n$, then we would have that the right hand side of (4.6) will be greater than,
+Notice that, as a function of the value of $\mathrm{Var}(\mathcal{E})$, the right-hand side increases when $\mathrm{Var}(\mathcal{E})\geq\mathrm{Var}(MT)$.
+Indeed, notice that if $\mathrm{Var}(MT)<\frac{c\delta}{8}n\log n$, then the right hand side of (4.7) will be greater than,
 $(\mathrm{Var}(\mathcal{E})^{1/2}-\mathrm{Var}(MT)^{1/2})^{2}\geq\left(\frac{\sqrt{c\delta}}{\sqrt{2}}\sqrt{n\log n}-\frac{\sqrt{c\delta}}{2\sqrt{2}}\sqrt{n\log n}\right)^{2}\geq\frac{c\delta}{8}n\log n.$
-Then, we have that by setting $\kappa$ sufficiently small that and applying the previous inequality that $\mathrm{Var}(X_{n})\geq\frac{c\delta}{8}n\log n$.
+Then, by setting $\kappa$ sufficiently small and applying the previous inequality, we have that $\mathrm{Var}(X_{n})\geq\frac{c\delta}{8}n\log n$.
 This is a contradiction to the assumption that we cannot produce a lower bound to the $\limsup$ of $\frac{\mathrm{Var}(X_{n})}{n\log n}$.
 ∎
-Now, we produce an upper bound to the variance of the effective resistance. Since the proof is the same as that of (4.1), we omit it.
+Now, we produce an upper bound to the variance of general variables with the same heavy tail on the cross terms. Since the proof is the same as that of (4.2), we omit it.
 Lemma 4.4.
@@
 $Z_{n_{\delta},i}:=\frac{X[(i-1)n/n_{\delta},in/n_{\delta}]-\mathbb{E}[X[(i-1)n/n_{\delta},in/n_{\delta}]]}{\sqrt{\mathrm{Var}(X_{n/n_{\delta}})}}.$
-The variables $Z_{n_{\delta},i}$ are all independent of each other and one
+The variables $Z_{n_{\delta},i}$ are all independent of each other and one obtains
 $\sum_{i=1}^{n_{\delta}}\mathbb{E}\left[Z_{n_{\delta},i}^{2}\right]=1.$
@@
 Thus, we see that $\mathbb{E}[\mathcal{M}^{2}]\leq 4n(\sqrt{\log n})^{2-2\delta}$. Hence, by Markov’s inequality, we have that
-(4.7) $\mathbb{P}(\mathcal{M}\geq\sqrt{n(\log n)^{1-\delta/2}})\to 0$
+(4.8) $\mathbb{P}\bigg(\mathcal{M}\geq\sqrt{n(\log n)^{1-\delta/2}}\bigg)\to 0$
 as $n\to\infty$ for any $\epsilon>0$.
@@
 as $n\to\infty$.
-Recall that $\mathcal{G}$ from Part 1 was a Gaussian random variable and the order of the variance of $\mathcal{G}$ was $n\log n$, which is much larger than the error terms of Parts 2 and 3.
-Then, combining Part 1, 2, 3, we can obtain the desired result using the Lindeberg–Feller central limit theorem.
-4.2.2. Effective resistance
-Next, we will give the proof of Theorem 1.4 for $d=6$.
+Recall that $\mathcal{G}$ from Part 1 was a Gaussian random variable, and the order of the variance of $\mathcal{G}$ was $n\log n$, which is much larger than the error terms of Parts 2 and 3.
+Then, combining Parts 1, 2, and 3, we can obtain the desired result using the Lindeberg–Feller central limit theorem.
+4.2.2. General Analysis with heavy tailed Cross Terms
+In this section, we will demonstrate a version of the phase transition demonstrated for the graph distance of the random walk range in six dimensions for variables whose cross terms have the same heavy tailed asymptotics: $\mathbb{P}(E_{n}\geq l)\asymp l^{-1}$.
+In particular, we do not make any assumptions about the covariance structure for the cross terms.
+We emphasize that the results of this section show that heavy tailed behavior is the key source of the phase transition.
+We note that our results will apply to the effective resistance; we briefly give the definitions here.
+We define a quadratic form $\mathcal{E}$ on functions $f,g:V(G)\to\mathbb{R}$ by
+$\mathcal{E}(f,g)=\frac{1}{2}\sum_{\begin{subarray}{c}x,y\in V\\ \{x,y\}\in E\end{subarray}}(f(x)-f(y))(g(x)-g(y)).$
+This corresponds to the energy dissipation of an electrical network on the graph $G$. We also introduce the space
+$H^{2}:=\{\,f\in\mathbb{R}^{V}:\mathcal{E}(f,f)<\infty\,\},$
+which serves as the natural domain of the energy form. The effective resistance between $A$ and $B$ is defined via the energy minimization
+$R_{G}(A,B)^{-1}\;=\;\inf\Biggl\{\mathcal{E}(f,f)\;\Big|\;\quad f|_{A}=1,\ f|_{B}=0,f\in H^{2}\Biggr\}.$
+This quantity represents the voltage difference required to send the current from $A$ to $B$, and therefore corresponds to the usual notion of effective resistance in an electrical network.
+For vertices $x,y\in V$, we write $R_{G}(x,y)$ to denote $R_{G}(\{x\},\{y\})$.
+Now, we will give the proof of Theorem 1.5 for $d=6$.
 We present only the necessary adjustments from the proof of Theorem 1.2.
 This involves changing Parts 1 and 4, while keeping Parts 2 and 3.
-We also remark that the results of this section only depend on the tail behavior of the cross terms $E_{n}$ behaving like $\mathbb{P}(E_{n}\geq l)\asymp\frac{1}{l}$.
-In particular, the results are general enough to not require any estimate on the correlation between the cross terms.
-Thus, the methods of this section are of general interest and can potentially be applied to other heavy tailed statistics of the random walk range.
+As before, we mention that the particular value of this result is that it demonstrates that the right order of heavy tails is what ultimately leads to a phase transition in dimension 6, rather than detailed estimates of the covariance structure.
 We now define the function $f(n)$ as,
@@
 $f(m_{k+1})-f(m_{k})\geq\frac{1}{\log\log m_{k+1}}\geq\frac{1}{(k+1)\log 4},$
-for all $k$ large enough. This will imply that $\lim_{k\to\infty}f(m_{k})=\infty$. This contradicts the upper bound on $f$ from Lemma 4.1.
+for all $k$ large enough. This will imply that $\lim_{k\to\infty}f(m_{k})=\infty$. This contradicts the upper bound of $f$ from Lemma 4.1.
 Thus, we see that this proof establishes that there is a subsequence where $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}},$ does not converge to 0.
-Moreover, along this subsequence $X_{n}-\mathbb{E}[X_{n}]$ will be distributed (after appropriate normalization) as a normal random variable whose variance is given by $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$.
+Moreover, along this subsequence, $X_{n}-\mathbb{E}[X_{n}]$ will be distributed (after appropriate normalization) as a normal random variable whose variance is given by $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$.
 ∎
 5. The fluctuations converge to 0 in dimension $d=5$
-In this section, we will give the proof of Propositions 1.1 and 1.3 and Theorems 1.2 and 1.4 in $d=5$.
-Since we do not show the correlation estimates (5.1) for the effective resistance, we present the weaker estimates of its variance.
-Since the proof is similar, we will omit the proof of Proposition 1.3 in $d=5$.
+In this section, we will give the proof of Propositions 1.1 and 1.4 and Theorems 1.2 and 1.5 in $d=5$. Since the proof is similar, we will omit the proof of Proposition 1.4 in $d=5$.
 Lemma 5.1.
-We have
+For $d=5$, we have
 $\mathrm{Var}(X_{n})\asymp n^{3/2}.$
@@
 For $0\leq k_{1},k_{2}\leq(1-\alpha)\log_{2}n$, $1\leq l_{1}\leq 2^{k_{1}}$, $1\leq l_{2}\leq 2^{k_{2}}$, let
-$\Lambda_{1}:=\{((k_{1},l_{1}),(k_{2},l_{2})):[2^{-k_{1}}(l_{1}-1)n,2^{-k_{1}}l_{1}n]\cap[2^{-k_{2}}l_{2}n,2^{-k_{2}}(l_{2}+1)n]\neq\emptyset\},$
+$\Lambda_{1}:=\{((k_{1},l_{1}),(k_{2},l_{2})):[2^{-k_{1}}(l_{1}-1)n,2^{-k_{1}}l_{1}n]\cap[2^{-k_{2}(l_{2}-1)n},2^{-k_{2}}l_{2}n]\neq\emptyset\},$
 $\Lambda_{2}:=\{((k_{1},l_{1}),(k_{2},l_{2})):[2^{-k_{1})}(l_{1}-1)n,2^{k_{1}}l_{1}n]\cap[2^{-k_{2}}(l_{2}-1)n,2^{-k_{2}}l_{2}n]=\emptyset\}.$
@@
 Hence, we obtain $\mathrm{Var}(\mathcal{E})\gtrsim n^{3/2}$.
-In addition, (2.2) yields $\sum_{k=1}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\leq cn^{3/2}$. Now by the same argument as that of Lemma 4.1, we have
+In addition, (2.2) yields $\sum_{k=1}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\leq cn^{3/2}$.
+Now by the same argument as that of Lemma 4.1, we have
 (5.1) $\sum_{k_{1}=0}^{(1-\alpha)\log_{2}n}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{(1-\alpha)\log_{2}n}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\leq n^{3/2-\alpha/2+o(1)}.$
@@
 ∎
-Proof of Theorems 1.2 and 1.4 in $d=5$.
-For small $\delta>0$, we use Markov’s inequality in the computation of the variance, to derive,
+Proof of Theorems 1.2 and 1.5 in $d=5$.
+For small $\delta>0$, we use Markov’s inequality in the computation of the variance to derive,
 $\sup_{\alpha\leq\beta\leq 1-\delta}\mathbb{P}\bigg(\frac{\sum_{l=1}^{n^{1-\beta}}(E_{n}^{((1-\beta)\log n,l)}-\mathbb{E}[E_{n}^{((1-\beta)\log n,l)}])}{\sqrt{\mathrm{Var}(X_{n})}}\geq\frac{\epsilon}{\log n}\bigg)$
@@
 $\lesssim$ $\sup_{\alpha\leq\beta\leq 1-\delta}(\log n)^{2}\frac{n^{1-\beta}n^{3\beta/2}}{n^{3/2}}\lesssim(\log n)^{2}n^{-\delta/2}.$
-Now, we use Markov’s inequality in the computation of the absolute value combined with the triangle inequality, to get,
+Now, we use Markov’s inequality, along with the triangle inequality, to get,
 $\sup_{1-\delta\leq\beta\leq 1}\mathbb{P}\bigg(\frac{\sum_{l=1}^{n^{1-\beta}}(E_{n}^{((1-\beta)\log n,l)}-\mathbb{E}[E_{n}^{((1-\beta)\log n,l)}])}{\sqrt{\mathrm{Var}(X_{n})}}\geq\frac{\epsilon}{\log n}\bigg)$
@@
 In this section, we estimate the variance and the limit in distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ in $d=4$.
-In what follows, we will simplify notation by reducing to a discussion of only the cross term $E_{n}$ since we will only apply a dyadic decomposition and the only other cross term that could appear, $X_{2n+1}-X_{n}-X[n,2n+1]$, will have a similar analysis.
-In subsection 6.1, we will give the lower bounds for the cross term, which holds for the graph distance, cut points and the effective resistance.
+In subsection 6.1, we will give the lower bounds for the cross term, which holds for the graph distance and the cut points.
 In subsections 6.2-6.4, we will show the upper bounds for the cross term and main results in $d=4$.
-In these subsections, we are considering only the graph distance, cut points since Lemma 6.8 does not work for the effective resistance.
-We will give our main result in $d=4$ for the effective resistance in subsection 6.5.
+In subsection 6.5, we will present our results for general random variables whose cross terms have the same heavy tailed behavior as the graph distance in dimension 4.
 6.1. Lower bounds for the Cross Term
 We will first start by obtaining a lower bound of $\mathbb{E}[E_{n}].$ Our heuristic argument is the following: if there is an intersection between $S[3n/2,2n]$ and $S[0,n]$, then this will form a shortcut in the graph distance between the points $S_{2n}$ and $S_{0}$.
-The amount that the graph distance would reduce via this shortcut will be related to the graph distance between $S_{n}$ and $S_{2n}$, which by a law of large numbers should be $O\left(\frac{n}{(\log n)^{1/2}}\right)$.
+The amount that the graph distance would reduce via this shortcut will be related to the graph distance between $S_{n}$ and $S_{2n}$, which, by a law of large numbers, should be $O\left(\frac{n}{(\log n)^{1/2}}\right)$.
 Furthermore, the probability of the intersection of $S[3n/2,2n]$ and $S[0,n]$ is approximately $O((\log n)^{-1})$.
@@
 However, the exact value of the reduction of the graph distance is a complicated function of the loop intersection structure of the two walks.
-Thus, knowing that there is a long-range intersection may not be enough the deduce the reduction of the graph distance.
-Furthermore, to get the exact order $\frac{n}{(\log n)^{3/2}}$, that we desire, then we would need to argue that there is exact independence between the event that there is an intersection between $S[3n/2,2n]$ and $S[0,n]$ and the event that the graph distance between $S_{n}$ and $S_{2n}$ is $O\left(\frac{n}{(\log n)^{1/2}}\right)$.
+Thus, knowing that there is a long-range intersection may not be enough to deduce the reduction of the graph distance.
+Furthermore, to get the exact order $\frac{n}{(\log n)^{3/2}}$, that we desire, we would need to argue that there is an exact independence between the event that there is an intersection between $S[3n/2,2n]$ and $S[0,n]$ and the event that the graph distance between $S_{n}$ and $S_{2n}$ is $O\left(\frac{n}{(\log n)^{1/2}}\right)$.
 The following lemma will address these issues by finding appropriate events that will allow us to constrain the geometry of the random walk range.
 Lemma 6.1.
-Let $E_{n}$ denote the cross term that appears when considering the number of cut points, the effective resistance or the graph distance. We have $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$.
+Let $E_{n}$ denote the cross term that appears when considering the number of cut points or the graph distance. We have $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$.
 Remark 6.2.
 In [21, Section 4], he has already shown $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$ for the effective resistance.
-We obtain the same result by a slightly simpler proof which derives intersection estimates by using the capacity of the random walk.
+We obtain the same result with a slightly simpler proof, which derives intersection estimates by using the capacity of the random walk.
 Proof.
@@
 $A_{4}=\{{\rm Cap}(\tilde{S}(n-n/(\log n)^{\epsilon},n))\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}\},$
-$A_{5}=\{\tilde{S}_{i}\leq C_{1}\sqrt{n},\quad\forall i\in[n-n/(\log n)^{\epsilon},n]\}$
-noting $\mathbb{E}[{\rm Cap}(\tilde{S}[0,n])]\sim\frac{\pi^{2}}{8}\frac{n}{\log n}$ (e.g., see [6]).
+$A_{5}=\{\tilde{S}_{i}\leq C_{1}\sqrt{n},\quad\forall i\in[n-n/(\log n)^{\epsilon},n]\}.$
 In the above equation, we use the notation $\tilde{X}[a,b]$ to denote the graph distance between the points $\tilde{S}_{a}$ and $\tilde{S}_{b}$ along the graph $\tilde{S}[a,b]$.
 By [11, Lemma 2.3], $\mathbb{P}(A_{1})=1-o(1)$.
-Similarly, the law of large numbers ensures that $\mathbb{P}(A_{2})$, $\mathbb{P}(A_{3})=1-o(1)$ noting [11, Theorem 1.2] and $\mathbb{P}(A_{4})=1-o(1)$ (see [6, Theorem 1.1]).
+Similarly, the law of large numbers ensures that $\mathbb{P}(A_{2})$, $\mathbb{P}(A_{3})=1-o(1)$ (noting [11, Theorem 1.2]) and $\mathbb{P}(A_{4})=1-o(1)$ (see [6, Theorem 1.1]).
 By the local central limit theorem, we have that $\mathbb{P}(A_{5})=1-\epsilon(C_{1})$, where $\epsilon(C_{1})\to 0$ as $C_{1}\to\infty$.
@@
 Now, we need to define an event on the intersection of $\hat{S}$ with $\tilde{S}$.
-Let $\tau$ be defined to be the first time of intersection between $\hat{S}$ and $\tilde{S}[n-n/(\log n)^{\epsilon},n]$ after time $n-n/(\log n)^{\epsilon}$.
+Let $\tau$ be defined as the first time of intersection between $\hat{S}$ and $\tilde{S}[n-n/(\log n)^{\epsilon},n]$ after time $n-n/(\log n)^{\epsilon}$.
 That is, $\tau:=\inf\{i>n-n/(\log n)^{\epsilon}:\hat{S}_{i}\in\tilde{S}[n-n/(\log n)^{\epsilon},n]\}$.
@@
 $=$ $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right){\rm Cap}(\mathcal{S}).$
-On the union $A_{5}\cap B_{2}$, we can apply the local central limit theorem to deduce that $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right)\geq cn^{-1}(\log n)^{-\epsilon}$, for some constant $c>0$ on $\mathcal{Q}$.
+On the union $A_{5}\cap B_{2}$, we can apply the local central limit theorem to deduce that $\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\geq cn^{-1}(\log n)^{-\epsilon}$, for some constant $c>0$ on $\mathcal{Q}$.
 Furthermore, the event $A_{5}$ ensures that ${\rm Cap}(\mathcal{S})\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}$ on $\mathcal{Q}$.
@@
 By multiplying these two terms together, we see that $\mathbb{E}[E_{n}]>n/(\log n)^{3/2+\epsilon/2+2\epsilon}$.
-This proof also works even for the effective resistance or the number of cut points.
+This proof also works even for the number of cut points.
 ∎
@@
 The third and fourth events force there to be no cut points in $\tilde{S}$ between $n/(\log n)^{4}$ and $n/(\log n)^{2}$.
-We claim both that $\Omega$ holds with probability $O\left(\frac{\log\log n}{\log n}\right)$ and that on the event $\Omega^{c}$, we must have that the cross term $E_{n}$ for number of cut points, the effective resistance, and the graph distance must by less than $\frac{n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$ by Lemma 6.1.
+We claim both that $\Omega$ holds with probability $O\left(\frac{\log\log n}{\log n}\right)$ and that on the event $\Omega^{c}$, we must have that the cross term $E_{n}$ for the number of cut points and the graph distance must be less than $\frac{n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$ by Lemma 6.1.
 By [15, Theorem 4.3.6 part(ii)], we must have that $\mathbb{P}(\hat{S}[n/(\log n)^{4},n]\cap\tilde{S}[0,n]\neq\emptyset)=\text{O}\left(\frac{\log\log n}{\log n}\right)$.
@@
 Notice that on $\Omega^{c}$, there must be some cut point in the interval $\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]$ for both $\tilde{S}$ and $\hat{S}$.
-Let $\mathcal{C}_{\tilde{S}}$ denote the time instance of the cut point of $\tilde{S}$ (cut time) in this interval and let $\mathcal{C}_{\hat{S}}$ denote the time instance of the cut point of $\hat{S}$ in this interval.
-Furthermore, since there is no intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$, $\mathcal{C}_{\hat{S}}$ must still be a cut point in the full graph $\hat{S}\cup\tilde{S}$.
-Notice that we have $\hat{X}_{n}=\hat{X}[0,\mathcal{C}_{\hat{S}}]+\hat{X}[\mathcal{C}_{\hat{S}},n]\leq\hat{X}[\mathcal{C}_{\hat{S}},n]+n/(\log n)^{2}.$ Similarly, we have $\tilde{X}_{n}\leq\tilde{X}[\mathcal{C}_{\tilde{S}},n]+n/(\log n)^{2}.$ Lastly, we see that $X_{2n}\geq\hat{X}[\mathcal{C}_{\hat{S}},n]+\tilde{X}[\mathcal{C}_{\tilde{S}},n]$.
-Altogether, we see that on $\Omega^{c}$, we have $E_{n}=\tilde{X}_{n}+\hat{X}_{n}-X_{2n}\leq\frac{2n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$.
+Let $\mathcal{C}_{\tilde{S}}$ denote the time instance of the cut point of $\tilde{S}$ (cut time) in this interval, and let $\mathcal{C}_{\hat{S}}$ denote the cut time of $\hat{S}$ in this interval.
+Furthermore, since there is no intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$, $\mathcal{C}_{\hat{S}}$ must still be a cut time in the full graph $\hat{S}\cup\tilde{S}$.
+Notice that we have $\hat{X}_{n}=\hat{X}[0,\mathcal{C}_{\hat{S}}]+\hat{X}[\mathcal{C}_{\hat{S}},n]\leq\hat{X}[\mathcal{C}_{\hat{S}},n]+n/(\log n)^{2}.$ Similarly, we have $\tilde{X}_{n}\leq\tilde{X}[\mathcal{C}_{\tilde{S}},n]+n/(\log n)^{2}.$ Lastly, $X_{2n}\geq\hat{X}[\mathcal{C}_{\hat{S}},n]+\tilde{X}[\mathcal{C}_{\tilde{S}},n]$.
+Altogether, on $\Omega^{c}$, $E_{n}=\tilde{X}_{n}+\hat{X}_{n}-X_{2n}\leq\frac{2n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$.
 ∎
@@
 Corollary 6.5.
-Let $E_{n}$ denote the cross term that appears when computing the number of cut points, the effective resistance, or the graph distance. Then, we have,
+Let $E_{n}$ denote the cross term that appears when computing the number of cut points or the graph distance. Then, we have,
 $\mathbb{E}[E_{n}^{2}]\geq\frac{n^{2}}{(\log n)^{2+o(1)}}.$
 Proof.
-Let $\Omega$ be the set from Lemma 6.3. Note that by the Cauchy-Schwarz inequality, we have that,
+Let $\Omega$ be the set from Lemma 6.3. Note that by the Cauchy-Schwarz inequality,
 $\mathbb{E}[E_{n}^{2}]\mathbb{P}(\Omega)\geq\mathbb{E}[E_{n}\mathbbm{1}[\Omega]]^{2}\geq(1-o(1))\mathbb{E}[E_{n}]^{2}.$
@@
 6.2.1. Our first reduction
-The goal of this section is to establish our reduction that showing $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$ is enough to establish $\mathbb{E}[E_{n}^{2}],\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$.
+The goal of this section is to establish our reduction that showing $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$ is sufficient to establish $\mathbb{E}[E_{n}^{2}],\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$.
 Lemma 6.6.
 Let $X_{n}$ denote the number of cut points or the graph distance. Assume that there is some $\epsilon>0$ such that for all $n$, we have
-(6.2) $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n}).$
+(6.2) $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n}).$
 Then, we must have that,
@@
 Proof.
-Fix any $\kappa>0$, we will first show that for $n$ large enough, we must have that,
+Fix any $\kappa>0$; we will first show that for $n$ large enough, we must have that,
 $\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2-\kappa}}.$
@@
 $\mathrm{Var}(X_{2n})=\mathrm{Var}(X_{n})+\mathrm{Var}(X_{[n,2n]})+\mathrm{Var}(\mathcal{E}_{n})+\text{Cov}(\mathcal{E}_{n},X_{n}+X_{[n,2n]}).$
-We then have that using the previous expansion on the variance and equation (6.2) that,
-$\frac{(2n)^{2}}{(\log(2n))^{f(2n)}}$ $\leq\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\frac{n^{2}}{(\log n)^{f(n)}}$
-$+\sqrt{\frac{(\log n)^{-\epsilon}n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}}\sqrt{\frac{2n^{2}}{(\log n)^{f(n)}}}.$
+By using the previous expansion of the variance and equation (6.2), we obtain
+$\frac{(2n)^{2}}{(\log(2n))^{f(2n)}}$ $\leq\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+(\log n)^{-\epsilon}\frac{n^{2}}{(\log n)^{f(n)}}$
+$+\sqrt{\frac{(\log n)^{-\epsilon}n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}}\sqrt{\frac{2n^{2}}{(\log n)^{f(n)}}}.$
 We can bound the right-hand side as,
-$\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+\frac{n^{2}}{(\log n)^{f(n)+\epsilon}}+C\frac{n^{2}}{(\log n)^{f(n)+\epsilon/2}}+C\frac{n^{2}\sqrt{\log\log n}}{(\log n)^{f(n)/2+1}}$
-$\leq\frac{(2+\delta)n^{2}}{(\log n)^{f(n)}}+C\frac{n^{2}\log\log n}{(\log n)^{2}},$
-where $\delta>0$ can be set to be any small constant and $C$ is some universal constant (that could depend on $\delta$) and could change from line to line.
+$\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+\frac{n^{2}}{(\log n)^{f(n)+\epsilon}}+C\frac{n^{2}}{(\log n)^{f(n)+\epsilon/2}}+C\frac{n^{2}\sqrt{\log\log n}}{(\log n)^{f(n)/2+1+o(1)}}$
+$\leq\frac{(2+\delta)n^{2}}{(\log n)^{f(n)}}+C\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}},$
+where $\delta>0$ can be set to be any small constant, and $C$ is some universal constant (that could depend on $\delta$) and could change from line to line.
 We consider two cases:
 Case 1: $f(n)\geq 2-\kappa$
-In this case, note that for $n$ large enough, we must have that $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{2-\kappa}}$. Thus, we obtain the inequality,
+In this case, note that for $n$ large enough, we must have that $\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}}\leq\delta\frac{n^{2}}{(\log n)^{2-\kappa}}$. Thus, we obtain the inequality,
 (6.3) $\frac{4n^{2}}{(\log 2n)^{f(2n)}}\leq\frac{(2+2\delta)n^{2}}{(\log n)^{2-\kappa}},$
@@
 Case 2: $f(n)\leq 2-\kappa$
-In this case, when $n$ is large enough, we can still obtain the inequality for any small $\delta>0$, $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{f(n)}}$.
+In this case, when $n$ is large enough, we can still obtain the following inequality for any small $\delta>0$: $\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}}\leq\delta\frac{n^{2}}{(\log n)^{f(n)}}$.
 Thus, we obtain the inequality,
@@
 where $c>0$ will be some universal constant, and the least line will hold provided $n$ is sufficiently large.
-Now, let $n_{0}$ be large enough integer such that both equations (6.3) and (6.4) hold. Furthermore, consider $k$ large enough so that
+Now, let $n_{0}$ be a large enough integer such that both equations (6.3) and (6.4) hold. Furthermore, consider $k$ large enough so that
 $\sum_{j=0}^{k-1}\frac{c}{\log\log(2^{j+1}n_{0})}\geq 2-\kappa.$
@@
 That is $\tilde{S}_{i}=S_{n+i}-S_{n}$.
-In this subsection, we will give the estimate only for cut points and the graph distance.
 Lemma 6.8.
-Let $X_{n}$ denote the number of cut points, or the graph distance. Assume that there exists some $\epsilon>0$ such that with probability significantly less than $(\log n)^{-2}$, we have that,
+Let $X_{n}$ denote the number of cut points, or the graph distance. Assume that there exists some $\epsilon>0$ such that, with probability significantly less than $(\log n)^{-2}$, we have that,
 (6.5) $\mathbb{P}(\tilde{S}[n/(\log n)^{4},n]\cap\hat{S}\neq\emptyset|\hat{S})\geq(\log n)^{-\epsilon}.$
@@
   1. (1)
-    Notice that we have $E_{n}>l$ implies that $\tilde{X}_{n}\geq\frac{l}{2}$ or $\hat{X}_{n}\geq\frac{l}{2}$. Furthermore, $\Omega$ also holds by Lemma 6.3. Then,
+    Notice that $E_{n}>l$ implies that $\tilde{X}_{n}\geq\frac{l}{2}$ or $\hat{X}_{n}\geq\frac{l}{2}$. Furthermore, $\Omega$ also holds by Lemma 6.3. Then,
     $\{E_{n}>l\}\subset(\{\tilde{X}_{n}\geq l/2\}\cap\Omega)\cup(\{\hat{X}_{n}\geq l/2\}\cap\Omega).$
   2. (2)
-    Furthermore, (again still considering that $E_{n}>l$) in the case $\tilde{X}_{n}\geq\frac{l}{2}$, it must be the case that there is an intersection between $\tilde{S}{[n/(\log n)^{4},n]}$ and $\hat{S}$ by the proof of Lemma 6.3.
+    Furthermore, (again still considering that $E_{n}>l$), in the case $\tilde{X}_{n}\geq\frac{l}{2}$, there must be an intersection between $\tilde{S}{[n/(\log n)^{4},n]}$ and $\hat{S}$ by the proof of Lemma 6.3.
 (Recall, we needed the event $\Omega$ to hold in order to ensure that there is a large reduction in the graph distance.)
@@
 […]
-Furthermore, observe that,
+Furthermore, observe,
+[…]
 […]
@@
 We used $\mathbb{E}[\tilde{X}_{n}]\leq\frac{n}{(\log n)^{1/2+o(1)}}$ to show the final inequality.
-(Note that in [11, Lemma 2.2], they stated that it should be obtained for the graph distance if one follows the proof of [21].) Now, assume that equation (6.5) holds with probability significantly less than $(\log n)^{-2}$.
+(Note that in [11, Lemma 2.2], they stated that this bound should be obtained for the graph distance if one follows the proof of [21].) Now, assume that equation (6.5) holds with probability significantly less than $(\log n)^{-2}$.
 Let $\Omega_{B}$ be the set of walks $\tilde{S}$ such that equation (6.5) holds.
@@
 $\mathbb{P}(\hat{S}\cap\tilde{S}[\frac{n}{(\log n)^{4}},n]\neq\emptyset|\tilde{S})\leq(\log n)^{-c}.$
-We start with the following observation.
-We define a sequence of scales $0=K_{0}<K_{1}<K_{2}<K_{3}<\ldots<K_{f}$ and consider the following bound.
-Also, let $A(R_{o},R_{i})$ denote the annulus around the origin with outer radius $R_{o}$ and inner radius $R_{i}$.
+We start by defining a sequence of scales $0=K_{0}<K_{1}<K_{2}<K_{3}<\ldots<K_{f}$; also, let $A(R_{o},R_{i})$ denote the annulus around the origin with outer radius $R_{o}$ and inner radius $R_{i}$.
 Let $\mathcal{A}_{i,n}:=A[\sqrt{n/K_{i}},\sqrt{n/K_{i+1}}]$.
@@
 (6.8) $+$ $\sum_{i=11}^{f-1}\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\max_{z\in\mathcal{A}_{i,n}}G(z).$
-Our heuristics for (6.8) are the following:
+Our heuristics for analyzing (6.8) are the following:
   1. (1)
@@
     For (6.7), we will use improved estimates for the capacity of random walks.
-Eventually, we will choose $K_{i}=\frac{1}{(\log n)^{i(1-\epsilon)}}$ for $i\leq 11$, and after this we choose $K_{i}=2K_{i-1}$ for $i\geq 12$.
+Eventually, we will choose $K_{i}=\frac{1}{(\log n)^{i(1-\epsilon)}}$ for $i\leq 11$, and after this, we will choose $K_{i}=2K_{i-1}$ for $i\geq 12$.
 Furthermore, we can choose $K_{f}=n^{1/10}$.
-Note that by the local central limit theorem, the probability a point $\tilde{S}_{k}$ for $k\in[n/(\log n)^{4},n]$ lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{n^{2/10}}{\frac{n^{2}}{(\log n)^{8}}}$.
+Note that by the local central limit theorem, the probability that a point $\tilde{S}_{k}$ for $k\in[n/(\log n)^{4},n]$ lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{n^{2/10}}{\frac{n^{2}}{(\log n)^{8}}}$.
 Taking a union bound, the probability that there is a point in $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ that lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{nn^{1/5}(\log n)^{8}}{n^{2}}\ll(\log n)^{-2}$.
@@
 $L(\mathcal{B}_{r}):=\sum_{t=\frac{n}{(\log n)^{4}}}^{n}\mathbbm{1}[S_{t}\in\mathcal{B}_{r}].$
-Namely, it is the amount of times that $S_{t}$ lies in the ball $\mathcal{B}_{r}$.
+Namely, it is the number of times that $S_{t}$ lies in the ball $\mathcal{B}_{r}$.
 Then, we have that
@@
 Proof.
-By using the local central limit theorem, one observes that the probability of a point $S_{t}$ being in the ball $\mathcal{B}_{r}$ is of order $\frac{r^{4}}{t^{2}}$.
+By using the local central limit theorem, the probability that a point, $S_{t}$, lies in the ball $\mathcal{B}_{r}$ is of order $\frac{r^{4}}{t^{2}}$.
 If we sum up this probability over all times $t$ from $n/(\log n)^{4}$ and $n$, we see that the expectation of the local time within this ball is $\frac{r^{4}(\log n)^{4}}{n}$.
@@
 By using subadditivity of the capacity and the fact that the capacity must be less than the volume, we must have that
-$\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\leq|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|$
-$\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}.$
+$\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\leq|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}.$
 Now, since $\max_{z\in\mathcal{A}_{i,n}}G(z)\leq C\frac{K_{i+1}}{n}$, then we have,
@@
 Proof.
-In [12, Lemma 4.6], we have that these are $c>0$ and $C<\infty$ such that
-$\mathbb{E}\left[\exp\left(c\frac{(\log n)^{2}}{n}({\rm Cap}(S[0,n])-\mathbb{E}[{\rm Cap}(S[0,n])])\right)\right]\leq C$
+In [12, Lemma 4.6], we have that there is $c>0$ such that
+$\limsup_{n\to\infty}\mathbb{E}\left[\exp\left(c\frac{(\log n)^{2}}{n}({\rm Cap}(S[0,n])-\mathbb{E}[{\rm Cap}(S[0,n])])\right)\right]<\infty$
 and $\mathbb{E}[{\rm Cap}(S[0,n])]\sim\frac{\pi^{2}}{8}\frac{n}{(\log n)}$ (e.g., see [6, Theorem 1.1]). Hence, we can derive (6.12) by applying Markov’s inequality. ∎
-In order to discuss the properties of the parts of the random walk inside each annulus, we need to consider the following stopping times.
-Define
-$T^{i}_{k}=\inf\{t\geq\tau^{i}_{k-1}:S_{t}\leq\sqrt{n/K_{i}}\},$
-$\tau^{i}_{k}=\inf\{t\geq T^{i}_{k}:S_{t}\geq 2\sqrt{n/K_{i}}\}.$
-These are a sequence of times indicating when the random walk hits that ball of radius $\sqrt{n/K_{i}}$ and later details when it escapes said ball.
-The following lemma will ensure that with high probability that there are not too many times in which the random walk will make an excursion into the ball of radius $\sqrt{n/K_{i}}$.
+In order to discuss the properties of the parts of the random walk inside each annulus, we need to consider the following stopping times. Define
+$T^{i}_{k}=\inf\{t\geq\tau^{i}_{k-1}:S_{t}\leq\sqrt{n/K_{i}}\},\quad\tau^{i}_{k}=\inf\{t\geq T^{i}_{k}:S_{t}\geq 2\sqrt{n/K_{i}}\}.$
+These are a sequence of times that indicate when the random walk hits that ball of radius $\sqrt{n/K_{i}}$, and later details when it escapes that ball.
+The following lemma will ensure that, with high probability, there are not too many times in which the random walk will make an excursion into the ball of radius $\sqrt{n/K_{i}}$.
 Furthermore, each excursion into this ball will lie very close to its expected time.
 Lemma 6.12.
-Let $M^{i}=\sup\{k:\tau^{i}_{k}\leq n\}$.
-Let $\mathcal{H}^{i}$ be the event that
+Let $M^{i}=\sup\{k:\tau^{i}_{k}\leq n\}$. Let $\mathcal{H}^{i}$ be the event that
 $\mathcal{H}^{i}:=\{M^{i}\leq(\log n)^{C}\}\bigcap_{k=1}^{(\log n)^{C}}\{|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1-\kappa}}\}\bigcap_{k=1}^{(\log n)^{C}}\{|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{n}{K_{i}^{1+\kappa}}\}.$
@@
 $\mathcal{S}^{i}:=\left\{L(\mathcal{B}_{2\sqrt{n/K_{i}}})\leq\frac{n}{(K_{i})^{1-\delta}}\right\},$
-where we recall the notation $L(\mathcal{B}_{r})$ from Lemma 6.10.
-By using equation (6.10) and Markov’s inequality with $k$ large enough, we would have that
+where we recall the notation $L(\mathcal{B}_{r})$ from Lemma 6.10. By using equation (6.10) and Markov’s inequality with $k$ large enough, we would have that
 $\mathbb{P}((\mathcal{S}^{i})^{c})\ll(\log n)^{-2}.$
-Now we consider the intersection of events,
-$\Omega^{i}=\mathcal{S}^{i}\cap\mathcal{G}^{i}\cap\mathcal{H}^{i}.$
+Now, we consider the intersection of events $\Omega^{i}:=\mathcal{S}^{i}\cap\mathcal{G}^{i}\cap\mathcal{H}^{i}$.
 By combining the probability bounds from earlier, we have $\mathbb{P}((\Omega^{i})^{c})\ll(\log n)^{-2}$.
 Notice that, on the event $\Omega^{i}$, we can show that,
-(6.13) ${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},\tau^{i}_{k}])\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}])$
+${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},\tau^{i}_{k}])\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}])$
 $\leq\sum_{k=1}^{M^{i}}\frac{n}{(\log n)^{1-\chi}K_{i}^{1-\kappa}}\leq K_{i}^{2\kappa}\sum_{k=1}^{M^{i}}\frac{n}{(\log n)^{1-\chi}K_{i}^{1+\kappa}}\leq\frac{K_{i}^{2\kappa}}{(\log n)^{1-\chi}}\sum_{k=1}^{M^{i}}|\tau^{i}_{k}-T^{i}_{k}|$
@@
 The fifth inequality used that $|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{1}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$.
-Note that the sum of these differences is a lower bound on the amount of times in the intersection between $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ and $\mathcal{B}_{2\sqrt{n/K_{i}}}$.
+Note that the sum of these differences is a lower bound on the number of times in the intersection between $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ and $\mathcal{B}_{2\sqrt{n/K_{i}}}$.
 The final inequality used the event $\mathcal{S}^{i}$.
@@
 $\leq\frac{K_{i}^{2\kappa}n}{(\log n)^{1-\chi}(K_{i})^{1-\delta}}\frac{K_{i+1}}{n}\leq(\log n)^{3(2\kappa+\delta)+\chi}(\log n)^{-\epsilon}.$
-By setting $3(2\kappa)+\delta+\chi\leq\epsilon/2$, we see that the line above will be less than $(\log n)^{-\epsilon/2}$.
+By setting $11(2\kappa+\delta)+\chi\leq\epsilon/2$, we see that the line above will be less than $(\log n)^{-\epsilon/2}$.
 We now consider $\hat{\Omega}:=\bigcap_{i=1}^{11}\Omega^{i}$, we must have $\mathbb{P}(\hat{\Omega}^{c})\ll(\log n)^{-2}$.
@@
 Now, consider the intersection $\tilde{\Omega}\cap\hat{\Omega}$, where we recall $\tilde{\Omega}$ from equation (6.11).
-We must have that $\mathbb{P}((\tilde{\Omega}\cap\hat{\Omega})^{c})\ll(\log n)^{-2}$ and, in addition, the right-hand side of equation (6.8) can be bounded by $(\log n)^{-c}$ for some $c>0$.
+We must have that $\mathbb{P}((\tilde{\Omega}\cap\hat{\Omega})^{c})\ll(\log n)^{-2}$ and, in addition, that the right-hand side of equation (6.8) can be bounded by $(\log n)^{-c}$ for some $c>0$.
 This completes the proof of Lemma 6.9.
@@
 We now derive the corresponding lower bound for $\mathrm{Var}(X_{n})$.
-Is this subsection, we will give the estimate only for cut points and graph distance.
 Lemma 6.13.
@@
 $X_{n}=\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}].$
-Notice, now that since $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]$ is a sum of i.i.d.
+Notice, that since $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]$ is a sum of i.i.d.
 terms where each individual term has variance at most $\frac{n^{2}}{(\log n)^{2-o(1)}}$, we can compute its variance as
@@
 Theorem 6.14.
-Let $X_{n}$ denote the number of cut points, or the graph distance. We have that as $n\to\infty$,
+Let $X_{n}$ denote the number of cut points or the graph distance. We have that, as $n\to\infty$,
 $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\to 0,$
@@
 Proof.
-We proceed with the same dyadic decomposition as we have done in the previous proof.
+We proceed with the same dyadic decomposition as in the previous proof.
 $X_{n}=\sum_{k=0}^{\epsilon\log_{2}(\log n)}\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}+\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}].$
@@
 Furthermore, $\mathbb{P}(\Omega_{A}\cap\Omega_{B})=1-o(1)$ by a union bound. This completes the proof. ∎
-6.5. An analysis of the effective resistance
+6.5. An Analysis of Variables with heavy tailed Cross Terms
 In this subsection, we present, for independent interest, a similar analysis done in the previous sections when we impose weaker conditions on the bounds we obtain for the variance and the expectation for the cross terms under consideration in $d=4$.
-Namely, we will assume that the cross-terms are very heavy tailed and that we have some almost optimal upper and lower bounds for the variance up to powers of $\log n$.
-In particular, we will give the proof of the desired result for the effective resistance in $d=4$ in Subsection 6.5.2.
-The following assumption will detail what we assume about the cross terms under consideration.
-Before this, in Subsection 6.5.1, we will give an estimate for the effective resistance, which assures that Assumption 6.17 parts (2) (3) will hold.
+Namely, we will assume that the cross terms are very heavy tailed and that we have some almost optimal upper and lower bounds for the variance up to powers of $\log n$.
+In particular, we will show that the effective resistance in $d=4$ satisfies these assumptions in Subsection 6.5.2.
+Assumption 6.16 will detail what we assume about the cross terms under consideration.
+Before this, in Subsection 6.5.1, we will briefly derive estimates for the effective resistance, which ensure that parts (2) and (3) of Assumption 6.16 will hold.
 6.5.1. Upper bound for the cross term of the effective resistance
-We will give the upper bound for the second moment of the cross term of the effective resistance $E[E_{n}^{2}]$.
+We will start by giving the upper bound for the second moment of the cross term of the effective resistance, $E[E_{n}^{2}]$.
 Lemma 6.15.
@@
 (6.14) $\mathbb{E}[E_{n}^{2}]\leq\frac{n^{2}}{(\log n)^{1-o(1)}}.$
-Remark 6.16.
-The constant that appears on the right-hand side of equation (6.15) is not the best possible (rather it should be improved to 1o(1)1-o(1)).
-However, obtaining the optimal constant is not necessary for the estimates that follow.
 Proof.
-Recall $\Omega$ in Lemma 6.3. On $\Omega^{c}$, we have that,
+Recall $\Omega$ from Lemma 6.3. The same proof will allow us to say that on $\Omega^{c}$, we have that,
 $E_{n}\leq 2\frac{n}{(\log n)^{b/2}}.$
-Hence, we obtain (6.14). Notice, also that since $E_{n}$ is deterministically less than $n$, we must have that,
+Hence, we obtain (6.13). Notice also that since $E_{n}$ is deterministically less than $n$, we must have that,
 $\mathbb{E}[E_{n}^{2}]=\mathbb{E}[E_{n}^{2}\mathbbm{1}[\Omega]]+\mathbb{E}[E_{n}^{2}\mathbbm{1}[\Omega^{c}]]\leq n^{2}\mathbb{P}(\Omega)+4\frac{n^{2}}{(\log n)^{b}}\leq\frac{n^{2}}{(\log n)^{1-o(1)}}.$
@@
 Proof of Proposition 1.4.
-By Lemmas 6.1 and 6.15, we obtain the result, which corresponds to (2) (3) for Assumption 6.17 for $b_{1}=2+o(1)$ and $b_{2}=1-o(1)$. ∎
+By Lemmas 6.1 and 6.15, which also hold for the effective resistance, we obtain results which correspond to parts (2) and (3) of Assumption 6.16 for $b_{1}=2+o(1)$ and $b_{2}=1-o(1)$. ∎
 6.5.2. General arguments
 Assumption 6.16.
-We let $X[a,b]$ for $a,b\in\mathbb{Z}^{+},b>a$ be a collection of random variables satisfying the following properties. First, define $E_{n}:=X_{n}+X[n,2n]-X_{2n}$
+We let $X[a,b]$ for $a,b\in\mathbb{Z}^{+},b>a$ be a collection of random variables satisfying the following properties. First, define $E_{n}:=X_{n}+X[n,2n]-X_{2n}.$
   1. (1)
@@
   2. (2)
-    If $a,b]\cap[c,d]=\emptyset$ independent, then $X[a,b]$ is independent of $X[c,d]$.
+    If $a,b]\cap[c,d]=\emptyset$, then $X[a,b]$ is independent of $X[c,d]$.
   3. (3)
@@
   4. (4)
-    There is a set $\Omega$ with probability $O(\frac{\log\log n}{\log n})$ such that
+    There is an event $\Omega$ with probability $O(\frac{\log\log n}{\log n})$ such that
     $E_{n}\mathbbm{1}[\Omega^{c}]\leq\mathbb{E}[E_{n}].$
@@
 $\mathrm{Var}(X_{n})=\frac{n^{2}}{(\log n)^{f(n)}}.$
-In order to characterize this, we need to first define the associated function $g(n)$ as as follows:
+In order to characterize this, we need to first define the associated function $g(n)$ as follows:
 $\mathbb{E}[E_{n}^{2}]:=\frac{n^{2}}{(\log n)^{g(n)}}.$
@@
 Thus, under the condition that $E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$ is negative, the expectation of $\mathbb{E}[(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])^{2}\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]^{1/2}\leq\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$.
-The second to last inequality used the definition of $g$ along with equation (6.16) for $n$ large enough.
-Finally, the last inequality used the fact that $\frac{n}{(\log n)^{k}}$ is increasing for $n$ large enough as well as the definition of $g(n,K)$.
+The second-to-last inequality used the definition of $g$ along with the equation (6.15) for $n$ large enough.
+Finally, the last inequality used the fact that $\frac{n}{(\log n)^{k}}$ is increasing for $n$ large enough, as well as the definition of $g(n,K)$.
 This should contrast with the positive contribution coming from,
@@
 The factor of $o(1)$ is to compensate for possible changes of the constant factor in front.
-We remark that due to these computations, there is a gap of of a factor of $(\log n)^{1/2}$ between the positive term above and the negative contribution from equation (6.17).
+We remark that due to these computations, there is a gap of a factor of $(\log n)^{1/2}$ between the positive term above and the negative contribution from equation (6.16).
 This will ensure a sufficiently good lower bound for the expectation of $\overline{\mathcal{E}}$.
@@
 In the first inequality, we used the fact that $\mathcal{I}$ is a sum of independent random variables.
-Since $g(n,K)$ lies between $b_{2}+o(1)$ and $b_{1}+o(1)$, the final inequality must be true for large enough $n$.
-Notice now, that we have, by the Cauchy-Schwarz inequality,
+Since $g(n,K)$ lies between $b_{2}+o(1)$ and $b_{1}+o(1)$, the final inequality must be true for a large enough $n$.
+Finally, notice that, by the Cauchy-Schwarz inequality,
 (6.19) $\mathbb{E}[(\overline{\mathcal{E}}+\mathcal{I})^{2}]\geq(\mathbb{E}[\overline{\mathcal{E}}^{2}]^{1/2}-\mathbb{E}[\mathcal{I}^{2}]^{1/2})^{2}\geq\frac{n^{2}}{(\log n)^{g(n,K)+o(1)}}.$
@@
   • We start by taking the high probability event $\Omega$ to be the event that all $E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]\leq\mathbb{E}[E_{n}^{(k,l)}]$.
-  • If there exists an infinite sequence of $m$ such that we have $f(n2^{-K})$ is greater than $g(n,K)+\frac{1}{4}$, then we have $f(n)\leq g(n,K)-o(1)$.
+  • If there exists an infinite sequence of $m$ such that $f(n2^{-K})$ is greater than $g(n,K)+\frac{1}{4}$, then $f(n)\leq g(n,K)-o(1)$.
 For such $n$, we will be able to show that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ will go to 0.
@@
 Theorem 6.19.
-Let $X_{n}$ be a series of random variables that satisfies the assumptions of 6.17. We have that there exists a subsequence such that as $n\to\infty$,
+Let $X_{n}$ be a series of random variables that satisfy the assumptions of 6.16. We have that there exists a subsequence such that as $n\to\infty$,
 $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\to}}0.$
@@
 $X_{n}=\sum_{l=1}^{2^{K}}\tilde{X}_{(K,l)}-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}.$
-Here, $\tilde{X}_{(K,l)}=X[(l-1)n2^{-K},ln2^{-K}]$ is a collection of independent copies with the same distribution as $X_{n2^{-K}}$ and for each given $K$, $E_{n}^{(k,l)}$ is a collection of independent random variables with the same distribution as $E_{n2^{-k}}$.
+Here, $\tilde{X}_{(K,l)}=X[(l-1)n2^{-K},ln2^{-K}]$ is a collection of independent copies with the same distribution as $X_{n2^{-K}}$, and for each given $K$, $E_{n}^{(k,l)}$ is a collection of independent random variables with the same distribution as $E_{n2^{-k}}$.
 We now define the event,
 $\Omega_{(k,l)}=\left\{|E_{n}^{(k,l)}-\mathbb{E}[E_{n}^{(k,l)}]|\leq\mathbb{E}[E_{n}^{(k,l)}]\right\},\quad\Omega=\bigcap_{k,l}\Omega_{(k,l)}.$
-By the third Assumption 6.17, we see that $\Omega$ occurs with probability 1o(1)1-o(1). We define $\mathcal{E}=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$ and observe that,
+By the third part of Assumption 6.16, we see that $\Omega$ occurs with probability 1o(1)1-o(1). We define $\mathcal{E}=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$ and observe that,
 (6.20) $\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]$
@@
 $\leq$ $2^{2K+2}4\max_{(k,l)}(\mathbb{E}[E_{n}^{(k,l)}]^{2})\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}.$
-In the last equation, we used equation (6.16) to bound $\mathbb{E}[E_{n}^{(k,l)}]^{2}$ relative to $\mathbb{E}[(E_{n}^{(k,l)})^{2}]$ and we placed irrelevant constant factors into an $o(1)$ exponent of $\log n$ in the denominator.
+In the last equation, we used equation (6.15) to bound $\mathbb{E}[E_{n}^{(k,l)}]^{2}$ relative to $\mathbb{E}[(E_{n}^{(k,l)})^{2}]$, and we placed irrelevant constant factors into an $o(1)$ exponent of $\log n$ in the denominator.
 In addition, we have that,
@@
 when $n$ is sufficiently large.
-Indeed, we used that for $n$ large and fixed $K$, we can bound $\frac{1}{(\log n2^{-K})^{f(n2^{-K})}}\leq\frac{1}{(\log n)^{g(n,K)+\frac{1}{4}+o(1)}}$ where the $o(1)$ term in the exponent is introduced to from $\log n2^{-K}\to\log n$ in the denominator.
+Indeed, we used that for $n$ large and fixed $K$, we can bound $\frac{1}{(\log n2^{-K})^{f(n2^{-K})}}\leq\frac{1}{(\log n)^{g(n,K)+\frac{1}{4}+o(1)}}$, where the $o(1)$ term in the exponent comes from the change $\log n2^{-K}\to\log n$ in the denominator.
 Recall, we assume $f(n2^{-K})\geq g(n,K)+\frac{1}{4}.$ In the numerator, we bounded $(\log n)^{f(n)}$ by $(\log n)^{g(n,K)+o(1)}$.
-Combining this estimate of equation (6.22), we see that
+Combining this estimate with equation (6.21), we see that
 $\mathbb{E}\left[\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)^{2}\mathbbm{1}[\Omega]\right]\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
-Thus, there will be some set $\Omega^{\prime}$, which will still hold with probability 1o(1)1-o(1) such that,
+Thus, there will be some set $\Omega^{\prime}$, which will still hold with probability 1o(1)1-o(1), such that,
 $\left(\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])\right)<\epsilon\sqrt{\mathrm{Var}(X_{m})}.$
-Furthermore, recall from equation (6.21) that the sums of the cross terms $\mathcal{E}$ satisfies,
+Furthermore, recall from equation (6.20) that the sums of the cross terms $\mathcal{E}$ satisfy,
 $\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
-Again, there will be a further subset $\Omega^{\prime\prime}\subset\Omega^{\prime}$ that will still hold with probability 1o(1)1-o(1) on which we have additionally that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|<\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
+Again, there will be a further sub-event $\Omega^{\prime\prime}\subset\Omega^{\prime}$, that will still hold with probability 1o(1)1-o(1), on which we have that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|<\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
 Thus, on $\Omega^{\prime\prime}$, we have $|X_{n}-\mathbb{E}[X_{n}]|<2\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
@@
 This can be divided into two more subcases.
-Case 2a: There does not exists an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
+Case 2a: There does not exist an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
 In this case, we must be able to find a sequence $f(a2^{mK})$ such that
@@
 Now, since $\sum_{m=1}^{\infty}\frac{1}{\log\log(a2^{mK})}$ diverges, this must mean that $f(a2^{mK})$ is unbounded from above.
-This contradicts our assumption that $f(n)\leq g(n,K)+\frac{1}{4}$ for $x$ large enough.
+This contradicts our assumption that $f(n)\leq g(n,K)+\frac{1}{4}$ for $n$ large enough.
 Thus, we do not have to worry about this case.
@@
 $\frac{(\log n)^{f(n)}}{(\log n2^{-K})^{f(n2^{-K}})}\leq(\log n)^{f(n)-f(n2^{-K})}\left(\frac{\log n}{\log n-K\log 2}\right)^{f(n2^{-K})}$
-$\leq$ $(\log n)^{\frac{1}{\log\log n}}\left(\frac{\log n}{\log n-K\log 2}\right)^{2+\frac{1}{4}}\leq 2e,$
-where the final inequality will hold for large enough $m$.
-Notice that we also used that $f(n2^{-K})\leq g(n,K)+\frac{1}{4}\leq\frac{3}{2}+\frac{1}{4}+o(1)$, by our assumption that eventually all $f(n)\leq g(n,K)+\frac{1}{4}$ and our upper bound on $g(n,K)$.
+$\leq$ $(\log n)^{\frac{1}{\log\log n}}\left(\frac{\log n}{\log n-K\log 2}\right)^{b_{2}+\frac{1}{4}}\leq 2e,$
+where the final inequality will hold for large enough $n$.
+Notice that we also used $f(n2^{-K})\leq g(n,K)+\frac{1}{4}\leq b_{2}+\frac{1}{4}+o(1)$, by our assumption that eventually all $f(n)\leq g(n,K)+\frac{1}{4}$ and our upper bound on $g(n,K)$.
 Returning to equation (6.21), we can bound the expectation of the square of $\sum_{l=1}^{2^{K}}(\tilde{X}_{(K,l)}-\mathbb{E}[\tilde{X}_{(K,l)}])$ by $2e2^{-K}\mathrm{Var}(X_{n})$ for large enough $n$.
@@
 $\mathbb{E}[(\mathcal{E}-\mathbb{E}[\mathcal{E}])^{2}\mathbbm{1}[\Omega]]\leq\frac{n^{2}}{(\log n)^{g(n,K)+1-o(1)}}\ll\frac{n^{2}}{(\log n)^{f(n)}}.$
-Again, here we used that first we are considering an $m$ such that first we have $f(n)\leq f(n2^{-K})+\frac{1}{\log\log n}$.
+Again, here we used that first we are considering an $n$ such that first we have $f(n)\leq f(n2^{-K})+\frac{1}{\log\log n}$.
 Secondly, under the general assumption of case 2, we have $f(n2^{-K})+\frac{1}{\log\log n}\leq g(n,K)+\frac{1}{4}+\frac{1}{\log\log n}\leq g(n,K)+1+o(1)$ for $n$ large enough.
 Thus, we have $f(n)\leq g(n,K)+1+o(1)$ for some infinite sequence of $n$.
-Thus (again by Markov’s inequality on the set $\Omega$), with probability $o(1)$, we have that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
-In total, we see that with probability at most $o(1)+2e2^{-K}\epsilon^{-2}$, we have that
+Thus (again by Markov’s inequality on the event $\Omega$), with probability $o(1)$, we have $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
+In total, we see that with probability at most $o(1)+2e2^{-K}\epsilon^{-2}$,
 $|X_{n}-\mathbb{E}[X_{n}]|\geq 2\epsilon\sqrt{\mathrm{Var}(X_{n})}.$
@@
 Notice that in all cases, for a fixed $K$, we can find a subsequence $X_{n_{1}},X_{n_{2}},\ldots$ such that with probability at least $1-4e2^{K(1-q)}\epsilon^{-2}$, we must have that
-$|X_{n_{i}}-\mathbb{E}[X_{n_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{n_{i}})}$
+$|X_{n_{i}}-\mathbb{E}[X_{n_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{n_{i}})}.$
 As $K\to\infty$, we see that $1-4e2^{-K}\epsilon^{-2}\to 0$.
@@
   • [5] K. Burdzy and G. F. Lawler. Rigorous exponent inequalities for random walks. J. Phys. A Math. Gen. 23(1), L23–L28 1990.
-  • [6] Y. Chang. Two observations on the capacity of the range of simple random walk on Z3 and Z4. Electron. Com. Probab. 22 (2017), No. 25, 9 pp.
+  • [6] Y. Chang. Two observations on the capacity of the range of simple random walk on $\mathbb{Z}^{3}$ and $\mathbb{Z}^{4}$. Electron. Com. Probab. 22, No. 25, 2017.
   • [7] D. A. Croydon. Random walk on the range of random walk. J. Stat. Phys. 136(2):349–372, 2009.